### v1

# Jarník-Besicovitch type theorems for semisimple algebraic groups

Abstract.
In this note, we initiate a study on Jarník-Besicovitch type theorems for semisimple algebraic groups from the representation-theoretic point of view. Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible $\mathbb{Q}$-rational representation of a connected semisimple $\mathbb{Q}$-algebraic group $\mathbf{G}$ on a complex vector space $V$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in a $\mathbb{Q}$-split torus in $\mathbf{G}$. We define a subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ in terms of the representation $\rho$ and the subgroup $\{a_{t}\}_{t\in\mathbb{R}}$, and prove formulas for the Hausdorff dimension of the complement of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$. As corollaries, we deduce several Jarník-Besicovitch type theorems.
2020 Mathematics Subject Classification:
Primary: 37A17; Secondary: 11J83
The author acknowledges the support by the Institute of Modern Analysis-A Frontier Research Centre of Shanghai.
1. Introduction and main results
1.1. Introduction
In this note, we initiate a study on the Jarník-Besicovitch theorem in the metric theory of Diophantine approximation. Recall that a real number $x\in\mathbb{R}$ is Diophantine of type $\kappa>0$ if there exists a constant $C>0$ such that
$\left|nx-m\right|\geq\frac{C}{n^{\kappa}}\;\left(\forall\frac{m}{n}\in\mathbb{Q}\right).$
Then necessarily $x\in\mathbb{R}\setminus\mathbb{Q}$ and the Dirichlet theorem implies that $\kappa\geq 1$. Denote by $S_{\kappa}$ the set of all Diophantine numbers of type $\kappa$ and $S_{\kappa}^{c}$ its complement. Then one can deduce from the Khintchine’s theorem on metric Diophantine approximation that $S_{\kappa}^{c}$ is a set of full measure if $\kappa=1$, and is null if $\kappa>1$. A refined result due to Jarník and Besicovitch [4, 22] states further that the Hausdorff dimension of $S_{\kappa}^{c}$ is equal to $2/(\kappa+1)$ ($\kappa>1$).
From the viewpoint of dynamical systems, the Jarník-Besicovitch theorem can be reformulated as a result of shrinking target problem, which was first observed by Hill and Velani [20]. Let $f:X\to X$ be a map on a metric space $X$ with a measure $\mu$. Generally speaking, in the shrinking target problem, one studies the set $S$ of points in $X$ whose trajectories under $f$ approach to a shrinking target infinitely often with certain rate, and seeks to establish results about the size (the measure or the Hausdorff dimension) of the set $S$. The measure theoretic version of the shrinking target problem is usually related to the Khintchine’s theorem [26, 41] while the Hausdorff dimension version is linked with the Jarník-Besicovitch theorem [20].
In homogeneous dynamics, various Diophantine sets can also be studied by the dynamical and ergodic properties of group actions on homogeneous spaces via Dani’s correspondence [14, 26]. Specifically, let $X_{m+n}=\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})$,
$a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})\quad(t\in\mathbb{R})$
[…]
$g\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})\to g\cdot\mathbb{Z}^{m+n}\quad(g\in\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})).$
Let $\delta:X_{m+n}\to\mathbb{R}_{+}$ be the systole function on $X_{m+n}$ defined by
$\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|$
which is also called the first minimum of the lattice $\Lambda\in X_{m+n}$. For any $m\times n$ matrix $A$, denote by
$u_{A}:=\left(\begin{array}[]{ccc}I_{m}&A\\ O&I_{n}\end{array}\right)\in N_{+}.$
Then it is well-known via Mahler’s compactness criterion and Dani’s correspondence [14, 26] that $A$ is badly approximable if and only if the orbit $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ is bounded in $X_{m+n}$, i.e.
$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\geq c\quad(t\geq 0)$
for some constant $c>0$; singular if and only if $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ diverges i.e.
$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\to 0\quad(\textup{as }t\to\infty);$
and very well approximable if and only if
$\limsup_{t\to\infty}\frac{-\log(\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n}))}{t}>0.$
We say that an $m\times n$ matrix $A$ is Diophantine of type $\kappa$ if there exists a constant $C>0$ such that
$\|Aq-p\|\geq C\|q\|^{-\kappa}$
for any $q\in\mathbb{Z}^{n}\setminus\{0\}$ and $p\in\mathbb{Z}^{m}$. Then necessarily $\kappa\geq n/m$, and again using Dani’s correspondence, one can deduce that $A$ is Diophantine of type $\kappa$ if and only if there exists a constant $C>0$ such that
$\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{m+n}))\geq Ce^{-\tau t}\;(\forall t>0)$
where
$\tau=\frac{m\kappa-n}{mn(\kappa+1)}.$
The set $S_{\kappa}$ of all Diophantine numbers of type $\kappa$ defined at the beginning of the introduction then corresponds to the case $m=n=1$ here. In particular, one can rephrase the Jarník-Besicovitch theorem for the homogeneous space $X_{m+n}=X_{2}$ $(m=n=1)$, and get that for any parameter $10\leq\tau<$ 1, the Hausdorff dimension of the complement of the set
$\{u_{A}\in N_{+}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{2}))\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $1-\tau$. Using the decomposition of $X_{2}$ into the stable, central and unstable submanifolds of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $X_{2}$, one can then obtain the following equivalent statement: the Hausdorff dimension of the complement of the set
$\{p\in X_{2}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $3-\tau$. In general, the matrix version of the Jarnik-Besicovitch theorem has also been established [9, 15], and it is equivalent to the statement that for any $0\leq\tau<1/n$, the Hausdorff dimension of the complement of the set
$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $mn(1-\tau)+m^{2}+n^{2}+mn-1$. We remark that one can further consider the case where $\{a_{t}\}_{t\in\mathbb{R}}$ is a generic one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})$ and study the Hausdorff dimension of the complement of the set
$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}.$
This type of question leads to a general weighted multidimensional Jarník-Besicovitch theorem on metric Diophantine approximation, which is one of the main topics we pursue in this note (See Theorem 1.6).
[…]
$\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|.$
Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter Ad-diagonalizable subgroup in $\mathbf{G}(\mathbb{R})$. Then the main object we would like to study is the Hausdorff dimension of the complement of the set
$\{g\in\mathbf{G}(\mathbb{R}):\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for some }C>0\}.$
Clearly the discussions above about the Jarnik-Besicovitch theorem with the homogeneous flow $a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})$ on $X_{m+n}$ can be reduced to this setting if we let $\mathbf{G}=\operatorname{\mathbf{SL}}_{m+n}$ and $\rho$ the standard representation of $\operatorname{\mathbf{SL}}_{m+n}$ on $\mathbb{C}^{m+n}$. Moreover, we will see later (Theorem 1.8) that by considering representations of algebraic groups, one will be able to establish some Jarnik-Besicovitch type theorems for certain algebraic subvarieties in an affine space. We remark that the question we study in this note starts from the sparse equidistribution problem, and a related shrinking target problem plays an important role. For more details, one can read [17, 46, 47, 48].
1.2. Main results
In the rest of this note, we address the question above under the following assumption. We assume that $\mathbf{G}$ is a connected semisimple $\mathbb{Q}$-algebraic group (in the Zariski topology) and $\mathbf{T}$ is a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$, and $\rho$ a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Let $d=\dim V$ and we may identify $V\cong\mathbb{C}^{d}$ so that $\mathbb{Z}^{d}\subset\mathbb{C}^{d}$ is compatible with the $\mathbb{Q}$-structure in $V$.
Definition 1.1.
An element $g\in\mathbf{G}(\mathbb{R})$ is Diophantine of type $\tau\geq 0$ if there exists a constant $C>0$ such that
$\delta(\rho(a_{t}\cdot g)\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for any }t>0$
where $\delta$ is the first minimum function. We denote by $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ the set of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ and $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$ its complement in $\mathbf{G}(\mathbb{R})$. If the representation $\rho$ and the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ are clearly stated in the contexts, we will simply write $S_{\tau}$ instead of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$.
Remark 1.2.
If $\mathbf{G}$ is $\mathbb{Q}$-anisotropic, then $\mathbf{T}=\{e\}$. So in the following, we assume that $\mathbf{G}$ is $\mathbb{Q}$-isotropic.
To state the first main theorem about $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$, we need to introduce some notation. We choose a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$. Then $\mathbf{P}_{0}$ and $\mathbf{T}$ defines a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ defined by $\Phi\setminus\Phi^{+}$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. One can write the space $V$ in the representation $\rho$ as a direct sum of weight spaces with respect to the action of $\mathbf{T}$
$V=\bigoplus_{\beta}V_{\beta}.$
By the structure theory of irreducible representations of complex semisimple groups and semisimple Lie algebras, there is a highest weight $\beta_{0}$ among the weights $\beta$’s (where the order is determined by the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$) and we denote by $V_{\beta_{0}}$ its corresponding weight space. (See §2 for more details). The stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$ is a parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{\beta_{0}}$ containing $\mathbf{P}_{0}$, and its unipotent radical is denoted by $R_{u}(\mathbf{P}_{\beta_{0}})$. The opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ and its unipotent radical are denoted by $\overline{\mathbf{P}}_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ respectively.
In the following, if an algebraic $\mathbb{Q}$-subgroup $\mathbf{F}\subset\mathbf{G}$ is normalized by $\mathbf{T}$, then we write $\Phi(\mathbf{F})$ for the set of $\mathbb{Q}$-roots in $\mathbf{F}$ relative to $\mathbf{T}$ while the summation $\sum_{\alpha\in\Phi(\mathbf{F})}$ (or $\prod_{\alpha\in\Phi(\mathbf{F})}$) means that we take the sum (or product) all over $\alpha\in\Phi(\mathbf{F})$ counted with multiplicities (i.e., the dimensions of the corresponding $\mathbb{Q}$-root spaces in the Lie algebra of $\mathbf{F}$). For a $\mathbb{Q}$-root or a $\mathbb{Q}$-weight $\lambda$ in $\rho$ relative to $\mathbf{T}$ (which is a $\mathbb{Q}$-character of $\mathbf{T}$) , we will often consider it as a linear functional on the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ and use the same symbol. In particular, we will write $\lambda(a_{t})$ ($t\in\mathbb{R}$) for the values of $\lambda$ (as a linear functional) on the Lie algebra of $\{a_{t}\}_{t\in\mathbb{R}}$ (so that $\lambda(a_{t})$ is linear on $t\in\mathbb{R}$). We denote by $\nu_{0}$ the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ relative to $\mathbf{T}$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
Now we can state the first main theorem in this note.
Theorem 1.3.
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Then
$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\geq$ $\dim\mathbf{G}-\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
for any $0\leq\tau<\beta_{0}(a_{-1})$.
To state the next theorem, we need to introduce another notation. Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
${}_{\mathbb{Q}}{W}=N(\mathbf{T})/Z(\mathbf{T}).$
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above, and we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. The Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
which implies that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
Let $\{w_{i}\}_{i\in I}$ be a set of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$ where $I$ is the index set of ${}_{\mathbb{Q}}W$. For each $w_{i}$ $(i\in I)$, define
$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
and we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}\quad(i\in I).$
Then one can further write the Bruhat decomposition as
$\mathbf{G}(\mathbb{Q})=$ $\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q})$
$=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
$=$ $\bigcup_{i\in I}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Note that the sets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$ $(i\in I)$ in the decomposition of $\mathbf{G}(\mathbb{Q})$ above may overlap, and for our purpose, we can choose any subset $\bar{I}$ of $I$ (as small as possible) such that
$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
In the statement of Theorem 1.4 below, we will fix any such subset $\bar{I}$ in $I$.
Let $\bigwedge^{\dim V_{\beta_{0}}}V$ be the $\dim V_{\beta_{0}}$-exterior product vector space of $V$ over $\mathbb{C}$, and $\rho_{\beta_{0}}$ the natural extension of $\rho$ on $\bigwedge^{\dim V_{\beta_{0}}}V$. Let
$\{e_{1},e_{2},\dots,e_{\dim V_{\beta_{0}}}\}\subset\mathbb{Z}^{d}$
be an integral basis in $V_{\beta_{0}}$ which spans $V_{\beta_{0}}\cap\mathbb{Z}^{d}$. We write
$e_{V_{\beta_{0}}}:=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}\in\bigwedge^{\dim V_{\beta_{0}}}V.$
For each $i\in\bar{I}$, define the following morphism
$\Psi_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\Psi_{w_{i}}$ is an isomorphism onto its image. We denote by $a_{w_{i}}$ the growth rate of the asymptotic volume estimate of the real variety $\textup{Im}(\Psi_{w_{i}})$, and by $A_{w_{i}}$ the growth rate of the number of rational points in $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})$ (See §6 Corollary 6.8 and equation ($\ast$6) for more details). Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
Now we can state the second main theorem in this note.
Theorem 1.4.
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Then
$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$
$\leq$ $\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{a_{w_{i}},A_{w_{i}}\}\cdot\dim V_{\beta_{0}}\Big{|}\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>\tau\right\}$
for any $0\leq\tau<\beta_{0}(a_{-1})$.
Remark 1.5.
It follows from the proof of Theorem 1.4 that for any $\tau\geq\beta_{0}(a_{-1})$, the subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=\emptyset$. (See §6.)
As corollaries of Theorems 1.3 and 1.4, we obtain the following results. The first result generalizes the Jarnik-Besicovitch type theorem on the homogneous space $X_{n}=\operatorname{\mathbf{SL}}_{n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{n}(\mathbb{Z})$ for any one-parameter diagonal flow.
Theorem 1.6.
Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the standard representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ on the complex vector space $V=\mathbb{C}^{n}$ defined by
$\rho(g)\cdot v=g\cdot v\quad(g\in\operatorname{\mathbf{SL}}_{n},v\in V)$
via matrix multiplication. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4. Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
The second result generalizes the main results in [17].
Theorem 1.7.
Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ where $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$, $V=\mathfrak{g}=\mathfrak{sl}_{n}$ is the Lie algebra of $\operatorname{\mathbf{SL}}_{n}$ and $\rho=\operatorname{Ad}$ is the adjoint representation of $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\mathbf{G}$, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4. Then for any $0\leq\tau<\nu_{0}(a_{1})$, we have
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}.$
We also deduce the following
Theorem 1.8.
Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible representation of $\operatorname{\mathbf{SL}}_{2}$ defined over $\mathbb{Q}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{2}$ defined as in Theorem 1.4. Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
$\dim_{H}S_{\tau}^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
Remark 1.9.
Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be the irreducible representation in Theorem 1.8 with $\dim V=n+1$ $(n\geq 1)$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ and let $\rho_{0}$ be the standard representation of $\operatorname{\mathbf{SL}}_{n+1}$ on $\mathbb{C}^{n+1}$. Note that $V\cong\mathbb{C}^{n+1}$ and $\rho(\operatorname{\mathbf{SL}}_{2})\subset\operatorname{\mathbf{SL}}_{n+1}\subset\operatorname{\mathbf{GL}}(V)$. Then by Theorem 1.8, we have
$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}\cap\rho(\operatorname{\mathbf{SL}}_{2}(\mathbb{R}))=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}$
where $\beta_{0}$ is the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3. On the other hand, by Theorem 1.6
$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}=\dim\operatorname{\mathbf{SL}}_{n+1}-\frac{(n+1)\cdot\tau}{\nu_{0}(\rho(a_{1}))}$
where $\nu_{0}$ is the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n+1}$ with respect to $\{\rho(a_{t})\}_{t\in\mathbb{R}}$ defined as in Theorem 1.4. This gives an example of Jarník-Besicovitch type theorems for some algebraic subvarieties in affine spaces.
Remark 1.10.
One can see from the arguments in this note that all the theorems stated above also work for any open bounded subset $U$ in $\mathbf{G}(\mathbb{R})$, i.e. $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\cap U=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$. (See Remark 5.6.)
1.3. Discussions and open problems
[…]
$\liminf_{t\to\infty}-\frac{1}{t}\log\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})=\tau.$
When $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and $\rho$ is the standard representation of $\mathbf{G}$, the Hausdorff dimension and the packing dimension of $\operatorname{Sing}(\tau)$ are discussed in [43] in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ is a singular one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, and the proofs rely on a variational principle in the parametric geometry of numbers. This variational principle is later generalized in [39] for any one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$.
For the lower bound and the upper bound we obtain in Theorem 1.3 and Theorem 1.4, we have computed in several cases (e.g. Theorems 1.61.7 and 1.8) that these two bounds are equal. It would be interesting to compute these two bounds in other cases and determine the representations $\rho$ for which these two bounds are equal.
2. Preliminaries
In this section, we list some preliminaries needed in this note. Then we give a definition of rational elements in $\mathbf{G}(\mathbb{R})$ and discuss some properties of rational elements.
We first need the reduction theory of arithmetic subgroups of $\mathbf{G}(\mathbb{R})$ [5]. Let $K$ be a maximal compact subgroup in $\mathbf{G}(\mathbb{R})$ and $\Gamma$ an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})$. We fix a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$. Denote by $\Phi$ the set of $\mathbb{Q}$-roots with respect to $\mathbf{T}$, $\Phi^{+}$ the set of positive $\mathbb{Q}$-roots corresponding to the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ and $\Delta$ the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. For any parabolic $\mathbb{Q}$-subgroup $\mathbf{P}$, denote by $R_{u}(\mathbf{P})$ its unipotent radical. Let $M$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\mathbf{G}(\mathbb{R})}(\mathbf{T}(\mathbb{R}))$ (= the centralizer of $\mathbf{T}(\mathbb{R})$ in $\mathbf{G}(\mathbb{R})$). For $\eta>0$, denote by
$T_{\eta}=\{a\in\mathbf{T}(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\Delta\}.$
A Siegel set in $\mathbf{G}(\mathbb{R})$ is a subset of the form $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ for some $\eta>0$ and a relatively compact open subset $\Omega$ containing identity in $M\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R})$, and the group $\mathbf{G}(\mathbb{R})$ can be written as
$\mathbf{G}(\mathbb{R})=S_{\eta,\Omega}\cdot\mathcal{K}\cdot\Gamma$
for some Siegel set $S_{\eta,\Omega}$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$. Moreover, the finite set $\mathcal{K}$ satisfies the property that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
where $\mathbf{P}_{0}$ is the minimal parabolic $\mathbb{Q}$-subgroup. Denote by
$\mathcal{K}=\{x_{1},x_{2},\dots,x_{k}\}=\{x_{j}\}_{j\in J}\subset\mathbf{G}(\mathbb{Q})$
and we may assume that $e\in\mathcal{K}$. In what follows, we choose $\Gamma$ to be an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})\cap\mathbf{G}(\mathbb{R})^{0}$, where $\mathbf{G}(\mathbb{R})^{0}$ denotes the connected component of identity in the Lie group $\mathbf{G}(\mathbb{R})$, so that $\rho(\Gamma)$ preserves the lattice $\mathbb{Z}^{d}$ in $V$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$ and its unstable horospherical subgroup is contained in $R_{u}(\overline{\mathbf{P}}_{0})$ where $\overline{\mathbf{P}}_{0}$ is the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ determined by $\Phi\setminus\Phi^{+}$. We also write $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the centralizer of $\mathbf{T}$ in $\mathbf{G}$, and write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$ so that $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$.
Now we choose a maximal $\mathbb{Q}$-torus $\mathbf{S}$ in $\mathbf{P}_{0}$ containing $\mathbf{T}$. The Lie algebra $\mathfrak{g}$ of $\mathbf{G}$ can be written as a direct sum
$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Psi}\mathfrak{g}_{\alpha}$
of root spaces relative to $\mathbf{S}$ via the adjoint representation of $\mathbf{G}$, where $\Psi$ denotes the set of roots in $\mathfrak{g}$ relative to $\mathbf{S}$. We may determine a set of positive roots in $\Psi$, which we denote by $\Psi^{+}$, such that
$\operatorname{Lie}(R_{u}(\mathbf{P}_{0}))\subset\sum_{\alpha\in\Psi^{+}}\mathfrak{g}_{\alpha}\textup{ and }\operatorname{Lie}(R_{u}(\overline{\mathbf{P}}_{0}))\subset\sum_{\alpha\in\Psi\setminus\Psi^{+}}\mathfrak{g}_{\alpha}.$
The set of simple roots in $\Psi^{+}$ is denoted by $\Pi$.
It is known that there are complete classifications of irreducible representations of complex semisimple groups and Lie algebras [21, 27]. Let $V=\bigoplus_{\lambda}V_{\lambda}$ be the decomposition of the vector space $V$ in $\rho$ into the direct sum of weight spaces $V_{\lambda}$ relative to $\mathbf{S}$. According to the theorem of highest weight, there is a unique highest weight $\lambda_{0}$ among the weights $\lambda$’s (the order is determined by $(\Psi,\Psi^{+},\Pi)$) such that
  1. (1)
    $\dim_{\mathbb{C}}V_{\lambda_{0}}=1$.
  2. (2)
    for any $\alpha\in\Psi^{+}$, any $E_{\alpha}\in\mathfrak{g}_{\alpha}$ annihilates $V_{\lambda_{0}}$ via the differential $d\rho$ of $\rho$, any $n_{\alpha}\in\exp(\mathfrak{g}_{\alpha})$ fixes elements in $V_{\lambda_{0}}$ via $\rho$ (where $\exp$ is the exponential map), and elements of $V_{\lambda_{0}}$ are the only vectors with this property.
  3. (3)
    every weight $\lambda$ in $\rho$ is of the form $\lambda_{0}-\sum_{i=1}^{l}n_{i}\alpha_{i}$ where $n_{i}\in\mathbb{Z}_{\geq 0}$ and $\alpha_{i}\in\Pi$.
Note that by our choices of $\Phi$ and $\Psi$, we have
$\{\alpha|_{\mathbf{T}}:\alpha\in\Psi,\alpha|_{\mathbf{T}}\neq 0\}=\{\alpha:\alpha\in\Phi\}.$
On the other hand, one can also write $V$ as a direct sum of weight spaces relative to $\mathbf{T}$ as follows
$V=\bigoplus_{\beta}V_{\beta}.$
Let $\beta_{0}$ be the weight in the decomposition above such that its weight space $V_{\beta_{0}}$ contains $V_{\lambda_{0}}$. Note that $\beta_{0}$ is defined over $\mathbb{Q}$ and $\beta_{0}=\lambda_{0}|_{\mathbf{T}}$. In particular, $\lambda_{0}(\mathbf{T}(\mathbb{R}))\subset\mathbb{R}$ and $\beta_{0}$ is the highest weight among $\beta$’s relative to $\mathbf{T}$ (the order is determined by the root system $(\Phi,\Phi^{+},\Delta)$) (as discussed in §1).
Since the stable subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$, we have $\alpha(a_{t})\leq 0$ $(t\geq 0)$ for any $\alpha\in\Phi^{+}$. From property (3) of the representation $\rho$ above and the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, one can deduce that $\beta_{0}(a_{t})=\lambda_{0}(a_{t})\leq 0$. If $\beta_{0}(a_{t})=\lambda_{0}(a_{t})=0$, then by the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, for any weight $\lambda$ in $\rho$ relative to $\mathbf{S}$, we have $\lambda(a_{t})=0$ and $\{a_{t}\}_{t\in\mathbb{R}}\subset\ker\rho$, which contradicts the assumption that $\dim\ker\rho=0$. Therefore we have $\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0$ and
$\rho(a_{t})\cdot v=e^{\beta_{0}(a_{t})}\cdot v\to 0\textup{ as }t\to\infty$
for any $v\in V_{\beta_{0}}$ (here $\beta_{0}(a_{-1})>0$ is the fastest contracting rate under the action of $\{a_{t}\}_{t\in\mathbb{R}}$).
Let $\mathbf{P}_{\beta_{0}}$ be the stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$. Then $\mathbf{P}_{\beta_{0}}$ contains the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ since $\beta_{0}$ is the highest weight among the weights $\beta$’s relative to $\mathbf{T}$ in $\rho$. Therefore, $\mathbf{P}_{\beta_{0}}$ is a parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$. Let $\overline{\mathbf{P}}_{\beta_{0}}$ be the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$, and denote by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ the unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$.
Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
${}_{\mathbb{Q}}{W}=N(\mathbf{T})/Z(\mathbf{T}).$
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above. Then the Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
which implies that
$\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
If we choose a set $\{w_{i}\}_{i\in I}$ of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$, then for each $w_{i}$, define
$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
and we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$
Then one can further write the Bruhat decomposition as
$\mathbf{G}(\mathbb{Q})=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
$=$ $\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
where the index sets $I$ and $\bar{I}$ are defined as in §1. We denote by
$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
Note that the group $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}\subset\mathbf{P}_{\beta_{0}}$ is generated by unipotent subgroups whose Lie algebras are direct sums of $\mathbb{Q}$-root spaces in $\mathbf{G}$ relative to $\mathbf{T}$, and $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ by the structure of the representation $\rho$.
Definition 2.1.
An element $g\in\mathbf{G}(\mathbb{R})$ is called rational if $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is Zariski dense in $V_{\beta_{0}}$.
Lemma 2.2.
An element $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ is rational if and only if $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$.
Proof.
Let $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ be a rational element in $\mathbf{G}(\mathbb{R})$. Then by definition, there exists a discrete subgroup $\Lambda_{g}\subset\mathbb{Z}^{d}$ such that $\rho(g)\cdot\Lambda_{g}$ is a lattice in $V_{\beta_{0}}(\mathbb{R})$. Choose $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Since $V_{\beta_{0}}$ is defined over $\mathbb{Q}$, we have
$\sigma(\rho(g)\cdot\Lambda_{g})=\rho(\sigma(g))\cdot\Lambda_{g}\subset\sigma(V_{\beta_{0}}(\mathbb{R}))=V_{\beta_{0}}(\mathbb{R})$
and $\rho(\sigma(g))\cdot\Lambda_{g}$ is Zariski dense in $V_{\beta_{0}}$. Then we obtain that
$\rho(\sigma(g)g^{-1})V_{\beta_{0}}=\rho(\sigma(g)g^{-1})\overline{\rho(g)\cdot\Lambda_{g}}=V_{\beta_{0}}$
and $\sigma(g)g^{-1}$ is in the stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$. Here $\overline{\rho(g)\cdot\Lambda_{g}}$ denotes the Zariski closure of $\rho(g)\cdot\Lambda_{g}$. On the other hand, $\sigma(g)g^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cap\mathbf{P}_{\beta_{0}}=\{e\}$. Therefore, $\sigma(g)=g$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$, and $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$. The other direction is clear. ∎
Corollary 2.3.
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if $g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$.
Proof.
Let $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that $g=h_{i}\cdot f_{i}$. Suppose that $w_{i}\cdot g$ is rational. Since $w_{i}\in\mathbf{G}(\mathbb{Q})$, by definition, $w_{i}\cdot h_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is also rational in $\mathbf{G}(\mathbb{R})$. We know that $w_{i}\cdot h_{i}\cdot w_{i}^{-1}$ preserves $V_{\beta_{0}}$, so $w_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is rational. By the fact that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and Lemma 2.2, we conclude that $w_{i}\cdot f_{i}\cdot w_{i}^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$.
Conversely, suppose that $g=h_{i}\cdot f_{i}$ where $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Then
$w_{i}\cdot g=(w_{i}h_{i}w_{i}^{-1})\cdot(w_{i}f_{i}w_{i}^{-1})\cdot w_{i}.$
Note that $w_{i}h_{i}w_{i}^{-1}\in\mathbf{P}_{\beta_{0}}$ and $w_{i}\in\mathbf{G}(\mathbb{Q})$. By definition, we know that $w_{i}\cdot g$ is rational. This completes the proof of the lemma. ∎
By properties of the subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ in the reduction theory, we can give another characterization of rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Lemma 2.4.
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if
$g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma).$
Proof.
If $w_{i}\cdot g$ is rational, then by Corollary 2.3
$g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{G}(\mathbb{Q})=\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
Suppose that $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma)$. Then there exist $h\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
$g=h\cdot f,\quad f\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
We write
$f=h_{i}\cdot p\cdot x\cdot\gamma$
for some $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$. Let $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Then we have
$\sigma(f)=\sigma(h_{i})\cdot\sigma(p)\cdot x\cdot\gamma.$
This implies that $x\cdot\gamma=p^{-1}\cdot h_{i}^{-1}\cdot f=\sigma(p)^{-1}\cdot\sigma(h_{i})^{-1}\cdot\sigma(f)$. Since the product map
$\mathbf{P}_{0}\times\mathbf{H}_{w_{i}}\times\mathbf{F}_{w_{i}}\to\mathbf{G}$
is injective, we have $\sigma(f)=f$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$ and hence $f\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. By Corollary 2.3, $w_{i}\cdot g$ is rational. ∎
Definition 2.5.
Let $g$ be a rational element in $\mathbf{G}(\mathbb{R})$. The denominator (or the height) $d(g)$ of $g$ is defined to be the co-volume of the lattice $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$.
In the following lemma, we discuss the relation between the shortest vector and the co-volume of a discrete subgroup of the form $\rho(g)\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}$.
Lemma 2.6.
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Suppose that $w_{i}\cdot g$ is rational. Then
$d(w_{i}\cdot g)\sim\delta(\rho(w_{i}\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})^{\dim V_{\beta_{0}}}.$
Here the implicit constant depends only on $\mathbf{G}$ and $\Gamma$.
Proof.
By Corollary 2.3, there exist $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$ such that $g=h_{i}\cdot f_{i}$. Note that
$f_{i}\in\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
and there exist $p\in\mathbf{P}_{0}(\mathbb{Q})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that $f_{i}=w_{i}^{-1}\cdot p\cdot x\cdot\gamma$. We know that $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the Levi factor of $\mathbf{P}_{0}$. We write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$. Then $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$, $\mathbf{T}$ commutes with $\mathbf{M}_{a}$ and $\mathbf{M}_{a}(\mathbb{R})/(\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma)$ is compact. So there exist $p_{1}\in\mathbf{T}$, $p_{2}$ in a compact fundamental domain of $\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ in $\mathbf{M}_{a}(\mathbb{R})$, $p_{3}\in\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ and $u\in R_{u}(\mathbf{P}_{0})$ such that
$p=u\cdot p_{1}\cdot p_{2}\cdot p_{3}.$
Then we have
$w_{i}\cdot g=(w_{i}\cdot h_{i}\cdot w_{i}^{-1})\cdot u\cdot p_{1}\cdot p_{2}\cdot p_{3}\cdot x\cdot\gamma.$
Since $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ and $R_{u}(\mathbf{P}_{0})$ fix every element in $V_{\beta_{0}}$, and $\mathbf{T}$ and $\mathbf{M}_{a}$ preserve the weight space $V_{\beta_{0}}$, we obtain
$\rho(w_{i}\cdot g)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2}\cdot p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2})(\rho(p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}).$
The lemma then follows from the facts that $p_{1}\in\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$, $p_{2}$ is in a fixed compact subset in $\mathbf{M}_{a}(\mathbb{R})$ and $\rho(p_{3}\cdot x)\mathbb{Z}^{d}$ is commensurable with $\mathbb{Z}^{d}$ by $\Gamma$ and $\mathcal{K}$. ∎
Now we consider the case that $w_{i}=e$ is the identity element in ${}_{\mathbb{Q}}W$. Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $\mathbf{F}_{w_{i}}=\mathbf{F}_{e}$ and $\mathbf{H}_{w_{i}}=\mathbf{H}_{e}$. By Lemma 2.4, there exist $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that
$g=h\cdot p\cdot x\cdot\gamma.$
Furthermore, we can write
$p=a\cdot m\cdot u$
where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$. Then we can compute the denominator $d(g)$ of $g$ as follows:
$(\rho(g)\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}=\rho(a\cdot m\cdot u\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(a\cdot m\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(a\cdot m)(\rho(x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})$
and
$d(g)=c_{x}\cdot e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}$
for some constant $c_{x}>0$ depending only on $x$. Here we use the facts that $\mathbf{M}_{a}$ and $R_{u}(\mathbf{P}_{0})$ stabilize $V_{\beta_{0}}$ and preserve the volumes of sets in $V_{\beta_{0}}$, $\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$ and $x\in\mathbf{G}(\mathbb{Q})$.
Definition 2.7.
Let $\mathcal{K}=\{x_{j}\}_{j\in J}$. A rational element $g$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is called $j$-rational for some $j\in J$ if it can be written as
$g=h\cdot p\cdot x_{j}\cdot\gamma$
for some $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x_{j}\in\mathcal{K}$ and $\gamma\in\Gamma$.
3. Counting rational points
In this section, we consider the problem of counting rational points in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. To estimate the Hausdorff dimension, one usually constructs boxes centered at certain rational points and obtains a tree-like subset. At each level of the tree-like subset, the boxes are disjoint. In our case, it may happen that the boxes at each level are not disjoint if we proceed in the usual way. It implies that the boxes constructed are too many, and we have to sieve out some proportion of the rational points so that the boxes centered at the remaining rational points are disjoint. It is also required that the number of the remaining rational points is not small compared to the total number of rational points in order to avoid any loss of Hausdorff dimension. For our purpose, in the following, we will first sieve some proportion of rational points according to several algebraic conditions, and then count the remaining rational points by the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$. The disjointness property of the boxes centered at the remaining rational points will follow from the transversal structure of some submanifolds in the homogeneous space $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
For any $\mathbb{Q}$-algebraic group $\mathbf{L}$ in $\mathbf{G}$ and $\delta>0$, we write $B_{\mathbf{L}}(\delta)$ for the open ball of radius $\delta>0$ around the identity in $\mathbf{L}(\mathbb{R})$. For any $t\in\mathbb{R}$, we write
$B_{\mathbf{L}}(\delta,t):=a_{-t}\cdot B_{\mathbf{L}}(\delta)\cdot a_{t}.$
The connected component of identity of $\mathbf{L}$ in the Zariski topology is denoted by $\mathbf{L}^{0}$, and the connected component of identity in the Lie group $\mathbf{L}(\mathbb{R})$ is denoted by $\mathbf{L}(\mathbb{R})^{0}$. We write $A\lesssim B$ $(A\gtrsim B)$ if there exists a constant $c>0$ such that
$A\leq c\cdot B\quad(A\geq c\cdot B).$
If $A\lesssim B$ and $A\gtrsim B$, then we write $A\sim B$. We will specify the implicit constants in the contexts if necessary.
In this section and next section, we assume that the action of the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. This holds true when the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into any $\mathbb{Q}$-simple factor of $\mathbf{G}$. Later in §5, we will explain how to establish the results of counting rational points and Hausdorff dimension estimates in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ appears only in some $\mathbb{Q}$-simple factors of $\mathbf{G}$.
Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. For any $0<A<B$, define
$S(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
We know that the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ has the following decomposition
$\mathfrak{a}=\operatorname{Lie}(a_{t})\oplus\ker(\beta_{0})$
where $\operatorname{Lie}(a_{t})$ is the Lie algebra of the one parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ and we have
$\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0.$
Denote by $\pi_{\ker(\beta_{0})}$ the projection from $\mathfrak{a}$ to $\ker(\beta_{0})$ according to this decomposition. We also write $\pi_{\ker(\beta_{0})}(a)$ for $\pi_{\ker(\beta_{0})}(\log(a))$ whenever $a\in\exp(\mathfrak{a})$. Here $\log$ is the inverse of the exponential map $\exp$. For any compact subset $L$ in $\operatorname{Lie}(a_{t})$ and any compact subset $K$ in $\ker(\beta_{0})$, we denote by
$\mathfrak{a}_{L,K}=\{x\in\mathfrak{a}:x=y_{1}+y_{2},y_{1}\in L,y_{2}\in K\}.$
For any $x_{j}\in\mathcal{K}$, any compact subset $K_{1}$ in $\ker(\beta_{0})$, any compact subset $K_{2}$ in $\mathbf{H}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$, we define
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$
to be the set of all rational points $q$ in $U$ such that
  1. (1)
    $A\leq d(q)\leq B$ and $q$ is $j$-rational for $x_{j}\in\mathcal{K}$;
  2. (2)
    $q=a\cdot h\cdot m\cdot u\cdot x_{j}\cdot\gamma$ for some $a\in\exp(\mathfrak{a})$, $\pi_{\ker(\beta_{0})}(a)\in K_{1}$ and $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\Gamma$.
Note that $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\subset S(U,A,B)$, and by definition one can check that if
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\neq\emptyset,$
then $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$ (and such elements exist as $e\in\mathcal{K}$). The elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ are the rational points we are interested in when we construct a tree-like subset in §4. We denote by
$S_{K_{1},K_{2},K_{3},K_{4}}(U,A,B):=\bigcup_{j\in J}S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B).$
To count the rational points in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$, we need results about limiting distributions of translates of unipotent orbits pushed by the one-parameter semisimple flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. The following is a direct consequence of the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
Proposition 3.1.
Let $x\in\mathbf{G}(\mathbb{R})^{0}/\Gamma$ and $W\subset\mathbf{G}(\mathbb{R})^{0}/\Gamma$ an open bounded subset whose boundary has measure zero with respect to the invariant probability measure $\mu_{\mathbf{G}^{0}(\mathbb{R})/\Gamma}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Let $\chi_{W}$ denote the characteristic function of $W$. Then for any bounded open subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})$ we have
[…]
where $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}$ is the Haar measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Remark 3.2.
Note that $R_{u}(\overline{\mathbf{P}}_{0})$ is not necessarily the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$. Here we can still apply the mixing property to obtain Proposition 3.1 as long as $R_{u}(\overline{\mathbf{P}}_{0})$ is contained in the group generated by the unstable subgroup and the central subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$. (See [25, §2].)
Now we can follow the same arguments as in [17, 48] and measure the size of the subset $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ for a fixed index $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$, and for some sufficiently large numbers $B>A>0$. Recall that
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{e}\cdot\mathbf{F}_{e}$
where $\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\cap\mathbf{P}_{\beta_{0}}$. We fix a Haar measure $\mu_{\mathbf{H}_{e}}$ on $\mathbf{H}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\mathbf{F}_{e}}$ on $\mathbf{F}_{e}(\mathbb{R})$ such that the product maps
$\mathbf{F}_{e}\times\mathbf{H}_{e}\to\mathbf{F}_{e}\cdot\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\textup{ and }\mathbf{H}_{e}\times\mathbf{F}_{e}\to\mathbf{H}_{e}\cdot\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{0})$
induce Haar measures on $R_{u}(\overline{\mathbf{P}}_{0}).$ Then for any $q\in\mathbf{F}_{e}(\mathbb{Q})$, we define $m_{\mathbf{H}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which is supported on $\mathbf{H}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\mathbf{H}_{e}}$ via the product map
$\mathbf{H}_{e}\times\{q\}\to\mathbf{H}_{e}\cdot q\subset R_{u}(\overline{\mathbf{P}}_{0}).$
We define
$m_{\mathbf{H}_{e}}:=\sum_{q\in\mathbf{F}_{e}(\mathbb{Q})}m_{\mathbf{H}_{e}q}.$
Note that $m_{\mathbf{H}_{e}}$ is not a locally finite measure on $R_{u}(\overline{\mathbf{P}}_{0})$, and it is defined by the leaf-wise measures on the countable leaves through the rational elements in $\mathbf{F}_{e}(\mathbb{Q})$ in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ induced by the group action of $\mathbf{H}_{e}$ on $R_{u}(\overline{\mathbf{P}}_{0})$. Note also that by Corollary 2.3, we have
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\subset\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\mathbf{H}_{e}(\mathbb{R})\cdot q=\mathbf{H}_{e}(\mathbb{R})\cdot\mathbf{F}_{e}(\mathbb{Q}).$
In the following, we estimate the size of the subset
$S_{K_{1},K_{2},K_{3},K_{4},j}(U,(l/2)^{\dim V_{\beta_{0}}},l^{\dim V_{\beta_{0}}})$
with respect to the measure $m_{\mathbf{H}_{e}}$ for sufficiently large $l>0$. For convenience, we write
$A_{l}=(l/2)^{\dim V_{\beta_{0}}}\textup{ and }B_{l}=l^{\dim V_{\beta_{0}}}\quad(\forall l>0).$
For any $l>1$, let $T=T(l)>0$ such that
$\beta_{0}(a_{T})=-\ln l.$
Let $q$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, we may write
$q=a\cdot h\cdot m\cdot u\cdot x_{k}\cdot\gamma\in\mathbf{H}_{e}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot x_{k}\cdot\Gamma$
for some $a\in\mathbf{T}(\mathbb{R})$, $h\in\mathbf{H}_{e}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$, $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, $x_{k}\in C$ and $\gamma\in\Gamma$. Then
$q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
$\iff$ $q\in U,\;A_{l}\leq d(q)\leq B_{l},\textrm{ and }a\in\exp(\mathfrak{a}),\pi_{\ker(\beta_{0})}(a)\in K_{1},h\in K_{2},m\in K_{3},u\in K_{4},k=j.$
$\iff$ $a_{T}\cdot q\Gamma\in a_{T}\cdot U\Gamma/\Gamma\textrm{ and }a_{T}\cdot q\Gamma\in\exp(\mathfrak{a}_{I_{0},K_{1}})\cdot K_{2}\cdot K_{3}\cdot K_{4}\cdot x_{j}\cdot\Gamma/\Gamma$
where $I_{0}$ is the following compact interval in the Lie algebra of $\left\{a_{t}\right\}_{t\in\mathbb{R}}$
$I_{0}:=\left\{x\in\mathbb{\operatorname{Lie}}(a_{t}):-\ln(2/{c_{x_{j}}^{\frac{1}{\dim V_{\beta_{0}}}}})\leq\beta_{0}(x)\leq-\ln(1/{c_{x_{j}}^{\frac{1}{\dim V_{\beta_{0}}}}})\right\}.$
Here we use the formula
$d(q)=c_{x_{j}}e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}.$
This implies that
$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
Since $\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$ is a compact subset in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$, there exist $\delta_{0}>0$ and a small neighborhood of identity $B_{\mathbf{F}_{e}}(\delta_{0})$ in $\mathbf{F}_{e}(\mathbb{R})$ such that
$B_{\mathbf{F}_{e}}(\delta_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
is a homeomorphism. We conclude that the following product map
$B_{\mathbf{F}_{e}}(\delta_{0})\times a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$
gives a transversal structure near the set $a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$ in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Consequently, for any $p,q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, the subsets $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot p\Gamma$ and $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot q\Gamma$ are disjoint, where
$B_{\mathbf{F}_{e}}(\delta_{0},T):=a_{-T}\cdot B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}.$
Now we estimate the size of $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ with respect to the measure $m_{\mathbf{H}_{e}}$. First we prove an upper bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$. Fix a sufficiently small number $0<\epsilon<\delta_{0}$ such that
$\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\leq\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\leq 2\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
Then for sufficiently large $l>0$, we have
$B_{\mathbf{F}_{e}}(\epsilon,T)\subset B_{\mathbf{F}_{e}}(\epsilon)\textup{ and}B_{\mathbf{F}_{e}}(\epsilon,T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\subset B_{\mathbf{F}_{e}}(\epsilon)\cdot U\Gamma/\Gamma.$
Since $B_{\mathbf{F}_{e}}(\delta_{0},T)$ and $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ form a transversal structure near $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, by Proposition 3.1, we deduce that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))$
[…]
$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\cdot\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}(B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma)$
$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))$
as $T\to\infty$. Here the implicit constant in the last equation depends only on the parameter $\delta_{0}>0$ (as $0<\epsilon<\delta_{0}$), and hence depends only on the compact subsets $K_{i}$’s. Note that
$\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))=$ $\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))\cdot e^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{T})}$
$=$ $\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))\cdot l^{\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{T})/\beta_{0}(a_{T})}$
$=$ $\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))\cdot l^{\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}.$
So for sufficiently large $l>0$ we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\lesssim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
where the implicit constant depends only on $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $I_{0}$, $\mathbf{G}$ and $\Gamma$. Similarly, using the arguments in [17, §4], we can prove a lower bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$ and for any sufficiently large $l>0$ we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\gtrsim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
To sum up, we have the following
Proposition 3.3.
Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})$. Then for any $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
Here the implicit constants depend only on the compact subsets $K_{i}$’s, $\mathbf{G}$ and $\Gamma$.
Remark 3.4.
For our purpose, in the following, we fix some compact subsets $K_{1}\subset\ker\beta_{0}$, $K_{2}\subset\mathbf{H}_{e}(\mathbb{R})$, $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$ discussed above.
4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
In this section, we prove Theorem 1.3 under the assumption that the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
Lemma 4.1.
Fix $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and
$\beta_{0}(a_{T})=-\ln l$
for some $T>0$ and $l>1$. Let $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\mathbf{F}_{e}(\mathbb{Q})$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$
Then there exist $\theta_{1}>0$ and $\theta_{2}>0$ such that for any $p\in\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
$B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p\cap U\subset\mathcal{F}_{q}\cap S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ and $B_{\mathbf{H}_{e}}(\theta_{1},T)=a_{-T}\cdot B_{\mathbf{H}_{e}}(\theta_{1})\cdot a_{T}$. Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\mathbf{G}$ and $\Gamma$.
Proof.
By the discussion in §3, we know that
$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
Choose $\theta_{1},\theta_{2}>0$ sufficiently small so that for any $a\in\exp(\mathfrak{a}_{I_{0},K_{1}})$ we have
$a^{-1}B_{\mathbf{H}_{e}}(\theta_{1})a\subset B_{\mathbf{H}_{e}}(\theta_{2}).$
Now for any $p\in\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, there exist $a\in\exp(\mathfrak{a}_{I_{0},K_{1}})$, $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\Gamma$ such that
$a_{T}\cdot p=a\cdot h\cdot m\cdot u\cdot x_{j}\cdot\gamma.$
Then
$a_{T}\cdot B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p=B_{\mathbf{H}_{e}}(\theta_{1})\cdot a\cdot h\cdot m\cdot u\cdot\gamma\subset\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. By definition, we have
$B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p\cap U\subset S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l}).$
This completes the proof of the lemma. ∎
Lemma 4.2.
We have
$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}>0.$
Proof.
Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{e})$, $\alpha(a_{1})=0$. Let $\Omega_{\mathbf{F}_{e}}$ and $\Omega_{\mathbf{H}_{e}}$ be small open neighborhoods of identity in $\mathbf{F}_{e}(\mathbb{R})$ and $\mathbf{H}_{e}(\mathbb{R})$ respectively, and
$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{e}}\cdot\Omega_{\mathbf{H}_{e}}.$
Note that $\mathbf{H}_{e}(\mathbb{R})$ stabilizes every element in $V_{\beta_{0}}$. Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{e}}$ and $h\in\Omega_{\mathbf{H}_{e}}$, we have
$a_{t}\cdot p=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}=f\cdot(a_{t}ha_{-t})\cdot a_{t}.$
The element $a_{t}$ is rational and
$\delta(\rho(a_{t})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(a_{t})}.$
This implies that
$\delta(\rho(a_{t}\cdot p)\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(a_{-1})t}$
for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{e}}$. Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$. Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. This contradicts the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. ∎
Let $X$ be a Riemannian manifold, $m$ a volume form on $X$ and $E$ a compact subset of $X$. Denote by $\operatorname{diam}(S)$ the diameter of a set $S\subset X$. A collection $\mathcal{A}$ of compact subsets of $E$ is said to be tree-like if $\mathcal{A}$ is the union of finite sub-collections $\mathcal{A}_{k}$ such that
  1. (1)
    $\mathcal{A}_{0}=\left\{E\right\}$;
  2. (2)
    For any $k$ and $S_{1},S_{2}\in\mathcal{A}_{j}$, either $S_{1}=S_{2}$ or $S_{1}\cap S_{2}=\emptyset$;
  3. (3)
    For any $k$ and $S_{1}\in\mathcal{A}_{k+1}$, there exists $S_{2}\in\mathcal{A}_{k}$ such that $S_{1}\subset S_{2}$;
  4. (4)
    $d_{k}(\mathcal{A}):=\sup_{S\in\mathcal{A}_{k}}\operatorname{diam}(S)\to 0$ as $k\to\infty$.
We write $\mathbf{A}_{k}:=\bigcup_{A\in\mathcal{A}_{k}}A$ and $\mathbf{A}_{\infty}:=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. We also define
$\Delta_{k}(\mathcal{A}):=\inf_{S\in\mathcal{A}_{k}}\frac{m(\mathbf{A}_{k+1}\cap S)}{m(S)}.$
Theorem 4.3 ([25, 31, 44]).
Let $(X,m)$ be a Riemannian manifold, where $m$ is the volume form on $X$. Then for any tree-like collection $\mathcal{A}$ of subsets of $E$
$\dim_{H}(\mathbf{A}_{\infty})\geq\dim_{H}X-\limsup_{k\to\infty}\frac{\sum_{i=0}^{k}\log(\Delta_{i}(\mathcal{A}))}{\log(d_{k+1}(\mathcal{A}))}.$
Proof of Theorem 1.3.
We fix the compact subsets $K_{1}\subset\ker(\beta_{0}),K_{2}\subset\mathbf{H}_{e}(\mathbb{R}),K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0},K_{4}\subset R_{u}(\mathbf{P}_{0})$ as in Remark 3.4, and also fix $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$. Let $\epsilon>0$ be a sufficiently small number. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
Then by Lemma 4.2, $\nu_{0}(a_{1})>0$.
We start with a small open bounded box $U_{0}$ in $R_{u}(\overline{\mathbf{P}}_{0})$.
For $k=0$, we set $\mathcal{A}_{0}=\left\{U_{0}\right\}$.
For $k=1$, we choose sufficiently large numbers $l_{1}>0$ and $T_{1}>0$ such that
$\beta_{0}(a_{T_{1}})=-\ln l_{1}.$
By Proposition 3.3, we know that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for some $\theta_{2}>0$ as defined in Lemma 4.1. Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset,$
we devide the region $\mathcal{F}_{q}\cap U_{0}$ into small cubes of side length
$\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes $R$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$, and denote the corresponding collection by
$\mathcal{G}_{1,q}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset\}.$
Note that $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$ is smaller than the minimum side length of the rectangle $B_{\mathbf{H}_{e}}(\theta_{1},T_{1})$, so by Lemma 4.1, we know that each cube in $\mathcal{G}_{1,q}$ is contained in
$S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
$\mathcal{H}_{1}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R.$
Then we have
$l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R)$
$\leq m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
$\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
and
$m_{\mathbf{H}_{e}}(\mathcal{H}_{1})\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0}).$
Note that each cube $R$ in $\mathcal{H}_{1}$ is contained in $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$. Let
$t_{1}=\frac{\ln l_{1}}{\beta_{0}(a_{-1})-(\tau+\epsilon)}.$
By the computations in §3, we can choose a sufficiently small number $\tilde{\delta}_{0}>0$ such that
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
is a homeomorphism, and the subsets in the collection
$\mathcal{F}_{1}(U_{0}):=\left\{\left(a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\right)\cdot q:q\in\mathcal{H}_{1}\right\}$
are disjoint by this homeomorphism as
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{1}).$
We write
$\Phi(\mathbf{F}_{e})=\Phi^{0}(\mathbf{F}_{e})\cup\Phi^{1}(\mathbf{F}_{e})$
where
$\Phi^{0}(\mathbf{F}_{e})=\{\alpha\in\Phi(\mathbf{F}_{e}):\alpha(a_{1})=0\}\textup{ and }\Phi^{1}(\mathbf{F}_{e})=\{\alpha\in\Phi(\mathbf{F}_{e}):\alpha(a_{1})\neq 0\}.$
We denote by
$\mathcal{P}_{1}=\bigcup_{E\in\mathcal{F}_{1}(U_{0})}E.$
Now we can divide the subset $\mathcal{P}_{1}$ into cubes of side length
$\tilde{\delta}_{0}\cdot l_{1}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
which is smaller than $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$ and the minimum side length of the rectangle
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}$
if $T_{1}>0$ and $l_{1}>0$ are sufficiently large. Note that the set $\Phi^{0}(\mathbf{F}_{e})$ may not be empty and the diameter of the set
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}$
may be larger than the diameter of $U_{0}$, so some cubes we obtain here from dividing the subset $\mathcal{P}_{1}$ may be outside the set $U_{0}$. For our purpose, we collect only those cubes which are inside the subset $U_{0}$. In this manner, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{1}(U_{0})$ inside $U_{0}$, which we denote by $\mathcal{A}_{1}$.
We remark here that if $\Phi^{0}(\mathbf{F}_{e})=\emptyset$, then all the cubes we obtain from dividing $\mathcal{P}_{1}$ are inside the subset $U_{0}$ if $T_{1}>0$ and $l_{1}>0$ are chosen to be sufficiently large (and also if we shrink or enlarge $U_{0}$ slightly to avoid the complexity caused by the boundary of $U_{0}$). We will see later that in the computations there is little difference between the case $\Phi^{0}(\mathbf{F}_{e})\neq\emptyset$ and the case $\Phi^{0}(\mathbf{F}_{e})=\emptyset$. Indeed, when we apply Theorem 4.3 and calculate $\Delta_{i}(\mathcal{A})$ and $d_{i}(\mathcal{A})$, the difference between these two cases may affect the value of the formula in finite steps, but eventually when we take the limit, this difference will disappear since we keep choosing sufficiently large numbers $T_{1},l_{1}$ and later $T_{2},l_{2},T_{3},l_{3},\dots$ to offset the effects by the difference at early stages.
Similarly, we can construct $\mathcal{A}_{k}$ inductively for any $k\in\mathbb{N}$. For $k>1$ we choose sufficiently large numbers $l_{k}>0$ and $T_{k}>0$ such that $\beta_{0}(a_{\tau_{k}})=-\ln l_{k}$. For each cube $S\in\mathcal{A}_{k-1}$, by the proposition above, we know that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for $\theta_{2}>0$ as defined in Lemma 4.1. Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset,$
we devide the region $\mathcal{F}_{q}\cap S$ into small cubes of side length
$\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$, and denote the corresponding collection by
$\mathcal{G}_{k,q,S}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset\}.$
Note that $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$ is smaller than the minimum side length of the rectangle $B_{\mathbf{H}_{e}}(\theta_{1},T_{k})$, so by Lemma 4.1, we know that each cube in $\mathcal{G}_{k,q,S}$ is contained in
$S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
$\mathcal{H}_{k,S}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R.$
Then we have
$l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R)$
$\leq m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
$\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
and
$m_{\mathbf{H}_{e}}(\mathcal{H}_{k,S})\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S).$
Note that each cube in $\mathcal{H}_{k,S}$ is contained in $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$. Let
$t_{k}=\frac{\ln l_{k}}{\beta_{0}(a_{-1})-(\tau+\epsilon)}.$
By the computations in §3, we can choose a sufficiently small number $\tilde{\delta}_{0}>0$ such that
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
is a homeomorphism, and the subsets in the collection
$\mathcal{F}_{k}(S):=\left\{(a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}})\cdot q:q\in\mathcal{H}_{k,S}\right\}$
are disjoint by this homeomorphism as
$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{k}).$
We denote by
$\mathcal{P}_{k,S}=\bigcup_{E\in\mathcal{F}_{k}(S)}E.$
Now we can divide the subset $\mathcal{P}_{k,S}$ into cubes of side length
$\tilde{\delta}_{0}\cdot l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
which is smaller than $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$ and the minimum side length of the rectangle
$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}$
if $T_{k}>0$ and $l_{k}>0$ are sufficiently large. Note that $\Phi^{0}(\mathbf{F}_{e})$ may not be empty, and some of these cubes may be outside $S$, and here we collect only those cubes which are inside $S$. Thus, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{k}(S)$ inside $S$. We denote by $\mathcal{A}_{k}$ the collection of all these disjoint cubes constructed from $\mathcal{F}_{k}(S)$ inside $S$ where $S$ ranges over all elements in $\mathcal{A}_{k-1}$.
In this manner, we obtain a sequence $\left\{l_{k}\right\}_{k\in\mathbb{N}}$ of sufficiently large numbers $l_{k}$ with
$l_{k+1}\gg l_{k}\;(\forall k\in\mathbb{N})$
and a tree-like structure $\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ of finite collections of cubes. Using the notation in Theorem 4.3, we have
$d_{k}(\mathcal{A})\sim l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
where $d_{k}(\mathcal{A})$ is the diameter of the family $\mathcal{A}_{k}$. Moreover, one can compute that
$\begin{cases}\Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})=\emptyset\\ \Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}\cdot\prod_{\alpha\in\Phi^{0}(\mathbf{F}_{e})}l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})\neq\emptyset.\end{cases}$
Now let $\mathbf{A}_{\infty}=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. By applying Theorem 4.3, we can compute (assuming that $l_{k+1}$ is much larger than $l_{k}$ for any $k\in\mathbb{N}$) that
$\dim_{H}(\mathbf{A}_{\infty})\geq$ $\dim_{H}X-\limsup_{k\to\infty}\frac{\sum_{i=0}^{k}\log(\Delta_{i}(\mathcal{A}))}{\log(d_{k+1}(\mathcal{A}))}$
$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
Lemma 4.4.
We have $\mathbf{A}_{\infty}\subset S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Proof.
By the construction, for any $p\in\mathbf{A}_{\infty}$, there exists a sequence of rational points
$q_{k}\in S_{K_{1},K_{2},K_{3},K_{4},j}(S_{k-1},A_{l_{k}},B_{l_{k}}),\quad S_{k-1}\in\mathcal{A}_{k-1}$
such that
$p\in(a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}})\cdot q_{k}$
where $t_{k}=\frac{\ln l_{k}}{\beta_{0}(a_{-1})-(\tau+\epsilon)}$. Then
$a_{t_{k}}\cdot p\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot(a_{t_{k}}\cdot q_{k}).$
Note that by definition, the denominator of the rational element $a_{t_{k}}\cdot q_{k}$ is equal to the co-volume of the lattice $a_{t_{k}}\cdot q_{k}\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$ and
$d(a_{t_{k}}\cdot q_{k})=e^{\beta_{0}(a_{t_{k}})\cdot\dim V_{\beta_{0}}}\cdot d(q_{k}).$
By Lemma 2.6, we deduce that
$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\sim\delta(\rho(a_{t_{k}}\cdot q_{k})\mathbb{Z}^{d})\lesssim d(a_{t_{k}}\cdot q_{k})^{\frac{1}{\dim V_{\beta_{0}}}}$
$\lesssim e^{\beta_{0}(a_{t_{k}})}\cdot l_{k}\leq e^{-(\tau+\epsilon)t_{k}}$
and $p\in S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. This completes the proof of the lemma. ∎
By Lemma 4.4 and the computation for $\dim_{H}\mathbf{A}_{\infty}$, we have
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
By taking $\epsilon\to 0$, we obtain
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Using the same argument as in [17, Section 10], we conclude that
$\dim_{H}(S_{\tau}^{c})\geq$ $\dim\mathbf{G}(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Note that by Lemma 4.2, we have $\nu_{0}(a_{1})>0$ and the argument here works for all $0\leq\tau<\beta_{0}(a_{-1})$. This completes the proof of Theorem 1.3 (provided that $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$). ∎
5. Proof of Theorem 1.3: $\{a_{t}\}_{t\in\mathbb{R}}$ in a proper normal $\mathbb{Q}$-subgroup of $\mathbf{G}$
In this section, we discuss the case when the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, and explain how to modify the arguments in §3 and §4 and give a proof of Theorem 1.3.
We first discuss the ergodic properties of group actions on homogeneous spaces, and one can refer to e.g. [29, 35, 37] for details. Let $\mathbf{G}_{i}$ $(1\leq i\leq k)$ be the $\mathbb{Q}$-simple factors of $\mathbf{G}$. Then $\mathbf{G}$ is an almost direct product of $\mathbf{G}_{i}$ $(1\leq i\leq k)$. Without loss of generality, we may assume that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ ($1\leq i\leq s$) for some $s<k$. It is known that for each $1\leq i\leq k$, any arithmetic lattice $\Gamma_{i}$ inside $\mathbf{G}_{i}(\mathbb{Z})\cap\mathbf{G}_{i}(\mathbb{R})^{0}$ is irreducible in $\mathbf{G}_{i}(\mathbb{R})^{0}$ (and we fix such an arithmetic lattice $\Gamma_{i}$ for later use). Moreover, since $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$, we have $\mathbf{T}\cap\mathbf{G}_{i}\neq\{e\}$ $(1\leq i\leq s)$. So $\mathbf{G}_{i}$ is $\mathbf{Q}$-isotropic, and $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is not compact. By Godement compactness criterion (Cf. [7, Theorem 11.6]) and [8, 6.21], every simple factor of $\mathbf{G}_{i}(\mathbb{R})^{0}$ is not compact. Let $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ be the projection of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$. Then $\{a_{t}^{i}\}_{t\in\mathbb{R}}\subset\mathbf{T}(\mathbb{R})\cap\mathbf{G}_{i}(\mathbb{R})$ is a non-compact subgroup in $\mathbf{G}_{i}(\mathbb{R})^{0}$. Using Moore’s ergodicity theorem [33] and Mautner phenomenon [30, 34], one can conclude that the action of $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ on $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is mixing $(1\leq i\leq s)$. Consequently, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\prod_{i=1}^{s}\mathbf{G}_{i}(\mathbb{R})^{0}/\prod_{i=1}^{s}\Gamma_{i}$ is mixing.
We denote by
$\tilde{\mathbf{G}}=\prod_{i=1}^{s}\mathbf{G}_{i}=\mathbf{G}_{1}\cdot\mathbf{G}_{2}\cdot\cdots\cdot\mathbf{G}_{s}.$
Note that in §2, §3 and §4, we don’t use any explicit expression of the lattice $\Gamma$ in $\mathbf{G}(\mathbb{R})^{0}$. So we may choose $\Gamma$ such that
$\Gamma\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}=\prod_{i=1}^{s}\Gamma_{i}.$
In the follwing, we fix these lattices $\Gamma_{i}$’s and $\Gamma$ and denote by
$\tilde{\Gamma}=\Gamma\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}.$
Therefore, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ is mixing.
Now all the arguments in §2, §3 and §4 can be carried over to $\tilde{\mathbf{G}}$ and the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ almost verbatim. Indeed, denote by
$\tilde{\mathbf{P}}_{0}=\mathbf{P}_{0}\cap\tilde{\mathbf{G}}$
which is a minimal parabolic $\mathbb{Q}$-subgroup of $\tilde{\mathbf{G}}$, and let $\overline{\tilde{\mathbf{P}}}_{0}=\overline{\mathbf{P}}_{0}\cap\tilde{\mathbf{G}}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and define
$\tilde{S}(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
Let $\tilde{\mathbf{T}}=\mathbf{T}\cap\tilde{\mathbf{G}}$ which is a maximal $\mathbb{Q}$-split torus in $\tilde{\mathbf{G}}$, and
$\tilde{\mathbf{H}}_{e}=\mathbf{H}_{e}\cap\tilde{\mathbf{G}},\quad\tilde{\mathbf{F}}_{e}=\mathbf{F}_{e}\cap\tilde{\mathbf{G}}.$
In the following, we need results in the reduction theory about $\tilde{\mathbf{G}}(\mathbb{R})/\tilde{\Gamma}$. Let $\tilde{K}$ be a maximal compact subgroup in $\tilde{\mathbf{G}}(\mathbb{R})$. Let $\tilde{M}$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\tilde{\mathbf{G}}(\mathbb{R})}(\tilde{\mathbf{T}}(\mathbb{R}))$ (= the centralizer of $\tilde{\mathbf{T}}(\mathbb{R})$ in $\tilde{\mathbf{G}}(\mathbb{R})$). Denote by
$\tilde{T}_{\eta}=\{a\in\tilde{\mathbf{T}}(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\tilde{\Delta}\}$
where $\tilde{\Delta}$ is the set of positive $\mathbb{Q}$-simple roots in $\tilde{\mathbf{G}}$. A Siegel set in $\tilde{\mathbf{G}}(\mathbb{R})$ is a subset of the form $\tilde{S}_{\eta,\Omega}=\tilde{K}\cdot\tilde{T}_{\eta}\cdot\tilde{\Omega}$ for some $\eta\in\mathbb{R}$ and a relatively compact open subset $\tilde{\Omega}$ containing identity in $\tilde{M}\cdot R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, and the group $\tilde{\mathbf{G}}(\mathbb{R})$ can be written as
$\tilde{\mathbf{G}}(\mathbb{R})=\tilde{S}_{\eta,\Omega}\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}$
for some Siegel set $\tilde{S}_{\eta,\Omega}$ and some finite subset $\tilde{\mathcal{K}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$. Moreover, the finite set $\tilde{\mathcal{K}}$ satisfies the property that
$\tilde{\mathbf{G}}(\mathbb{Q})=\tilde{\mathbf{P}}_{0}(\mathbb{Q})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}.$
Denote by $\tilde{\mathcal{K}}=\{\tilde{x}_{j}\}_{j\in\tilde{J}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$ and we may assume that $e\in\tilde{\mathcal{K}}$.
Lemma 5.1.
We have
  1. (1)
    An element $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is rational if and only if $g\in\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot\tilde{\mathbf{F}}_{e}(\mathbb{Q})$.
  2. (2)
    An element $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is rational if and only if
    $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})\cap(\tilde{\mathbf{H}}_{1}(\mathbb{R})\cdot\tilde{\mathbf{P}}_{0}(\mathbb{R})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}).$
Definition 5.2.
Let $j\in\tilde{J}$. A rational element $g$ in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is called $j$-rational if it can be written as
$g=h\cdot p\cdot\tilde{x}_{j}\cdot\gamma$
for some $h\in\tilde{\mathbf{H}}_{e}(\mathbb{R})$, $p\in\tilde{\mathbf{P}}_{0}(\mathbb{R})$ and $\gamma\in\tilde{\Gamma}$.
Let $\tilde{\mathfrak{a}}$ be the Lie algebra of $\tilde{\mathbf{T}}(\mathbb{R})$. For any $\tilde{x}_{j}\in\tilde{\mathcal{K}}$, any compact subset $K_{1}$ in $\ker(\beta_{0})\cap\tilde{\mathfrak{a}}$, any compact subset $K_{2}$ in $\tilde{\mathbf{H}}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, we define
$\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$
to be the set of all rational points $q$ in $U$ such that
  1. (1)
    $A\leq d(q)\leq B$ and $q$ is $j$-rational for $\tilde{x}_{j}\in\tilde{C}$;
  2. (2)
    $q=a\cdot h\cdot m\cdot u\cdot\tilde{x}_{j}\cdot\gamma$ for some $a\in\exp(\tilde{\mathfrak{a}})$, $\pi_{\ker(\beta_{0})}(a)\in K_{1}$, $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\tilde{\Gamma}$.
Note that $\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\neq\emptyset$ implies that $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$.
We fix a Haar measure $\mu_{\tilde{\mathbf{H}}_{e}}$ on $\tilde{\mathbf{H}}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\tilde{\mathbf{F}}_{e}}$ on $\tilde{\mathbf{F}}_{e}(\mathbb{R})$. Then for any $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$, we define $m_{\tilde{\mathbf{H}}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ which is supported on $\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\tilde{\mathbf{H}_{e}}}$ via the product map
$\tilde{\mathbf{H}}_{e}\times\{q\}\to\tilde{\mathbf{H}}_{e}\cdot q\subset R_{u}(\overline{\tilde{\mathbf{P}}}_{0}).$
We define
$m_{\tilde{\mathbf{H}}_{e}}:=\sum_{q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})}m_{\tilde{\mathbf{H}}_{e}q}.$
Then using the same argument as in Proposition 3.3, we have
Proposition 5.3.
Let $U$ be a small open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$. Then for any $j\in\tilde{J}$ with $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U)$
and
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U).$
Here the implicit constants depend only on the compact subsets $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
The analogues of Lemma 4.1 and Lemma 4.2 hold as well in $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
Lemma 5.4.
Fix $j\in\tilde{J}$ with $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and
$\beta_{0}(a_{T})=-\ln l$
for some $T>0$ and $l>1$. Let ~$\tilde{\mathcal{F}}_{q}=\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$ such that
~$\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$
Then there exist $\theta_{1}>0$ and $\theta_{2}>0$ such that for any $p\in\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
$B_{\tilde{\mathbf{H}}_{e}}(\theta_{1},T)\cdot p\cap U\subset\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
where $\tilde{K}_{2}=B_{\tilde{\mathbf{H}}_{1}}(\theta_{2})\cdot K_{2}$ and $B_{\tilde{\mathbf{H}}_{1}}(\theta_{1},T)=a_{-T}\cdot B_{\tilde{\mathbf{H}}_{e}}(\theta_{1})\cdot a_{T}$. Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
Lemma 5.5.
We have
$\max\{\alpha(a_{1}):\alpha\in\Phi(\tilde{\mathbf{F}}_{e})\}>0.$
Consequently, using the same argument as in §4, we can conclude that
$\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))\geq\dim\tilde{\mathbf{G}}-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\tilde{\nu}_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
where $\tilde{\nu}_{0}$ is the $\mathbb{Q}$-root in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$ such that
$\tilde{\nu}_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\tilde{\mathbf{P}}}_{0}))\}.$
Note that $\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$ all commute with $\{a_{t}\}_{t\in\mathbb{R}}$, and an element $g\in S_{\tau}^{c}$ if and only if $h\cdot g\in S_{\tau}^{c}$ for any $h\in\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$. Therefore, we have
$\dim_{H}(S_{\tau}^{c})=$ $\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))+\sum_{i=s+1}^{k}\dim\mathbf{G}_{i}(\mathbb{R})$
$\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
where $\nu_{0}(a_{1}):=\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}$. Here we use the fact that $\nu_{0}(a_{1})=\tilde{\nu}_{0}(a_{1})$ and
$\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})=\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1}).$
This completes the proof of Theorem 1.3.
Remark 5.6.
One can see from the arguments in §4 and §5 that we actually prove that for any open bounded subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$
$\dim_{H}(S_{\gamma}^{c}\cap U)\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Note that $\mathbf{P}_{0}\cdot R_{u}(\overline{\mathbf{P}}_{0})$ is a Zariski open dense subset in $\mathbf{G}$, and if $x\in S_{\gamma}^{c}\cap U$, then for any $g\in\mathbf{P}_{0}(\mathbb{R})$, $g\cdot x$ is also an element in $S_{\gamma}^{c}$. This implies that for any open bounded subset $\tilde{U}\subset\mathbf{G}(\mathbb{R})$
$\dim_{H}(S_{\gamma}^{c}\cap\tilde{U})\geq\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
In this section, we prove Theorem 1.4. We will compute an upper bound of $\dim_{H}S_{\gamma}^{c}$ by constructing open covers of the subset $S_{\gamma}^{c}$. Recall that
$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
Proposition 6.1.
Suppose that
$\delta(\rho(g)\cdot\mathbb{Z}^{d})\leq r$
for some $g\in\mathbf{G}(\mathbb{R})$ and $r>0$. Then there exist $\tilde{g}\in\mathcal{C}_{\bar{I}}$ and a discrete primitive subgroup
$\Lambda_{g}\subset\rho(g)\cdot\mathbb{Z}^{d}$
such that
$\operatorname{Vol}(\Lambda_{g})\leq C\cdot r^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}\textup{ is Zariski densein }V_{\beta_{0}}$
where the constant $C>0$ depends only on $\mathbf{G}$, $\Gamma$ and $\rho$.
Proof.
By the reduction theory of arithmetic subgroups, $\mathbf{G}(\mathbb{R})$ can be written as
$\mathbf{G}(\mathbb{R})=S_{\eta,\Omega}\cdot\mathcal{K}\cdot\Gamma$
for some Siegel set $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ where
$T_{\eta}=\{a\in T(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\Delta\}$
and the finite set $\mathcal{K}$ satisfies the property that
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma.$
Then there exist $k\in K$, $a\in T_{\eta}$, $u\in\Omega$, $\gamma\in\Gamma$ and $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ such that
$g=k\cdot a\cdot u\cdot x\cdot\gamma.$
Note that by the definition of $T_{\eta}$, the subset
$\{aua^{-1}:a\in T_{\eta}\textup{ and }u\in\Omega\}$
is compact. So we can write
$g=\tilde{k}\cdot a\cdot x\cdot\gamma$
for some $\tilde{k}$ in a fixed compact subset in $\mathbf{G}(\mathbb{R})$.
Now since $\delta(\rho(g)\cdot\mathbb{Z}^{d})\leq r$, there exists $y\in\mathbb{Z}^{d}\setminus\{0\}$ such that $\rho(\tilde{k}\cdot a\cdot x)\cdot y$ is a shortest vector in $\rho(g)\cdot\mathbb{Z}^{d}$ and
$\|\rho(a\cdot x)\cdot y\|\sim\|\rho(\tilde{k}\cdot a\cdot x)\cdot y\|\leq r.$
We write
$\rho(x)\cdot y=\sum_{\text{weight }\beta}y_{\beta}$
according to the decomposition of the complex vector space $V=\oplus_{\beta}V_{\beta}$ in $\rho$ into weight spaces $V_{\beta}$ relative to $\mathbf{T}$, where $y_{\beta}\in V_{\beta}$. Let $\tilde{\beta}$ be a weight among the weights $\beta$’s in the summation such that $y_{\tilde{\beta}}\neq 0$. Then by the structure of the representation $\rho$, we compute that
(1) $\|\rho(a\cdot x)\cdot y\|=\|e^{\tilde{\beta}(a)}y_{\tilde{\beta}}+\cdots\|\lesssim r.$
Note that $y\in\mathbb{Z}^{d}$ and $x\in\mathcal{K}$, and there is a positive lower bound for $\|y_{\tilde{\beta}}\|$ depending only on $\rho$ and $\mathcal{K}$. On the other hand, there exists a unique discrete primitive subgroup $\tilde{\Lambda}\subset\mathbb{Z}^{d}$ such that $\rho(x)\cdot\tilde{\Lambda}$ is a discrete and Zariski-dense subgroup in $V_{\beta_{0}}$, and for any $w\in\tilde{\Lambda}\setminus\{0\}$ we have
(2) $\|\rho(\tilde{k}\cdot a\cdot x)\cdot w\|\sim\|\rho(a)\cdot\rho(x)\cdot w\|=e^{\beta_{0}(a)}\|\rho(x)\cdot w\|.$
Comparing equations (1) and (2) and using the fact that $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, $a\in T_{\eta}$ and the relation between $\tilde{\beta}$ and $\beta_{0}$ according to the structure of the representation $\rho$, one can deduce that
$e^{\beta_{0}(a)}\lesssim_{\eta}e^{\tilde{\beta}(a)}\lesssim r$
and for any $w\in\tilde{\Lambda}\setminus\{0\}$
$\|\rho(\tilde{k}\cdot a\cdot x)w\|\lesssim r\cdot\|\rho(x)w\|.$
Let $\Lambda_{g}=\rho(g)(\rho(\gamma^{-1})\tilde{\Lambda})$. Then we have
$\Lambda_{g}\subset\rho(g)\cdot\mathbb{Z}^{d},\quad\operatorname{Vol}(\Lambda_{g})\lesssim r^{\operatorname{rank}\Lambda_{g}}=r^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{k}^{-1})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
By the Bruhat decomposition, we know that
$\tilde{k}^{-1}\in\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
By the fact that $\mathbf{P}_{\beta_{0}}$ stabilizes $V_{\beta_{0}}$, we conclude that there exists $\tilde{g}\in\mathcal{C}_{\bar{I}}$ such that
$\tilde{k}^{-1}\in\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot\tilde{g}\textup{ and }\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
This completes the proof of the proposition. ∎
Let $\epsilon_{0}>0$ which we will determine later. Let $g$ be an element in $S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then by definition, there exists a sequence $\{t_{k}\}\subset\mathbb{N}$ such that
$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq\epsilon_{0}\cdot e^{-\tau t_{k}}.$
By Proposition 6.1, for each $k\in\mathbb{N}$, there exist $\tilde{g}_{k}\in\mathcal{C}_{\bar{I}}$ and a discrete subgroup $\Lambda_{k}\subset\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d}$ such that
$\operatorname{Vol}(\Lambda_{k})\lesssim(\epsilon_{0}e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g}_{k})\cdot\Lambda_{k}\textup{ is Zariski dense in }V_{\beta_{0}}.$
Let $\tilde{g}_{k}=w_{i,k}\cdot f_{i,k}\in\mathcal{C}_{\bar{I}}$ where $w_{i,k}\in\{w_{i}\}_{i\in\bar{I}}$ and $f_{i,k}\in\mathbf{F}_{w_{i,k}}(\mathbb{R})$. Then it implies that
$\tilde{q}_{k}:=w_{i,k}\cdot f_{i,k}\cdot a_{t_{k}}\cdot g=w_{i,k}\cdot a_{t_{k}}\cdot(a_{-t_{k}}f_{i,k}a_{t_{k}})\cdot g$
$q_{k}:=(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\tilde{q}_{k}=w_{i,k}(a_{-t_{k}}f_{i,k}a_{t_{k}})g$
are rational elements in $\mathbf{G}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, we can compute that
$d(q_{k})=e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}\cdot d(\tilde{q}_{k})$
$d(\tilde{q}_{k})\cdot e_{V_{\beta_{0}}}=\rho_{\beta_{0}}(\tilde{g}_{k})\cdot\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right),\quad\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}=\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right)/d(\tilde{q}_{k})$
$\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|=$ $\operatorname{Vol}(\Lambda_{k})/d(\tilde{q}_{k})$
$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
where $e_{V_{\beta_{0}}}=\bigwedge^{\dim V_{\beta_{0}}}V_{\beta_{0}}$ is the unit vector which represents the vector space $V_{\beta_{0}}$ in the $\dim V_{\beta_{0}}$-exterior product space of $V$, and $\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}$ is the vector constructed from a $\mathbb{Z}$-basis of $\Lambda_{k}$ and represents the lattice $\Lambda_{k}$. The implicit constants here depend only on $\mathbf{G}$, $\Gamma$ and $\rho$.
Definition 6.2.
Fix an arbitrary $i\in\bar{I}$. For any $R>0$, define
$E_{w_{i}}(R):=\{f\in\mathbf{F}_{w_{i}}(\mathbb{R}):\|\rho_{\beta_{0}}(w_{i}fw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq R\}.$
We also define the following morphism
$\Psi_{w_{i}}:\mathbf{F}_{w_{i}}(\mathbb{R})\to\bigwedge_{j=1}^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(w_{i}xw_{i}^{-1})\cdot e_{V_{\beta_{0}}}.$
Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$. Hence $\Psi_{w_{i}}$ is an isomorphism onto its image.
From the discussion above, we know that
$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim$ $\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|$
$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
and by definition of $E_{w_{i,k}}(R)$, we have
$g=$ $(a_{-t_{k}}f_{i,k}^{-1}a_{t_{k}})\cdot w_{i,k}^{-1}\cdot q_{k}$
$\in$ $\left(a_{-t_{k}}\cdot E_{w_{i,k}}\left(C_{1}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k})\right)\cdot a_{t_{k}}\right)\cdot w_{i,k}^{-1}\cdot q_{k}$
for some constant $C_{1}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Moreover, by the structure of the representation $\rho$, we know that
$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\Psi_{w_{i,k}}(f_{i,k}^{-1})\|\gtrsim\|e_{V_{\beta_{0}}}\|=1$
so
$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
for some constant $C_{2}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Therefore, $q_{k}$ is a rational element in $\mathbf{G}(\mathbb{R})$ with
$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i,k}^{-1}\cdot q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By passing to a subsequence, we may assume that $w_{i,k}$ is a fixed element in $\{w_{i}\}_{i\in\bar{I}}$ independent of $k$. In the following, we take $C_{0}=\max\{C_{1},C_{2}\}$.
Definition 6.3.
For any $i\in\bar{I}$, define $\mathcal{E}_{i}(\tau)$ to be the subset of elements $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
Lemma 6.4.
Let $i\in\bar{I}$. Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $\theta_{i}>0$ depending only on $i$ such that $d(q)\geq\theta_{i}$.
Proof.
By definition, we know that $R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$ By Corollary 2.3, let
$w_{i}^{-1}\cdot q=u\cdot v$
for some $u\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Let $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ be a bounded fundamental domain of $\mathbf{F}_{w_{i}}(\mathbb{R})/(\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R}))$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$. We may write
$v=\tilde{v}\cdot\gamma$
for some $\tilde{v}\in\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R})$. Then
$q\cdot\mathbb{Z}^{d}=(w_{i}\cdot u\cdot v)\cdot\mathbb{Z}^{d}=(w_{i}uw_{i}^{-1})\cdot(w_{i}\tilde{v}\cdot\mathbb{Z}^{d}).$
By the fact that $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ and $\tilde{v}$ is in the bounded subset $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, we can conclude that the co-volume of $(w_{i}\tilde{v}\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}$ has a lower bound depending only on $w_{i}$ and $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, and so does $q\cdot\mathbb{Z}^{d}$. This completes the proof of the lemma. ∎
Now we choose $\epsilon_{0}>0$ so that
$\min_{i\in\bar{I}}\{\theta_{i}\}>C_{0}\cdot\epsilon_{0}^{\dim V_{\beta_{0}}}.$
Lemma 6.5.
If $\mathcal{E}_{i}(\tau)\neq\emptyset$ for some $i\in\bar{I}$, then $0\leq\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$.
Proof.
Let $g\in\mathcal{E}_{i}(\tau)$. By definition, there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 6.4, we have
$\theta_{i}\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
If $\tau\geq\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$, then by taking $t_{k}\to\infty$, we obtain that
$\theta_{i}\leq C_{0}\cdot(\epsilon_{0})^{\dim V_{\beta_{0}}}$
which contradicts the choice of $\epsilon_{0}$. This completes the proof of the lemma. ∎
From the discussion above, we obtain the following
Proposition 6.6.
We have
$S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{i\in\bar{I}}\mathcal{E}_{i}(\tau).$
In the following, for each $i\in\bar{I}$, we will consider the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ $(\mathcal{E}_{i}(\tau)\neq\emptyset)$ and compute an upper bound for the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$. These upper bounds will give an upper bound for $\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ by Proposition 6.6. For notational convenience, we may set $\epsilon_{0}=1$ as it does not affect the computations in the rest of this section.
Now fix $i\in\bar{I}$. To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we first need to estimate the volume of the subset $E_{w_{i}}(R)$ in ${\mathbf{F}}_{w_{i}}(\mathbb{R})$, which can be written as
$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))$
where $B_{R}$ is the ball of radius $R$ centered at 00 in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ and $\mu_{\mathbf{F}_{w_{i}}}$ is the Haar measure on ${\mathbf{F}}_{w_{i}}(\mathbb{R})$. We need the following result about the asymptotic volume estimates of algebraic varieties.
Theorem 6.7 ([3, Corollary 16.3]).
[…]
$\mu(B_{R})\sim cR^{a}(\log R)^{b}\;(\textup{as }R\to\infty)$
for some $a\in\mathbb{Q}_{\geq 0}$, $b\in\mathbb{Z}_{\geq 0}$ and $c>0$.
We apply Theorem 6.7 to the morphism $\Psi_{w_{i}}$ where the closed orbit $\mathcal{O}=\Psi_{w_{i}}(\mathbf{F}_{w_{i}}(\mathbb{R}))$ and the measuer $\mu$ is the push-forward $\Psi_{w_{i}}^{*}(\mu_{\mathbf{F}_{w_{i}}})$ of the Haar measure on $\mathbf{F}_{w_{i}}(\mathbb{R})$, and obtain the following
Corollary 6.8.
There exist constants $a_{w_{i}}\in\mathbb{Q}_{\geq 0}$, $b_{w_{i}}\in\mathbb{Z}_{\geq 0}$ and $c_{w_{i}}>0$ such that
$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))\sim c_{w_{i}}R^{a_{w_{i}}}(\log R)^{b_{w_{i}}}\;(\textup{as }R\to\infty).$
To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we also need an upper bound for the asymptotic number of rational elements in
$(w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q}).$
Consider the morphism
$\tilde{\Psi}_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V,\quad\tilde{\Psi}_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
Let $g\in w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in a bounded set $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$. By definition, $d(g)$ is the co-volume of $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, this implies that there exists a constant $C>0$ depending only on $\tilde{U}$, such that the length of the primitive integral vector in the line spanned by $\rho(g^{-1})\cdot e_{V_{\beta_{0}}}$ is less than $C\cdot d(g)$. So the number of rational points $g$ in $\tilde{U}$ whose denominators are less than $l>0$ is less than or equal to the total number of primitive integral points in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ whose lengths are less than $C\cdot l$. This leads us to the results about the Manin’s conjecture [2] (for relevant results about the Manin’s conjecture, one may refer to [11, 16, 18, 19, 32, 36, 38, 40] and the references therein.) In particular, there exists a constant $A_{w_{i}}>0$ such that for any $\epsilon>0$
($\ast$) #|$\#\left|\{g\in\tilde{U}\cap w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q}):d(g)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{A_{w_{i}}+\epsilon}$
where the implicit constant depends only on $\epsilon$, $\tilde{U}$, $\mathbf{G}$ and $\rho$. In the following, we will fix the constants $a_{w_{i}},b_{w_{i}},c_{w_{i}}$ and $A_{w_{i}}>0$.
Lemma 6.9.
Let $i\in\bar{I}$ such that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0.$ Then
$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{w_{i}})\}>0.$
Proof.
The proof is similar to those of Lemma 4.2 and Lemma 5.5. Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{w_{i}})$, $\alpha(a_{1})=0$. Let $\Omega_{\mathbf{F}_{w_{i}}}$ and $\Omega_{\mathbf{H}_{w_{i}}}$ be small open neighborhoods of identity in $\mathbf{F}_{w_{i}}(\mathbb{R})$ and $\mathbf{H}_{w_{i}}(\mathbb{R})$ respectively, and
$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{w_{i}}}\cdot\Omega_{\mathbf{H}_{w_{i}}}.$
Note that $w_{i}\mathbf{H}_{w_{i}}(\mathbb{R})w_{i}^{-1}$ stabilizes every element in $V_{\beta_{0}}$. Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{w_{i}}}$ and $h\in\Omega_{\mathbf{H}_{w_{i}}}$, we have
$a_{t}\cdot p\cdot w_{i}^{-1}=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}w_{i}^{-1}=fw_{i}^{-1}\cdot(w_{i}a_{t}ha_{-t}w_{i}^{-1})\cdot w_{i}a_{t}w_{i}^{-1}.$
The element $w_{i}a_{t}w_{i}^{-1}$ is rational and
$\delta(\rho(w_{i}a_{t}w_{i}^{-1})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(w_{i}a_{t}w_{i}^{-1})}.$
This implies that
$\delta(\rho(a_{t}\cdot p\cdot w_{i}^{-1})\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(w_{i}a_{-1}w_{i}^{-1})t}$
for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{w_{i}}}$ and $w_{i}$. Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$. Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot w_{i}^{-1}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})/\Gamma$. Set $\Gamma_{w_{i}}=\Gamma\cap w_{i}^{-1}\Gamma w_{i}$. Then $\Gamma_{w_{i}}$ is commensurable with $\Gamma$ and $w_{i}^{-1}\Gamma w_{i}$, and we can conclude that for any $p\Gamma_{w_{i}}\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot\Gamma_{w_{i}}$, the orbit $a_{t}\cdot p\Gamma_{w_{i}}$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma_{w_{i}}$. Similiar to Lemma 4.2 and Lemma 5.5, this contradicts the ergodic property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\Gamma_{w_{i}}\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$, where $\tilde{\mathbf{G}}$ is the product of some $\mathbb{Q}$-simple factors $\mathbf{G}_{i}$ $(1\leq i\leq s)$ of $\mathbf{G}$ such that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})$ $(1\leq i\leq s)$. This completes the proof of the lemma. ∎
Now we compute an upper bound for the Hausdorff dimension of the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$. In the computation, we project the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by the natural projection map
$\pi_{R_{u}(\overline{\mathbf{P}}_{0})}:R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\to R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
and compute the Hausdorff dimension of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$. Since $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}$ is a local isometry, by the countable stability of Hausdorff dimension, we then obtain the upper bounds for $\dim_{H}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}^{-1}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))))$ and $\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$.
Let $\tilde{\Omega}$ be a bounded fundamental domain of $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We choose a bounded subset $\Omega_{\mathbf{H}_{w_{i}}}$ in $\mathbf{H}_{w_{i}}(\mathbb{R})$ and a bounded subset $\Omega_{\mathbf{F}_{w_{i}}}$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w_{i}}}\cdot\Omega_{\mathbf{F}_{w_{i}}}.$
Now suppose that $q\in\mathbf{G}(\mathbb{R})$ is a rational element with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Let $\tilde{q}\in\mathbf{G}(\mathbb{R})$ such that $w_{i}^{-1}\tilde{q}$ is a representative of $w_{i}^{-1}q$ in $\tilde{\Omega}$ with
$w_{i}^{-1}\tilde{q}=w_{i}^{-1}q\gamma$
for some $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then one can check that $\tilde{q}$ is still a rational element in $\mathbf{G}(\mathbb{R})$ and $d(\tilde{q})=d(q)$ since $\Gamma$ preserves the lattice $\mathbb{Z}^{d}$. So by Corollary 2.3, we still have
$w_{i}^{-1}\tilde{q}\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Since $w_{i}^{-1}\tilde{q}$ is an element in the bounded fundamental domain $\tilde{\Omega}$, it implies that
$w_{i}^{-1}\tilde{q}\in\Omega_{\mathbf{H}_{w_{i}}}\cdot(\mathbf{F}_{w_{i}}(\mathbb{Q})\cap\Omega_{\mathbf{F}_{w_{i}}}).$
We then obtain the following
Proposition 6.10.
The set $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ consists of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$d(q_{k})\leq C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in\tilde{\Omega}$.
Lemma 6.11.
Let $\delta>0$ and $B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ be the small open ball of radius $\delta$ centered at identity in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $C_{\delta}>0$ depending only on $\delta$ such that for any $R>0$
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}(R)\subset E_{w_{i}}(C_{\delta}\cdot R)\cdot\mathbf{H}_{w_{i}}(\mathbb{R}).$
Proof.
Let $x\in B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ and $y\in E_{w_{i}}(R)$. Then $x\cdot y\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We write
$x\cdot y=u\cdot v$
where $u\in\mathbf{F}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{H}_{w_{i}}(\mathbb{R})$. Then we have
$\|\Psi_{w_{i}}(u)\|=$ $\|\rho_{\beta_{0}}(w_{i}uw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\rho(w_{i}uvw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|$
$=$ $\|\rho(w_{i}xyw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq C_{\delta}\cdot R$
for some constant $C_{\delta}>0$ depending only on $\delta>0$. This completes the proof of the lemma. ∎
Note that by definition, for any rational element $g\in\mathbf{G}(\mathbb{R})$, every element $x$ in the $\Gamma$-coset $g\Gamma$ of $g$ is also rational, and $d(x)=d(g)$. We will use this fact in the following lemma.
Lemma 6.12.
Let $i\in\bar{I}$ and $L>0$. Then the set of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$ is equal to
$\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}.$
Proof.
Let $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in R_{u}(\overline{\mathbf{P}}_{0})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$. We write
$g=(u\cdot v)\cdot\gamma$
for some $u\in\Omega_{\mathbf{H}_{w_{i}}}$, $v\in\Omega_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})$. Then we have
$w_{i}\cdot g=(w_{i}\cdot u\cdot w_{i}^{-1})\cdot w_{i}\cdot v\cdot\gamma$
and $w_{i}\cdot v$ is rational with
$d(w_{i}\cdot g)=d(w_{i}\cdot v)=L.$
By Corollary 2.3, we know that $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. This completes the proof of the lemma. ∎
Remark 6.13.
For any $L>0$, we define
$\mathcal{S}(L):=\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}$
$\mathcal{F}(L):=\{y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}):d(w_{i}\cdot y)=L\}.$
Now we construct open covers of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ with arbitrarily small diameters. Here we use the right-invariant metric $d_{R_{u}(\overline{\mathbf{P}}_{0})}$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which induces a metric $d_{R_{u}(\overline{\mathbf{P}}_{0})/\Gamma}$ on the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$. Note that by Corollary 6.8, for any $\epsilon>0$ and $R>0$, we have
$\operatorname{Vol}(E_{w_{i}}(R))\lesssim_{\epsilon}R^{a_{w_{i}}+\epsilon}$
where the implicit constant depends only on $\epsilon$ and $i\in\bar{I}$. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
By Lemma 6.9, we have $\nu_{0}(a_{1})>0$.
Fix $\delta>0$. Let $L>0$ and $t\in\mathbb{N}$ such that
$L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ such that $d(q)=L$ and $w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ The following subset
$\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
can be covered by disjoint boxes of diameter at most $\delta\cdot\exp(\nu_{0}(a_{-t}))$, and by Lemma 6.11 and Lemma 6.12, we know that these boxes are contained in
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)\cdot a_{t}\right)\cdot\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L).$
Since every point $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)$ satisfies the property that $w_{i}\cdot g$ is rational and $d(w_{i}\cdot g)=L$, it follows again from Lemma 6.12 that
$\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)=\mathcal{S}(L).$
We conclude that the subset
$\left(a_{-t}E_{w_{i}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q))a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
can be covered by disjoint boxes with diameter at most $\delta\cdot\exp(\nu_{0}(a_{-t}))$ and these boxes are contained in
$a_{-t}E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)a_{t}\cdot\mathcal{S}(L).$
We collect all these boxes constructed above in a family which is denoted by $\mathcal{P}_{t,L}(q)$, and define
$\mathcal{P}_{t,L}=\bigcup_{d(q)=L\textup{ and }w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})}\mathcal{P}_{t,L}(q).$
Now since every box in $\mathcal{P}_{t,L}$ is contained in the same bounded set
$a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t}\cdot\mathcal{S}(L)$
we can choose a maximal finite sub-collection $\mathcal{Q}_{t,L}$ of disjoint boxes in $\mathcal{P}_{t,L}$. Then by the maximality of $\mathcal{Q}_{t,L}$ we have
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\mathcal{Q}_{t,L}\supset\bigcup_{S\in\mathcal{P}_{t,L}}S.$
Meanwhile, the number of boxes in $\mathcal{Q}_{t,L}$ is at most
$\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}.$
Now we define $\mathcal{G}_{t}$ to be the collection of boxes in
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\mathcal{Q}_{t,L}$
where
$\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
Note that $\operatorname{diam}\mathcal{G}_{t}\sim e^{\nu_{0}(a_{-t})}$ and $\nu_{0}(a_{-t})\neq 0$ by Lemma 6.9. Define
$\mathcal{G}^{k}=\bigcup_{t\geq k}\mathcal{G}_{t}.$
Then by definition and Proposition 6.10, the subset $\mathcal{G}^{k}$ is a cover of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ for any $k\in\mathbb{N}$ and $\operatorname{diam}\mathcal{G}^{k}\sim e^{\nu_{0}(a_{-k})}$.
Now according to the construction of the open covers $\mathcal{G}^{k}$ $(k\in\mathbb{N})$ above, we consider the following series with respect to the parameter $s$
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}\leq\sum_{t\in\mathbb{N}}\sum_{B\in\mathcal{G}_{t}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}\cdot e^{s\nu_{0}(a_{-t})}.$
By Corollary 6.8, for any $\epsilon>0$, we have
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)^{a_{w_{i}}+\epsilon}\cdot|\mathcal{F}(L)|$
$e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}.$
Fix $t\in\mathbb{N}$. Note that $\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$. Choose $t_{0}\in\mathbb{Z}$ such that
$C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot t_{0}})^{\dim V_{\beta_{0}}}\leq\theta_{i}.$
Then by equation ($\ast$6) before Lemma 6.9, we have
$\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\sum_{C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}}\leq L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}\sum_{L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|$
$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}})^{A_{w_{i}}+\epsilon}$
$\sim$ $\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
Combining all these equations, we obtain that
$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
$\lesssim$ $\sum_{t\in\mathbb{N}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)t})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}$
$\cdot\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
By computing the series above in the cases $A_{w_{i}}<a_{w_{i}}$, $A_{w_{i}}=a_{w_{i}}$ and $A_{w_{i}}>a_{w_{i}}$, one can conclude that the series above converges if
$s>\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
This implies that for any $\epsilon>0$
$\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))\leq\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
By taking $\epsilon\to 0$ and Proposition 6.6, we conclude that
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\leq\max_{i\in\bar{I}}\left\{\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}$
and
$\dim_{H}S_{\tau}^{c}\leq\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}.$
This completes the proof of Theorem 1.4.
7. Proofs of Theorems 1.61.7 and 1.8
In this section, we prove Theorems 1.61.7 and 1.8.
Proof of Theorem 1.6.
Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Without loss of generality, we may write
$\{a_{t}\}_{t\in\mathbb{R}}=\left\{\operatorname{diag}(e^{b_{1}t},e^{b_{2}t},\cdots,e^{b_{n}t}):t\in\mathbb{R}\right\}$
where $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$. We write
$\mathbf{T}(\mathbb{R})^{0}=\{\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}}):t_{1}+t_{2}+\cdots+t_{n}=0\}$
where $\mathbf{T}(\mathbb{R})^{0}$ is the connected component of identity in $\mathbf{T}(\mathbb{R})$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$ and $R_{u}(\overline{\mathbf{P}}_{0})$ is the upper triangular unipotent subgroup in $\mathbf{G}$. Then all the $\mathbb{Q}$-roots in $\mathbf{G}$ with respect to $\mathbf{T}$ are
$\alpha_{i,j}(a)=t_{i}-t_{j}\quad(1\leq i\neq j\leq n)$
where $a=\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}})$. Let $\{e_{1},e_{2},...,e_{n}\}$ be the standard basis in $V$ where
$e_{1}=\left(\begin{array}[]{c}1\\ 0\\ 0\\ \vdots\\ 0\end{array}\right)\quad e_{2}=\left(\begin{array}[]{c}0\\ 1\\ 0\\ \vdots\\ 0\end{array}\right)\quad\cdots\cdots\quad e_{n}=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\\ 1\end{array}\right).$
We choose $V_{\lambda_{0}}=V_{\beta_{0}}=\mathbb{C}\cdot e_{n}$ to be the highest weight space and
$V=\mathbb{C}\cdot e_{1}\oplus\mathbb{C}\cdot e_{2}\oplus\cdots\oplus\mathbb{C}\cdot e_{n}$
is the weight space decomposition of $V$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are
$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{cccc}*&*&\cdots&0\\ &*&\cdots&0\\ \vdots&\vdots&\cdots&\vdots\\ &*&\cdots&*\end{array}\right)\textup{ and }R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\left(\begin{array}[]{cc}I_{n-1}&*\\ 0&1\end{array}\right).$
Suppose that $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ is a rational element in $\mathbf{G}(\mathbb{R})$. Let
$x=\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)$
where $p_{i}\in\mathbb{Z},q\in\mathbb{N}$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then $x\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}\neq\{0\}$ implies that there exists a primitive integral point $(k_{1},k_{2},\dots,k_{n-1},l)^{T}\in\mathbb{Z}^{n}\setminus\{0\}$ such that
$\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)\cdot\left(\begin{array}[]{c}k_{1}\\ k_{2}\\ \vdots\\ l\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ \end{array}\right).$
Then we have
$k_{i}+\frac{p_{i}}{q}\cdot l=0,\quad 1\leq i\leq n-1.$
Hence $|l|=q$ and the denominator $d(x)$ of $x$ is equal to $q$. One can then deduce from the formula $d(x)=q$ that for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and any sufficiently large $l>0$, the number of rational elements in $\tilde{U}$ whose denominators are less than $l$ is bounded by $\mu_{R_{u}(\overline{\mathbf{P}}_{\beta_{0}})}(\tilde{U})\cdot l^{n}.$ Therefore, we may choose $A_{e}=n$ in Equation ($\ast$6) in §6. One can also compute that $a_{e}=n-1$.
Now consider the upper bound of the Hausdorff dimension of $S_{\tau}^{c}$. In the computation, we actually use the Bruhat decomposition as
$\mathbf{G}(\mathbb{R})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
In our case, we may choose the index set $\bar{I}$ as small as possible (as the double cosets in the original decomposition with respect to Weyl group ${}_{\mathbb{Q}}W$ overlap). Note that ${}_{\mathbb{Q}}W$ is isomorphic to the symmetric group of permutations of $\{1,2,\dots,n\}$ and many elements of ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}(\mathbb{R})$. From this observation, one can deduce that if $w_{i}=(i,n)$ $(1\leq i\leq n)$ is the permutation of $i$ and $n$, then
$\mathbf{G}(\mathbb{R})=\bigcup_{i=1}^{n}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
Then by definition, one may compute that $A_{w_{i}}\leq A_{e}$, $a_{w_{i}}\leq a_{e}$ for all $1\leq i\leq n$, and
$\mathbf{F}_{w_{i}}=\left(\begin{array}[]{cccccc}1&0&\cdots&*&\cdots&0\\ 0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&\cdots&1\end{array}\right)$
where the only non-zero entries are in the $i$-th column.
Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$. Fix $w_{i}$ $(1\leq i\leq n)$. Note that the roots in $\mathbf{F}_{w_{i}}$ are $\alpha_{1,i},\alpha_{2,i},\dots,\alpha_{i-1,i}$. According to Theorem 1.4 and the discussion in §6, we may assume that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0$, i.e. $b_{i}<0$; otherwise there is nothing to compute. Then $\nu_{0}=\alpha_{1,n}$ and by the fact that $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+\frac{-b_{i}-\tau}{b_{1}-b_{n}}\cdot\max\{A_{w_{i}},a_{w_{i}}\}$
$\leq$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+n\cdot\frac{-b_{i}-\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})+nb_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{-b_{i+1}\cdots-b_{n}+(n-i)b_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{(b_{i}-b_{i+1})+\cdots(b_{i}-b_{n})}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$\leq$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}.$
Therefore, by Theorem 1.4, we have
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
On the other hand, for the lower bound of $\dim_{H}S_{\tau}^{c}$, by Theorem 1.3, we have
$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
$=$ $\dim\mathbf{G}-((b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n}))\cdot\frac{\tau}{(-b_{n})(b_{1}-b_{n})}$
$=$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
We conclude that
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
This completes the proof of Theorem 1.6. ∎
Proof of Theorem 1.7.
Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and we may write
$\{a_{t}\}_{t\in\mathbb{R}}=\left\{\operatorname{diag}(e^{b_{1}t},e^{b_{2}t},\cdots,e^{b_{n}t}):t\in\mathbb{R}\right\}$
where $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$. We write
$\mathbf{T}(\mathbb{R})^{0}=\{\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}}):t_{1}+t_{2}+\cdots+t_{n}=0\}$
where $\mathbf{T}(\mathbb{R})^{0}$ is the connected component of identity in $\mathbf{T}(\mathbb{R})$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$ and $R_{u}(\overline{\mathbf{P}}_{0})$ is the upper triangular unipotent subgroup in $\mathbf{G}$. Let $V_{\beta_{0}}$ be the highest weight space in $\mathfrak{sl}_{n}$
$V_{\beta_{0}}=\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ &0&\cdots&0\end{array}\right).$
Note that here $\nu_{0}=-\beta_{0}$ is the highest root in $\mathfrak{sl}_{n}$ and $\beta_{0}(a_{t})=(b_{n}-b_{1})t<0$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are
$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{ccccc}*&0&\cdots&0&0\\ &*&\cdots&*&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ &*&\cdots&*&0\\ &*&\cdots&*&*\end{array}\right)\textup{ and }R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\left(\begin{array}[]{ccccc}1&*&\cdots&*&*\\ 0&1&\cdots&0&*\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&*\\ 0&0&\cdots&0&1\end{array}\right).$
Let $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in $\mathbf{G}(\mathbb{R})$. We write
$x=\left(\begin{array}[]{ccccc}1&\frac{a_{1}}{b}&\cdots&\frac{a_{n-2}}{b}&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&0&\cdots&0&\frac{p_{1}}{q}\\ 0&1&\cdots&0&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&\frac{p_{n-1}}{q}\\ 0&0&\cdots&0&1\end{array}\right)$
where $(a_{1},a_{2},\dots,a_{n-2},b)=1$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then by definition,
$x\cdot\mathfrak{sl}_{n}(\mathbb{Z})\cdot x^{-1}\cap V_{\beta_{0}}\neq\{0\}.$
It implies that there exists $l\in\mathbb{R}$ such that
$x^{-1}\cdot\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ l&0&\cdots&0\end{array}\right)\cdot x\in\mathfrak{sl}_{n}(\mathbb{Z}).$
One can compute that
$\left(\begin{array}[]{ccccc}-\frac{p_{1}}{q}\cdot l&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ -\frac{p_{n-1}}{q}\cdot l&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ l&l\cdot\frac{a_{1}}{b}&\cdots&l\cdot\frac{a_{n-2}}{b}&l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\end{array}\right)\in\mathfrak{sl}_{n}(\mathbb{Z})$
and consequently
$\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q},\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{a_{i}}{b}(1\leq i\leq n-2),\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{d}{bq}\in\mathbb{Z}$
$l,\;l\cdot\frac{a_{1}}{b},\dots,l\cdot\frac{a_{n-2}}{b},\;l\cdot\frac{d}{bq}\in\mathbb{Z}$
where $d=bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}$. Then one can deduce from the formulas above that
$l=s\cdot\frac{bq^{2}}{\gcd(d,q)}\;(s\in\mathbb{Z}).$
Therefore, the denominator $d(x)$ of $x$ is equal to
$d(x)=\frac{bq^{2}}{\gcd(d,q).}$
It follows from the same argument as in [17, Lemma 8.6] that
|$\left|\{x\in\tilde{U}:d(x)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{n-1+\epsilon}$
for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$, and hence $A_{e}=n-1$.
Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$ using Theorem 1.4. We choose elements $\{w_{i}\}_{i\in\bar{I}}$ in the Weyl group ${}_{\mathbb{Q}}W$ such that the Bruhat decomposition stated in Theorem 1.4 holds
$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Note that many Weyl elements in ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}$. Here we choose $\{w_{i}\}_{i\in\bar{I}}$ as follows. We identify ${}_{\mathbb{Q}}W$ with the symmetric group of permutations on $\{1,2,...,n\}$. Then there are three types of elements in $\{w_{i}\}_{i\in\bar{I}}$:
Type I:
$(1,2,\dots,n-1,n)$, $(n,2,\dots,n-1,1)$
Type II:
$(1,2\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$, $(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$, $(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1)
Type III:
$(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1).
The index set $\bar{I}$ has cardinality $n(n-1)$, and one may check that $\{w_{i}\}_{i\in\bar{I}}$ meets our requirement. Note that $A_{w_{i}}\leq A_{e}$ for any $i\in\bar{I}$.
For a Weyl element $w$ of type I in $\{w_{i}\}_{i\in\bar{I}}$, if $w=(1,2,\dots,n)$ is the identity element, then the number $a_{w}$ in Corollary 6.8 comes from the morphism $\Psi_{w}$ defined by
$\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}0&0&\cdots&0&0\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ 1&0&\cdots&0&0\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)^{-1}$
$=$ $\left(\begin{array}[]{ccccc}z&-x_{1}\cdot z&\cdots&-x_{n-2}\cdot z&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)z\\ y_{1}&-x_{1}\cdot y_{1}&\cdots&-x_{n-2}\cdot y_{1}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ y_{n-2}&-x_{1}\cdot y_{n-2}&\cdots&-x_{n-2}\cdot y_{n-2}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{n-2}\\ 1&-x_{1}&\cdots&-x_{n-2}&x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z\end{array}\right)$
where $x_{1},x_{2},\dots,x_{n-2},y_{1},y_{2},\dots,y_{n-2},z$ are the variables of $\Psi_{w}$. One can estimate $\operatorname{Vol}(E_{w}(R))$ and compute that $a_{w}\leq n-1$. Then according to Theorem 1.4, for $w$, we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$=$ $\dim\mathbf{G}-(n-1)+\frac{(b_{1}-b_{n})-\tau}{b_{1}-b_{n}}\cdot(n-1)=\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Note that for $w=(n,2,3,\cdots,n-1,1)$, $\beta_{0}(wa_{-1}w^{-1})\leq 0$ and there is nothing to compute in this case.
Now let us check a Weyl element $w$ of type II in $\{w_{i}\}_{i\in\bar{I}}$. If $w=(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$ or $w=(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1), then $\beta_{0}(w\cdot a_{-1}\cdot w^{-1})\leq 0$ and there is nothing to compute. If $w=(1,2,\dots,j-1,n,j+1,\dots,n-1,j)$ $(2\leq j\leq n-1)$, one can compute that
$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccc}1&*&*&\cdots&*&\cdots&*\\ 0&1&0&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&0&\cdots&1\end{array}\right).$
One can then deduce that $a_{w}\leq n-1$, which is similar to the type I case. Then under the condition that $(\beta_{0}(wa_{-1}w^{-1})>\tau$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{1}-b_{2})+(b_{1}-b_{3})+\cdots+(b_{1}-b_{n})}{b_{1}-b_{n}}+\frac{(b_{2}-b_{j})+(b_{3}-b_{j})+\cdots+(b_{j-1}-b_{j})}{b_{1}-b_{n}}\right)$
$\hskip 28.45274pt+\frac{(b_{1}-b_{j})-\tau}{b_{1}-b_{n}}\cdot(n-1)$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{j+1})+(b_{1}-b_{j+2})+\cdots+(b_{1}-b_{n})-(n-j)(b_{1}-b_{j})}{b_{1}-b_{n}}-\frac{\tau}{b_{1}-b_{n}}\cdot(n-1)$
$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Similarly, for $w=(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$ ($12\leq j\leq n-$ 1),
$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccc}1&0&\cdots&0&0&\cdots&*\\ 0&1&\cdots&0&0&\cdots&*\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&*&\cdots&*\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&1\end{array}\right)$
and $a_{w}\leq n-1$. So under the condition that $(\beta_{0}(wa_{-1}w^{-1})>\tau$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n})}{b_{1}-b_{n}}+\frac{(b_{j}-b_{j+1})+(b_{j}-b_{j+2})+\cdots+(b_{j}-b_{n-1})}{b_{1}-b_{n}}\right)$
$\hskip 28.45274pt+\frac{(b_{j}-b_{n})-\tau}{b_{1}-b_{n}}\cdot(n-1)$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{j}-b_{n})-j(b_{j}-b_{n})}{b_{1}-b_{n}}-\frac{\tau}{b_{1}-b_{n}}\cdot(n-1)$
$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Let us check a Weyl element $w$ of type III. If $w=(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1), then
$\beta_{0}(w\cdot a_{-1}\cdot w^{-1})\leq 0$
and there is nothing to prove. If $w=(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, then one can compute that
$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccccc}1&0&\cdots&0&0&\cdots&*&\cdots&0\\ 0&1&\cdots&0&0&\cdots&*&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1&*&\cdots&*&\cdots&*\\ 0&0&\cdots&0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&0&\cdots&1\end{array}\right)$
and $a_{w}\leq n-1$. Then under the condition that $\beta_{0}(wa_{-1}w^{-1})>\tau$, we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{i}-b_{i+1})+(b_{i}-b_{i+2})+\cdots+(b_{i}-b_{n})}{\nu_{0}(a_{1})}+\frac{(b_{1}-b_{j})+(b_{2}-b_{j})+\cdots+(b_{j-1}-b_{j})}{\nu_{0}(a_{1})}-\frac{b_{i}-b_{j}}{\nu_{0}(a_{1})}\right)$
$\hskip 28.45274pt+\frac{(b_{i}-b_{j})-\tau}{\nu_{0}(a_{1})}\cdot(n-1)$
$=$ $\dim\mathbf{G}-\frac{(n-i)b_{i}+b_{1}+\cdots+b_{i}-b_{j}-\cdots-b_{n}-(j-1)b_{j}-n(b_{i}-b_{j})}{\nu_{0}(a_{1})}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}$
$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
Combining type I, type II, type III cases, we conclude that
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
On the other hand, by Theorem 1.3, we know that
$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
$=$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
$=$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
This completes the proof of Theorem 1.7. ∎
Proof of Theorem 1.8.
It is known that there is a complete classification of complex-linear finite-dimensional irreducible representations of $\mathbf{G}=\operatorname{\mathbf{SL}}_{2}$ described as follows. Let $V_{n}$ be the complex vector space of homogeneous polynomials of degree $n$ in two variables $x$ and $y$ ($\dim V_{n}=n+1$). Define $\rho_{n}:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{SL}}_{n+1}(V_{n})$ by
$\rho(g)\cdot f\left(\begin{array}[]{c}x\\ y\end{array}\right):=f\left(g^{-1}\cdot\left(\begin{array}[]{c}x\\ y\end{array}\right)\right)$
for any $g\in\operatorname{\mathbf{SL}}_{2}$ and $f\in V_{n}$. Then $\{\rho_{n}\}_{n\in\mathbb{N}}$ consists of all the complex-linear finite dimensional irreducible representations of $\operatorname{\mathbf{SL}}_{2}$ up to equivalence.
Without loss of generality, we may assume that $\rho=\rho_{n}$ for some $n\in\mathbb{N}$ and write $\{a_{t}\}_{t\in\mathbb{R}}\subset\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ as
$\left\{a_{t}=\left(\begin{array}[]{cc}e^{\alpha t}&0\\ 0&e^{-\alpha t}\end{array}\right):t\in\mathbb{R}\right\}$
for some $\alpha>0$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$. The standard basis of $V_{n}$ is $\{x^{n},x^{n-1}y,\cdots,xy^{n-1},y^{n}\}.$ We choose $\mathbb{C}\cdot x^{n}$ to be the highest weight space $V_{\beta_{0}}$ in $V_{n}$ where $\beta_{0}(a_{t})=-n\alpha\cdot t$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ is the subgroup
$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{cc}*&0\\ &*\end{array}\right)$
which coincides with the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ of $\operatorname{\mathbf{SL}}_{2}$, and
$R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=R_{u}(\overline{\mathbf{P}}_{0})=\left(\begin{array}[]{cc}1&*\\ 0&1\end{array}\right).$
The Bruhat decomposition of $\operatorname{\mathbf{SL}}_{2}$ is the following
$\operatorname{\mathbf{SL}}_{2}(\mathbb{R})=\mathbf{P}_{\beta_{0}}\cdot w_{0}\cdot R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cup\mathbf{P}_{\beta_{0}}\cdot w_{1}$
where $w_{0}$ is the identity element
$w_{1}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$
and ${}_{\mathbb{Q}}W=\{w_{0},w_{1}\}$.
Now we consider the upper bound of $\dim_{H}S_{\tau}^{c}$. Let $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and write
$g=\left(\begin{array}[]{cc}1&t\\ 0&1\end{array}\right).$
Then one can compute that
$\rho(g)\cdot(x^{n},x^{n-1}y,\dots,y^{n})$
$=$ $(x^{n},x^{n-1}y,\dots,y^{n})\cdot\left(\begin{array}[]{cccccc}{n\choose n}(-t)^{0}&0&\cdots&0&\cdots&0\\ {n\choose n-1}(-t)^{1}&{n-1\choose n-1}(-t)^{0}&\cdots&0&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose i}(-t)^{i}&{n-1\choose i}(-t)^{i-1}&\cdots&{i\choose i}(-t)^{0}&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose 0}(-t)^{n}&{n-1\choose 0}(-t)^{n-1}&\cdots&{i\choose 0}(-t)^{i}&\cdots&{0\choose 0}(-t)^{0}\end{array}\right).$
By Corollary 6.8 and the formula above, we have $a_{w_{0}}=1/n$ for the morphism $\Psi_{w_{0}}$ defined by
$\rho(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R}))\cdot\{x^{n}\}.$
Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ with
$g=\left(\begin{array}[]{cc}1&\frac{p}{q}\\ 0&1\end{array}\right)\quad(p\in\mathbb{Z},\;q\in\mathbb{N},\;\gcd(p,q)=1).$
Then $\rho(g)\cdot\mathbb{Z}^{n+1}\cap V_{\beta_{0}}\neq\{0\}$ and there exists a primitive integral vector $(k_{n},k_{n-1},\dots,k_{1},l)\in\mathbb{Z}^{n+1}\setminus\{0\}$ such that
($\left(\begin{array}[]{cccccc}{n\choose n}(-\frac{p}{q})^{0}&0&\cdots&0&\cdots&0\\ {n\choose n-1}(-\frac{p}{q})^{1}&{n-1\choose n-1}(-\frac{p}{q})^{0}&\cdots&0&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose i}(-\frac{p}{q})^{n-i}&{n-1\choose i}(-\frac{p}{q})^{n-i-1}&\cdots&{i\choose i}(-\frac{p}{q})^{0}&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose 0}(-\frac{p}{q})^{n}&{n-1\choose 0}(-\frac{p}{q})^{n-1}&\cdots&{i\choose 0}(-\frac{p}{q})^{i}&\cdots&{0\choose 0}(-\frac{p}{q})^{0}\end{array}\right)\cdot\left(\begin{array}[]{c}l\\ k_{1}\\ \vdots\\ \vdots\\ \vdots\\ k_{n}\end{array}\right)=\left(\begin{array}[]{c}*\\ 0\\ \vdots\\ 0\\ \vdots\\ 0\end{array}\right).$
One can deduce from the equation above that
$l\cdot{n\choose i}\cdot\left(\frac{p}{q}\right)^{n-i}=k_{n-i}\;(\forall 1\leq i\leq n),\quad|l|=q^{n}$
and $d(g)=q^{n}$. According to equation ($\ast$6) in §6, one can compute that $A_{1}=2/n$. Note that for the Weyl element $w_{1}$, $\beta_{0}(w_{1}a_{-1}w_{1}^{-1})<0$. So by Theorem 1.4, we obtain
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-1+\frac{n\alpha-\tau}{2\alpha}\cdot\frac{2}{n}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
On the other hand, by Theorem 1.3, we have
$\dim_{H}S_{\tau}^{c}\geq\dim\mathbf{G}-\frac{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
Therefore, we can conclude that $\dim_{H}S_{\tau}^{c}=3-\tau/\beta_{0}(a_{-1})$. ∎
Let us explain how to deduce [17, Corollary 1.3] (or equivalently [17, Theorem 1.2]) from Theorem 1.7. In [17], we consider a regular one-parameter diagonal subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ acting on the homogeneous space $X_{3}=\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$. According to [17, Definition 1.1], a point $p=g\cdot\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\in X_{3}$ is Diophantine of type $\tau$ if and only if there exists a constant $C>0$ such that for any $t>0$
$\eta(a_{t}\cdot p)\geq Ce^{-\tau t}$
where $\eta$ is the injectivity radius function on $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$. Now let $p=g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\in X_{3}$ which is not Diophantine of type $\tau$. Then by definition, for any $\epsilon>0$, there exists $t_{\epsilon}>0$ such that
$\eta(a_{t_{\epsilon}}\cdot p)<\epsilon e^{-\tau t_{\epsilon}}.$
By [37, Corollary 11.18], for sufficiently small $\epsilon>0$, there exists a unipotent element $u_{\epsilon}\in\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\setminus\{e\}$ such that $g\cdot u_{\epsilon}\cdot g^{-1}\in\operatorname{Stab}(p)=g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})g^{-1}$ (the stabilizer of $p$) and
$d_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g\cdot u_{\epsilon}\cdot g^{-1}a_{-t_{\epsilon}},e)<\epsilon e^{-\tau t_{\epsilon}}$
where $d_{\operatorname{\mathbf{SL}}_{3}}$ is the metric on $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$ induced by a norm $\|\cdot\|_{\mathfrak{sl}_{3}}$ on the Lie algebra $\mathfrak{sl}_{3}(\mathbb{R})$ of $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$. Note that $u_{\epsilon}$ is unipotent
$\log u_{\epsilon}=(u_{\epsilon}-I_{3})-\frac{(u_{\epsilon}-I_{3})^{2}}{2}$
and $2\log u_{\epsilon}\in\mathfrak{sl}_{3}(\mathbb{Z})\setminus\{0\}$. Then one can deduce that there exists a constant $\tilde{C}>0$ depending only on $\operatorname{\mathbf{SL}}_{3}$ such that
$\|\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g)(2\log u_{\epsilon})\|_{\mathfrak{sl}_{3}}<\tilde{C}\epsilon e^{-\tau t_{\epsilon}}$
where $\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}$ is the adjoint representation of $\operatorname{\mathbf{SL}}_{3}$. This implies that $g\in S_{\tau}(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}},\{a_{t}\}_{t\in\mathbb{R}})^{c}$ by Definition 1.1.
Conversely, let $g\in\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$ such that $g\in S_{\tau}(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}},\{a_{t}\}_{t\in\mathbb{R}})^{c}$. Then by Definition 1.1, for any $\epsilon>0$, there exists $t_{\epsilon}>0$ such that
$\delta(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}\cdot g)\mathfrak{sl}_{3}(\mathbb{Z}))<\epsilon\cdot e^{-\tau t_{\epsilon}}.$
By [42, Proposition 3.3], for sufficiently small $\epsilon>0$, there exists a nilpotent element $n_{\epsilon}\in\mathfrak{sl}_{3}(\mathbb{Z})\setminus\{0\}$ such that
$\|\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}\cdot g)n_{\epsilon}\|_{\mathfrak{sl}_{3}}<\epsilon\cdot e^{-\tau t_{\epsilon}}$
where $\|\cdot\|_{\mathfrak{sl}_{3}}$ is a norm on the Lie algebra $\mathfrak{sl}_{3}(\mathbb{R})$ of $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$. Note that
$\exp(2n_{\epsilon})=I_{3}+(2n_{\epsilon})+(2n_{\epsilon})^{2}/2\in\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\setminus\{e\}.$
Then one can deduce that there exists a constant $\tilde{C}>0$ depending only on $\operatorname{\mathbf{SL}}_{3}$ such that
$d_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g\cdot\exp(2n_{\epsilon})g^{-1}a_{-t_{\epsilon}},e)<\tilde{C}\cdot\epsilon\cdot e^{-\tau t_{\epsilon}}.$
By [17, Definition 1.1], this implies that $g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ is not Diophantine of type $\tau$. Consequently, Theorem 1.7 and Remark 1.10 imply [17, Corollary 1.3] when $\mathbf{G}=\operatorname{\mathbf{SL}}_{3}$.

### v2

@@
-Jarník-Besicovitch type theorems for semisimple algebraic groups
+A shrinking target problem in homogeneous spaces of semisimple algebraic groups
 Abstract.
-In this note, we initiate a study on Jarník-Besicovitch type theorems for semisimple algebraic groups from the representation-theoretic point of view.
-Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible $\mathbb{Q}$-rational representation of a connected semisimple $\mathbb{Q}$-algebraic group $\mathbf{G}$ on a complex vector space $V$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in a $\mathbb{Q}$-split torus in $\mathbf{G}$.
-We define a subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ in terms of the representation $\rho$ and the subgroup $\{a_{t}\}_{t\in\mathbb{R}}$, and prove formulas for the Hausdorff dimension of the complement of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$.
-As corollaries, we deduce several Jarník-Besicovitch type theorems.
+In this paper, we study a shrinking target problem with target at infinity in a homogeneous space of a semisimple algebraic group from the representation-theoretic point of view.
+Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible $\mathbb{Q}$-rational representation of a connected semisimple $\mathbb{Q}$-algebraic group $\mathbf{G}$ on a complex vector space $V$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in a $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a positive function on $\mathbb{R}_{+}$.
+We define a subset $S_{\rho}(\psi)$ of $\psi$-Diophantine elements in $\mathbf{G}(\mathbb{R})$ in terms of the representation $\rho$ and $\{a_{t}\}_{t\in\mathbb{R}}$, and prove formulas for the Hausdorff dimension of the complement of $S_{\rho}(\psi)$.
+We also discuss the connections of our results to Diophantine approximation on flag varieties and rational approximation to linear subspaces in Grassmann varieties.
+Key words and phrases:
+shrinking target problem, Dani correspondence, structures of irreducible representations of semisimple algebraic groups, reduction theory of arithmetic subgroups, mixing of a semisimple flow, asymptotic estimates of the number of rational points, Diophantine approximation on flag varieties
 2020 Mathematics Subject Classification:
@@
 1. Introduction and main results
-1.1. Introduction
-In this note, we initiate a study on the Jarník-Besicovitch theorem in the metric theory of Diophantine approximation.
-Recall that a real number $x\in\mathbb{R}$ is Diophantine of type $\kappa>0$ if there exists a constant $C>0$ such that
-$\left|nx-m\right|\geq\frac{C}{n^{\kappa}}\;\left(\forall\frac{m}{n}\in\mathbb{Q}\right).$
-Then necessarily $x\in\mathbb{R}\setminus\mathbb{Q}$ and the Dirichlet theorem implies that $\kappa\geq 1$.
-Denote by $S_{\kappa}$ the set of all Diophantine numbers of type $\kappa$ and $S_{\kappa}^{c}$ its complement.
-Then one can deduce from the Khintchine’s theorem on metric Diophantine approximation that $S_{\kappa}^{c}$ is a set of full measure if $\kappa=1$, and is null if $\kappa>1$.
-A refined result due to Jarník and Besicovitch [4, 22] states further that the Hausdorff dimension of $S_{\kappa}^{c}$ is equal to $2/(\kappa+1)$ ($\kappa>1$).
-From the viewpoint of dynamical systems, the Jarník-Besicovitch theorem can be reformulated as a result of shrinking target problem, which was first observed by Hill and Velani [20].
-Let $f:X\to X$ be a map on a metric space $X$ with a measure $\mu$.
-Generally speaking, in the shrinking target problem, one studies the set $S$ of points in $X$ whose trajectories under $f$ approach to a shrinking target infinitely often with certain rate, and seeks to establish results about the size (the measure or the Hausdorff dimension) of the set $S$.
-The measure theoretic version of the shrinking target problem is usually related to the Khintchine’s theorem [26, 41] while the Hausdorff dimension version is linked with the Jarník-Besicovitch theorem [20].
-In homogeneous dynamics, various Diophantine sets can also be studied by the dynamical and ergodic properties of group actions on homogeneous spaces via Dani’s correspondence [14, 26].
-Specifically, let $X_{m+n}=\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})$,
-$a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})\quad(t\in\mathbb{R})$
-[…]
-$g\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})\to g\cdot\mathbb{Z}^{m+n}\quad(g\in\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})).$
-Let $\delta:X_{m+n}\to\mathbb{R}_{+}$ be the systole function on $X_{m+n}$ defined by
-$\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|$
-which is also called the first minimum of the lattice $\Lambda\in X_{m+n}$. For any $m\times n$ matrix $A$, denote by
-$u_{A}:=\left(\begin{array}[]{ccc}I_{m}&A\\ O&I_{n}\end{array}\right)\in N_{+}.$
-Then it is well-known via Mahler’s compactness criterion and Dani’s correspondence [14, 26] that $A$ is badly approximable if and only if the orbit $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ is bounded in $X_{m+n}$, i.e.
-$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\geq c\quad(t\geq 0)$
-for some constant $c>0$; singular if and only if $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ diverges i.e.
-$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\to 0\quad(\textup{as }t\to\infty);$
-and very well approximable if and only if
-$\limsup_{t\to\infty}\frac{-\log(\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n}))}{t}>0.$
-We say that an $m\times n$ matrix $A$ is Diophantine of type $\kappa$ if there exists a constant $C>0$ such that
-$\|Aq-p\|\geq C\|q\|^{-\kappa}$
-for any $q\in\mathbb{Z}^{n}\setminus\{0\}$ and $p\in\mathbb{Z}^{m}$.
-Then necessarily $\kappa\geq n/m$, and again using Dani’s correspondence, one can deduce that $A$ is Diophantine of type $\kappa$ if and only if there exists a constant $C>0$ such that
-$\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{m+n}))\geq Ce^{-\tau t}\;(\forall t>0)$
-where
-$\tau=\frac{m\kappa-n}{mn(\kappa+1)}.$
-The set $S_{\kappa}$ of all Diophantine numbers of type $\kappa$ defined at the beginning of the introduction then corresponds to the case $m=n=1$ here.
-In particular, one can rephrase the Jarník-Besicovitch theorem for the homogeneous space $X_{m+n}=X_{2}$ $(m=n=1)$, and get that for any parameter $10\leq\tau<$ 1, the Hausdorff dimension of the complement of the set
-$\{u_{A}\in N_{+}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{2}))\geq Ce^{-\tau t}\;(\forall t>0)\}$
-is equal to $1-\tau$.
-Using the decomposition of $X_{2}$ into the stable, central and unstable submanifolds of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $X_{2}$, one can then obtain the following equivalent statement: the Hausdorff dimension of the complement of the set
-$\{p\in X_{2}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
-is equal to $3-\tau$.
-In general, the matrix version of the Jarnik-Besicovitch theorem has also been established [9, 15], and it is equivalent to the statement that for any $0\leq\tau<1/n$, the Hausdorff dimension of the complement of the set
-$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
-is equal to $mn(1-\tau)+m^{2}+n^{2}+mn-1$.
-We remark that one can further consider the case where $\{a_{t}\}_{t\in\mathbb{R}}$ is a generic one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})$ and study the Hausdorff dimension of the complement of the set
-$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}.$
-This type of question leads to a general weighted multidimensional Jarník-Besicovitch theorem on metric Diophantine approximation, which is one of the main topics we pursue in this note (See Theorem 1.6).
-[…]
+1.1. Introduction and main problem
+In this paper, we study a shrinking target problem, the prototype of which was first proposed by Hill and Velani [39].
+Let $f:X\to X$ (or $f_{t}:X\to X$) be a map (or a flow) on a metric space $X$ with a measure $\mu$.
+Generally speaking, in the shrinking target problem, one studies the set $S$ of points in $X$ whose orbits under $f$ (or $f_{t}$) hit a shrinking target infinitely often, and seeks to establish results about the size (the $\mu$-measure or the Hausdorff dimension) of $S$.
+The question about the measure of $S$, especially when $f_{t}$ is a flow on a homogeneous space, is closely related to Khintchine-type theorems in the metric theory of Diophantine approximation [47, 48, 35], and has been studied extensively in the past few decades (e.g.
+[2, 32, 45, 46, 50, 54, 77, 79, 53, 5]).
+In particular, the shrinking target problem of this type for a diagonalizable homogeneous flow $f_{t}$ has been addressed in full generality by Kleinbock and Margulis [50] where a null-conull law on $\mu(S)$ and a logarithm law are established.
+On the other hand, the study of the Hausdorff dimension of $S$ for a homogeneous flow $f_{t}$ (when $\mu(S)=0$) is a finer problem, and it is usually related to the Jarník-Besicovitch theorem in Diophantine approximation [6, 42].
+As the analysis of the Hausdorff dimension of $S$ is more delicate and requires more information about the geometric structure of the space $X$ and the dynamics of the homogeneous flow $f_{t}$, only a few cases were known (which we will discuss below in this subsection).
+In this paper, we focus on this Hausdorff dimension version of shrinking target problem for homogeneous flows.
+(For discussions about other maps or flows, one may refer to e.g.
+[23, 40, 58, 70, 74, 81, 57].)
+In [82, 25], we consider a shrinking target problem with target at infinity for the homogeneous flow $a_{t}:G/\Gamma\to G/\Gamma$ where $G/\Gamma$ is a finite-volume quotient of a rank-one simple Lie group $G$ [82] or $G/\Gamma=\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ [25], and $\{a_{t}\}_{t\in\mathbb{R}}$ is a one-parameter diagonalizable subgroup in $G$.
+In particular, we obtain a formula for the Hausdorff dimension of the set $S$ as defined above, and establish a Jarník-Besicovitch theorem on Diophantine approximation in Heisenberg groups [82].
+One may also refer to [38, 36, 37, 62] for related discussions.
+This paper is a continuation of the works [82, 25], and we aim to generalize the main results in [82, 25] for diagonalizable homogeneous flows on finite-volume quotients of semisimple algebraic groups.
+We will see that our generalization has close connections to Diophantine approximation on flag varieties [56, 17] and rational approximation to linear subspaces in Grassmann varieties [71, 16], and it can imply certain Jarník-Besicovitch type theorems about well approximable subsets.
+Currently there are references [11, 20, 15, 75] which may be related to the topic of this paper with different emphases.
+In [11, 20], the main results are Jarník-Besicovitch theorems about Diophantine matrices, but with the help of a generalized Dani correspondence developed in [14, 50], these results can be reformulated as a shrinking target problem for the homogeneous flow $h_{t}:X_{m+n}\to X_{m+n}$ where $X_{m+n}=\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})$ and
+$h_{t}=\operatorname{diag}(\underbrace{e^{t/m},\dots,e^{t/m}}_{m\textup{ times}},\underbrace{e^{-t/n},\dots,e^{-t/n}}_{n\textup{ times}})\quad(t\in\mathbb{R}).$
+In [15], a variational principle is established to analyze $h_{t}$-orbits with various behaviors in $X_{m+n}$, which makes it possible to study the Hausdorff dimensions of a variety of Diophantine subsets in the space of $m\times n$ matrices.
+This variational principle is generalized in [75] for any diagonalizable flow on $X_{m+n}$ with respect to a nonstandard Hausdorff dimension.
+We remark that the results in [11, 20, 15, 75] mainly deal with diagonalizable flows on $X_{m+n}$.
+Now we propose the main problem of this paper.
+Let $\mathbf{G}$ be a semisimple algebraic group defined over $\mathbb{Q}$ and $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ a finite-dimensional irreducible representation of $\mathbf{G}$ over $\mathbb{Q}$ on a complex vector space $V$ with a $\mathbb{Q}$-structure.
+The vector space $V$ may be identified with $\mathbb{C}^{d}$ ($d=\dim_{\mathbb{C}}V$) equipped with a norm $\|\cdot\|$ so that $\mathbb{Z}^{d}\subset\mathbb{C}^{d}$ is compatible with the $\mathbb{Q}$-structure in $V$.
+For any discrete subgroup $\Lambda$ in $V$, define the first minimum of $\Lambda$ by
 $\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|.$
-Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter Ad-diagonalizable subgroup in $\mathbf{G}(\mathbb{R})$.
-Then the main object we would like to study is the Hausdorff dimension of the complement of the set
-$\{g\in\mathbf{G}(\mathbb{R}):\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for some }C>0\}.$
-Clearly the discussions above about the Jarnik-Besicovitch theorem with the homogeneous flow $a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})$ on $X_{m+n}$ can be reduced to this setting if we let $\mathbf{G}=\operatorname{\mathbf{SL}}_{m+n}$ and $\rho$ the standard representation of $\operatorname{\mathbf{SL}}_{m+n}$ on $\mathbb{C}^{m+n}$.
-Moreover, we will see later (Theorem 1.8) that by considering representations of algebraic groups, one will be able to establish some Jarnik-Besicovitch type theorems for certain algebraic subvarieties in an affine space.
-We remark that the question we study in this note starts from the sparse equidistribution problem, and a related shrinking target problem plays an important role.
-For more details, one can read [17, 46, 47, 48].
+Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter Ad-diagonalizable subgroup in $\mathbf{G}(\mathbb{R})$, $\mathbb{R}_{+}$ the set of positive real numbers and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a positive function on $\mathbb{R}_{+}$.
+Then we aim to estimate the Hausdorff dimension of the complement of the set
+$\{g\in\mathbf{G}(\mathbb{R}):\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq C\cdot\psi(t)\;(\forall t>0)\textup{ for some }C>0\}\subset\mathbf{G}(\mathbb{R})$
+with respect to a standard Riemannian metric on $\mathbf{G}(\mathbb{R})$.
+By Mahler’s compactness criterion, the main problem above is clearly a shrinking target problem with target at infinity for the homogeneous flow $a_{t}:\mathbf{G}(\mathbb{R})/\Gamma\to\mathbf{G}(\mathbb{R})/\Gamma$, where $\Gamma$ is an arithmetic lattice in $\mathbf{G}(\mathbb{R})$ preserving $\mathbb{Z}^{d}$.
+Note that the results in [11, 20] can be translated in the setting of this problem if we let $\mathbf{G}=\operatorname{\mathbf{SL}}_{m+n}$, $\rho$ the standard representation of $\operatorname{\mathbf{SL}}_{m+n}$ on $\mathbb{C}^{m+n}$ and $\{a_{t}\}_{t\in\mathbb{R}}=\{h_{t}\}_{t\in\mathbb{R}}$.
+We will see later in our theorems that the main problem also includes the cases in [82, 25].
 1.2. Main results
-In the rest of this note, we address the question above under the following assumption.
-We assume that $\mathbf{G}$ is a connected semisimple $\mathbb{Q}$-algebraic group (in the Zariski topology) and $\mathbf{T}$ is a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$.
-Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$, and $\rho$ a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$.
-Let $d=\dim V$ and we may identify $V\cong\mathbb{C}^{d}$ so that $\mathbb{Z}^{d}\subset\mathbb{C}^{d}$ is compatible with the $\mathbb{Q}$-structure in $V$.
+In this paper, we address the main problem above under the following assumption.
+We assume that $\mathbf{G}$ is a connected semisimple algebraic group over $\mathbb{Q}$ (with respect to the Zariski topology) and $\mathbf{T}$ is a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$.
+Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$, $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a positive function on $\mathbb{R}_{+}$ and $\rho$ a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ with $\dim\ker\rho=0$.
 Definition 1.1.
-An element $g\in\mathbf{G}(\mathbb{R})$ is Diophantine of type $\tau\geq 0$ if there exists a constant $C>0$ such that
-$\delta(\rho(a_{t}\cdot g)\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for any }t>0$
-where $\delta$ is the first minimum function.
-We denote by $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ the set of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ and $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$ its complement in $\mathbf{G}(\mathbb{R})$.
-If the representation $\rho$ and the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ are clearly stated in the contexts, we will simply write $S_{\tau}$ instead of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$.
+An element $g\in\mathbf{G}(\mathbb{R})$ is called $\psi$-Diophantine if there exists a constant $C>0$ such that
+$\delta(\rho(a_{t}\cdot g)\mathbb{Z}^{d})\geq C\cdot\psi(t)\textup{ for any }t>0$
+where $\delta$ is the first minimum function. We denote by $S_{\rho}(\psi)$ the set of all $\psi$-Diophantine elements, and $S_{\rho}(\psi)^{c}$ its complement in $\mathbf{G}(\mathbb{R})$.
 Remark 1.2.
-If $\mathbf{G}$ is $\mathbb{Q}$-anisotropic, then $\mathbf{T}=\{e\}$. So in the following, we assume that $\mathbf{G}$ is $\mathbb{Q}$-isotropic.
-To state the first main theorem about $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$, we need to introduce some notation.
-We choose a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$.
-Then $\mathbf{P}_{0}$ and $\mathbf{T}$ defines a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$.
-Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ defined by $\Phi\setminus\Phi^{+}$.
-Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$.
+In the rest of the paper, we assume that $\mathbf{G}$ is $\mathbb{Q}$-isotropic since otherwise $\mathbf{T}=\{e\}$.
+Note that for any $g\in\mathbf{G}(\mathbb{R})$, $\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d}$ is a unimodular lattice in the vector space of real points in $V$, and $\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})$ is bounded above by a constant depending only on $\rho$.
+Due to this fact, we may also assume that the function $\psi$ is bounded since otherwise $S_{\rho}(\psi)=\emptyset$ by Definition 1.1.
+We denote by $\dim_{H}(S)$ the Hausdorff dimension of a subset $S$ in a smooth manifold $\mathcal{M}$ with respect to a Riemannian metric on $\mathcal{M}$.
+To state the first two theorems about $\dim_{H}S_{\rho}(\psi)^{c}$, we need to introduce some notation.
+We choose a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$ with the Levi subgroup $Z(\mathbf{T})$ (the centralizer of $\mathbf{T}$ in $\mathbf{G}$).
+Then $\mathbf{P}_{0}$ and $\mathbf{T}$ define a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$.
+Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ defined by $\Phi\setminus\Phi^{+}$ with the same Levi subgroup $Z(\mathbf{T})$.
+Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$.
 One can write the space $V$ in the representation $\rho$ as a direct sum of weight spaces with respect to the action of $\mathbf{T}$
 $V=\bigoplus_{\beta}V_{\beta}.$
-By the structure theory of irreducible representations of complex semisimple groups and semisimple Lie algebras, there is a highest weight $\beta_{0}$ among the weights $\beta$’s (where the order is determined by the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$) and we denote by $V_{\beta_{0}}$ its corresponding weight space.
+By the structure theory of irreducible representations of complex semisimple groups and semisimple Lie algebras, there is a highest weight $\beta_{0}$ among the weights $\beta$’s (where the order is determined by $\Phi^{+}$), and we denote by $V_{\beta_{0}}$ its corresponding weight space.
 (See §2 for more details).
 The stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$ is a parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{\beta_{0}}$ containing $\mathbf{P}_{0}$, and its unipotent radical is denoted by $R_{u}(\mathbf{P}_{\beta_{0}})$.
-The opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ and its unipotent radical are denoted by $\overline{\mathbf{P}}_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ respectively.
-In the following, if an algebraic $\mathbb{Q}$-subgroup $\mathbf{F}\subset\mathbf{G}$ is normalized by $\mathbf{T}$, then we write $\Phi(\mathbf{F})$ for the set of $\mathbb{Q}$-roots in $\mathbf{F}$ relative to $\mathbf{T}$ while the summation $\sum_{\alpha\in\Phi(\mathbf{F})}$ (or $\prod_{\alpha\in\Phi(\mathbf{F})}$) means that we take the sum (or product) all over $\alpha\in\Phi(\mathbf{F})$ counted with multiplicities (i.e., the dimensions of the corresponding $\mathbb{Q}$-root spaces in the Lie algebra of $\mathbf{F}$).
-For a $\mathbb{Q}$-root or a $\mathbb{Q}$-weight $\lambda$ in $\rho$ relative to $\mathbf{T}$ (which is a $\mathbb{Q}$-character of $\mathbf{T}$) , we will often consider it as a linear functional on the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ and use the same symbol.
-In particular, we will write $\lambda(a_{t})$ ($t\in\mathbb{R}$) for the values of $\lambda$ (as a linear functional) on the Lie algebra of $\{a_{t}\}_{t\in\mathbb{R}}$ (so that $\lambda(a_{t})$ is linear on $t\in\mathbb{R}$).
-We denote by $\nu_{0}$ the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ relative to $\mathbf{T}$ such that
+It is known that there exists a $\mathbb{Q}$-torus $\mathbf{T}_{\beta_{0}}$ in $\mathbf{T}$ such that $Z(\mathbf{T}_{\beta_{0}})$ (the centralizer of $\mathbf{T}_{\beta_{0}}$ in $\mathbf{G}$) is a Levi subgroup of $\mathbf{P}_{\beta_{0}}$.
+We denote by $\overline{\mathbf{P}}_{\beta_{0}}$ the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ with the same Levi subgroup $Z(\mathbf{T}_{\beta_{0}})$.
+The unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$ is denoted by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$.
+In the following, if an algebraic $\mathbb{Q}$-subgroup $\mathbf{F}\subset\mathbf{G}$ is normalized by $\mathbf{T}$, then we write $\Phi(\mathbf{F})$ for the set of $\mathbb{Q}$-roots in $\mathbf{F}$ relative to $\mathbf{T}$, while the symbol $\sum_{\alpha\in\Phi(\mathbf{F})}$ (or $\prod_{\alpha\in\Phi(\mathbf{F})}$) stands for the sum (or product) over all the $\mathbb{Q}$-roots $\alpha\in\Phi(\mathbf{F})$ counted with multiplicities (i.e., the dimensions of the corresponding $\mathbb{Q}$-root spaces associated to $\alpha\in\Phi(\mathbf{F})$ in the Lie algebra of $\mathbf{F}$).
+For a $\mathbb{Q}$-root $\lambda$ in $\mathbf{G}$ or a $\mathbb{Q}$-weight $\lambda$ in $\rho$ relative to $\mathbf{T}$ (which is a $\mathbb{Q}$-character of $\mathbf{T}$), we will often consider it as a linear functional on the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ and use the same symbol.
+In particular, we will write $\lambda(a_{t})$ ($t\in\mathbb{R}$) for the values of $\lambda$ (as a linear functional) on the Lie algebra of $\{a_{t}\}_{t\in\mathbb{R}}$ (so that $\lambda(a_{t})$ is linear on the parameter $t\in\mathbb{R}$).
+We denote by $\nu_{0}$ any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $R_{u}(\overline{\mathbf{P}}_{0})$ satisfying the condition that
 $\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
-Now we can state the first main theorem in this note.
+Since the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, the value $\nu_{0}(a_{1})>0$.
+Now we can state our first two theorems. Here for any $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, we define
+$\tau(\psi):=\liminf_{t\to\infty}\left(-\frac{\ln(\psi(t))}{t}\right)$
+which will be an important quantity in the formula of the Hausdorff dimension of $S_{\rho}(\psi)^{c}$. We prove that the effective range of $\tau(\psi)$ is $0\leq\tau(\psi)\leq\beta_{0}(a_{-1})$.
+Theorem 1.3.
+We have $\tau(\psi)\geq 0$ and $\beta_{0}(a_{-1})>0$. If $\tau(\psi)>\beta_{0}(a_{-1})$, then $S_{\rho}(\psi)^{c}=\emptyset$.
 Theorem 1.4.
 Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$.
-Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$.
+Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$, and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$.
 Then
-$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\geq$ $\dim\mathbf{G}-\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
-for any $0\leq\tau<\beta_{0}(a_{-1})$.
+$\dim_{H}S_{\rho}(\psi)^{c}\geq$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
+for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$. In the case $\tau(\psi)=\beta_{0}(a_{-1})$ we have
+  1. (1)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
+  2. (2)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
+    $\dim_{H}S_{\rho}(\psi)^{c}\geq\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
 To state the next theorem, we need to introduce another notation.
-Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively.
+Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ in $\mathbf{G}$ respectively.
 Then the Weyl group relative to $\mathbb{Q}$ is defined by
 $_{\mathbb{Q}}{W}=N(\mathbf{T})/Z(\mathbf{T}).$
-Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above, and we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$.
-The Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
-$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
-which implies that
-$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
-Let $\{w_{i}\}_{i\in I}$ be a set of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$ where $I$ is the index set of ${}_{\mathbb{Q}}W$. For each $w_{i}$ $(i\in I)$, define
-$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
-and we have
-$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}\quad(i\in I).$
-Then one can further write the Bruhat decomposition as
-$\mathbf{G}(\mathbb{Q})=$ $\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q})$
-$=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
-$=$ $\bigcup_{i\in I}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
-Note that the sets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$ $(i\in I)$ in the decomposition of $\mathbf{G}(\mathbb{Q})$ above may overlap, and for our purpose, we can choose any subset $\bar{I}$ of $I$ (as small as possible) such that
-$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
-In the statement of Theorem 1.4 below, we will fix any such subset $\bar{I}$ in $I$.
-Let $\bigwedge^{\dim V_{\beta_{0}}}V$ be the $\dim V_{\beta_{0}}$-exterior product vector space of $V$ over $\mathbb{C}$, and $\rho_{\beta_{0}}$ the natural extension of $\rho$ on $\bigwedge^{\dim V_{\beta_{0}}}V$.
+Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ defined as in Theorem 1.4 such that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$.
+It is known that the Bruhat decomposition of $\mathbf{G}$ is the following [8, Theorem 21.15]
+$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q}).$
+Let $_{\mathbb{Q}}\mathcal{W}$} be a set of representatives of $_{\mathbb{Q}}{W}$ in $N(\mathbf{T})(\mathbb{Q})$.
+We may assume that $e\in{{}_{\mathbb{Q}}\mathcal{W}}$.
+Then one can deduce from the Bruhat decomposition that (see §2)
+$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{\mathcal{W}}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{\mathcal{W}}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
+For each $w\in{{}_{\mathbb{Q}}\mathcal{W}}$, define
+$\mathbf{F}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad\mathbf{H}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}\mathbf{P}_{\beta_{0}}w$
+and by [8, Proposition 21.9], we have
+$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w}\cdot\mathbf{F}_{w}.$
+It follows that
+$\mathbf{G}(\mathbb{Q})=$ $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{\mathcal{W}}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q})$
+$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot(w\mathbf{H}_{w}w^{-1})(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$
+$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
+Note that the subsets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$ $(w\in{{}_{\mathbb{Q}}\mathcal{W}})$ above may overlap, and for our purpose, we choose a subset $_{\mathbb{Q}}\overline{\mathcal{W}}$} of $_{\mathbb{Q}}\mathcal{W}$} (as small as possible) such that
+$\mathbf{G}(\mathbb{Q})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
+In the statement of Theorem 1.5 below, we fix any such subset $_{\mathbb{Q}}\overline{\mathcal{W}}$} of $_{\mathbb{Q}}\mathcal{W}$}.
+Let $\bigwedge^{\dim V_{\beta_{0}}}V$ be the $\dim V_{\beta_{0}}$-exterior product space of $V$ over $\mathbb{C}$, and denote by $\rho_{\beta_{0}}$ the natural extension of $\rho$ on $\bigwedge^{\dim V_{\beta_{0}}}V$.
 Let
@@
 $e_{V_{\beta_{0}}}:=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}\in\bigwedge^{\dim V_{\beta_{0}}}V.$
-For each $i\in\bar{I}$, define the following morphism
-$\Psi_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
-Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\Psi_{w_{i}}$ is an isomorphism onto its image.
-We denote by $a_{w_{i}}$ the growth rate of the asymptotic volume estimate of the real variety $\textup{Im}(\Psi_{w_{i}})$, and by $A_{w_{i}}$ the growth rate of the number of rational points in $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})$ (See §6 Corollary 6.8 and equation ($\ast$6) for more details).
-Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
+For each $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, define the following morphism
+$\Psi_{w}:\mathbf{F}_{w}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w}(x)=\rho_{\beta_{0}}(wxw^{-1})\cdot e_{V_{\beta_{0}}}.$
+Note that $w\mathbf{F}_{w}w^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and the stabilizer of the weight space $V_{\beta_{0}}$ is $\mathbf{P}_{\beta_{0}}$.
+So $\Psi_{w}$ is an isomorphism onto its image.
+For any $R>0$, let
+$B_{R}=\left\{v\in\bigwedge^{\dim V_{\beta_{0}}}V:\|v\|\leq R\right\}$
+where $\|\cdot\|$ is a Euclidean norm on $\bigwedge^{\dim V_{\beta_{0}}}V$. We denote by $a_{w}$ the growth rate of the asymptotic volume estimate of the real variety $\textup{Im}(\Psi_{w})$
+$a_{w}=\lim_{R\to\infty}\frac{\log\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))}{\log R},$
+where $\mu_{\mathbf{F}_{w}}$ is the Haar measure on $\mathbf{F}_{w}(\mathbb{R})$, and define by
+$A_{w}=A_{w}(\Omega_{\beta_{0}}):=\limsup_{H\to\infty}\frac{\log|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq H\}|}{\log H}$
+the growth rate of the number of rational elements in $\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q})$ where $\Omega_{\beta_{0}}$ is a fixed open bounded subset in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ containing a fundamental domain of $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})\cap\Gamma)$ (See §6).
+The definition of the height function $0pt(\cdot)$ is given in §2 (Definition 2.5).
+As before, let $\nu_{0}$ be any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
 $\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
-Now we can state the second main theorem in this note.
 Theorem 1.5.
 Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$.
-Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$.
+Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$, and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$.
 Then
-$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$
-$\leq$ $\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{a_{w_{i}},A_{w_{i}}\}\cdot\dim V_{\beta_{0}}\Big{|}\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>\tau\right\}$
-for any $0\leq\tau<\beta_{0}(a_{-1})$.
+$\dim_{H}S_{\rho}(\psi)^{c}$
+$\leq$ $\max_{\begin{subarray}{c}w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}\\ \beta_{0}(wa_{-1}w^{-1})\\ \geq\tau(\psi)\end{subarray}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{w},A_{w}\}\cdot\dim V_{\beta_{0}}\right\}$
+for any $0\leq\tau(\psi)\leq\beta_{0}(a_{-1})$.
+Combining Theorem 1.4, Theorem 1.5 and [63, Theorem 4], we obtain the following theorem.
+Theorem 1.6.
+Let $\mathbf{G}$ be a connected $\mathbb{Q}$-simple group, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$.
+Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$, and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$.
+Let $\beta_{0}$ and $\nu_{0}$ be defined as in Theorems 1.4 and 1.5.
+Suppose that the highest weight $\beta_{0}$ is a multiple of $\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}$.
+Then
+$\dim_{H}S_{\rho}(\psi)^{c}=$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
+for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$. In the case $\tau(\psi)=\beta_{0}(a_{-1})$, we have
+  1. (1)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
+  2. (2)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
+    $\dim_{H}S_{\rho}(\psi)^{c}=\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
+Any irreducible $\mathbb{Q}$-rational representation of a $\mathbb{Q}$-rank one $\mathbb{Q}$-simple group satisfies the conditions of Theorem 1.6.
+In this case, Theorem 1.6 recovers and largely extends [82, Theorem 1.1] in the arithmetic setting.
+We will discuss the relation between Theorem 1.6 and [82, Theorem 1.1] in §7.
+We remark that in [82], we also deal with the non-arithmetic case due to the fact that there are non-arithmetic lattices in $\mathbb{R}$-rank one simple groups.
+Besides $\mathbb{Q}$-rank one $\mathbb{Q}$-simple groups, in the following we also list some other examples of Theorem 1.6.
+The first example concerns the standard representation of $\operatorname{\mathbf{SL}}_{n}$ and any one-parameter diagonal subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$.
+It can be considered as a generalization of [11, 20] from the viewpoint of shrinking target problem.
+Theorem 1.7.
+Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the standard representation of $\operatorname{\mathbf{SL}}_{n}$ on $V=\mathbb{C}^{n}$ defined via matrix multiplication
+$\rho(g)\cdot v=g\cdot v\quad(g\in\operatorname{\mathbf{SL}}_{n},v\in V).$
+Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ and $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n}$ defined as in Theorem 1.6.
+Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a function on $\mathbb{R}_{+}$.
+Then we have
+$\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{n\cdot\tau(\psi)}{\nu_{0}(a_{1})}$
+for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$. In the case $\tau(\psi)=\beta_{0}(a_{-1})$, we have
+  1. (1)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
+  2. (2)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
+    $\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{n\cdot\tau(\psi)}{\nu_{0}(a_{1})}.$
+The second example deals with the adjoint representation of $\operatorname{\mathbf{SL}}_{n}$. We will prove in §7 that it generalizes the main result of [25].
+Theorem 1.8.
+Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the adjoint representation of $\operatorname{\mathbf{SL}}_{n}$ where $V=\mathfrak{sl}_{n}$ is the Lie algebra of $\operatorname{\mathbf{SL}}_{n}$.
+Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n}$ defined as in Theorem 1.6 and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$.
+Then we have
+$\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{(n-1)\cdot\tau(\psi)}{\nu_{0}(a_{1})}$
+for any $0\leq\tau(\psi)<\nu_{0}(a_{1})$. In the case $\tau(\psi)=\nu_{0}(a_{1})$, we have
+  1. (1)
+    If $\psi(t)\cdot e^{\nu_{0}(a_{1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
+  2. (2)
+    If $\psi(t)\cdot e^{\nu_{0}(a_{1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
+    $\dim_{H}S_{\rho}(\psi)^{c}=\dim\operatorname{\mathbf{SL}}_{n}-\frac{(n-1)\cdot\tau(\psi)}{\nu_{0}(a_{1})}.$
 Remark 1.9.
-It follows from the proof of Theorem 1.4 that for any $\tau\geq\beta_{0}(a_{-1})$, the subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=\emptyset$. (See §6.)
-As corollaries of Theorems 1.3 and 1.4, we obtain the following results.
-The first result generalizes the Jarnik-Besicovitch type theorem on the homogneous space $X_{n}=\operatorname{\mathbf{SL}}_{n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{n}(\mathbb{Z})$ for any one-parameter diagonal flow.
+One can see from the proofs that all the theorems stated above also hold for $\dim_{H}(S_{\rho}(\psi)^{c}\cap U)$ where $U$ is any open bounded subset in $\mathbf{G}(\mathbb{R})$. (See Remark 5.5.)
+1.3. Connections to Diophantine approximation
+In this subsection, we discuss two connections of Theorem 1.6 to the metric theory of Diophantine approximation on certain algebraic varieties, and present some related Jarník-Besicovitch type theorems.
+1.3.1. Connection to Diophantine approximation on flag varieties
+Diophantine approximation on group varieties is an open problem raised by Lang [56].
+In recent years, many important results have been established about Diophantine approximation on various group varieties including spheres, quadrics and algebraic groups [28, 29, 22, 31, 1, 18, 51, 21, 52, 30, 27].
+Here we discuss a connection between Theorem 1.6 and Diophantine approximation on generalized flag varieties, and derive an analogue of Jarník-Besicovitch theorem.
+We mainly follow the exposition in [17] by de Saxcé.
+For more details about Diophantine approximation on flag varieties, one may refer to [17].
+Let $\mathbf{X}=\mathbf{P}\backslash\mathbf{G}$ be a flag variety where $\mathbf{G}$ is a connected $\mathbb{Q}$-simple group and $\mathbf{P}$ a parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$.
+Let $V$ be a finite dimensional vector space defined over $\mathbb{Q}$, $\|\cdot\|$ a Euclidean norm on $V$ and $(e_{i})_{1\leq i\leq d}$ a rational orthonormal basis of $V$.
+For any rational point $v$ in $\mathbf{P}(V)(\mathbb{Q})$, we define the height $H(v)$ of $v$ by
+$H(v)=\|\mathbf{v}\|$
+where $\mathbf{v}$ is the representative primitive integral vector (up to sign) of $v$ in $\oplus_{1\leq i\leq d}\mathbb{Z}e_{i}$.
+Note that the height function $H(\cdot)$ is defined up to a multiplicative constant.
+If $\rho_{\chi}:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ is an irreducible $\mathbb{Q}$-rational representation of $\mathbf{G}$ generated by a one-dimensional weight space $V_{\chi}\subset V$ of highest weight $\chi$ such that the stabilizer of $V_{\chi}$ is $\mathbf{P}$, then we may identify $\mathbf{X}=\mathbf{P}\backslash\mathbf{G}$ with orbit $\mathbf{G}\cdot V_{\chi}$ in $\mathbf{P}(V)$ by
+$\mathbf{X}\to\mathbf{G}\cdot V_{\chi},\quad\mathbf{P}\cdot g\mapsto\rho_{\chi}(g^{-1})\cdot V_{\chi}\in\mathbf{G}\cdot V_{\chi}$
+and the height of any rational point in $\mathbf{X}=\mathbf{P}\backslash\mathbf{G}$ can be defined by restriction of the height function $H(\cdot)$ on $\mathbf{G}\cdot V_{\chi}\subset\mathbf{P}(V)$.
+As in Theorem 1.6, we may choose a maximal $\mathbb{Q}$-split torus $\mathbf{T}$ and a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ such that $\mathbf{T}\subset\mathbf{P}_{0}\subset\mathbf{P}$.
+Then $\mathbf{P}_{0}$ and $\mathbf{T}$ define a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$.
+We denote by $R_{u}(\mathbf{P})$ (resp.
+$R_{u}(\mathbf{P}_{0})$) the unipotent radical of $\mathbf{P}$ (resp.
+$\mathbf{P}_{0}$), and write $\Phi(R_{u}(\mathbf{P}))$ for the set of $\mathbb{Q}$-roots in $R_{u}(\mathbf{P})$ relative to $\mathbf{T}$.
+The symbol $\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}$ stands for the sum over all $\mathbb{Q}$-roots $\alpha\in\Phi(R_{u}(\mathbf{P}))$ counted with multiplicities (i.e., the dimensions of the $\mathbb{Q}$-root spaces associated to $\alpha\in\Phi(R_{u}(\mathbf{P}))$ in the Lie algebra of $R_{u}(\mathbf{P})$).
+In order to apply Theorem 1.6, we choose a one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{T}(\mathbb{R})$ as follows: $a_{t}=\exp(t\cdot Y)$ where $Y$ is an element in the Lie algebra of $\mathbf{T}(\mathbb{R})$ satisfying
+$\begin{cases}\alpha(Y)=\alpha(a_{1})=0,\textup{ if }\alpha\in\Delta_{\mathbf{P}}\\ \alpha(Y)=\alpha(a_{1})=-1,\textup{ if }\alpha\in\Delta\setminus\Delta_{\mathbf{P}}\end{cases}.$
+Here $\Delta_{\mathbf{P}}\subset\Delta$ is the set of simple $\mathbb{Q}$-roots associated to $\mathbf{P}$ .
+We denote by $\nu_{0}$ any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $\Phi^{+}$ satisfying
+$\nu_{0}(a_{-1})=\max\{\alpha(a_{-1}):\alpha\in\Phi^{+}\}.$
+On the variety $\mathbf{X}(\mathbb{R})$, there is a Carnot-Carathéodory distance $d_{\textup{CC}}$, and its explicit definition can be found in [17, Chapter 2, §2].
+With this Carnot-Carathéodory distance, for any $x\in\mathbf{X}(\mathbb{R})$, one may define the Diophantine exponent $\beta_{\chi}(x)$ of $x$ by
+$\beta_{\chi}(x)=\inf\{\beta>0:\exists\textup{$C>0$ such that }d_{\textup{CC}}(x,v)\geq C\cdot H(v)^{-\beta}\;(\forall v\in\mathbf{X}(\mathbb{Q}))\}.$
+We have the following theorem about the Diophantine exponent $\beta_{\chi}(x)$.
+Theorem 1.10 ([17, Theorem 2.4.5]).
+There exits a positive constant $\beta_{\chi}>0$ such that for almost every $x\in\mathbf{X}(\mathbb{R})$ (with respect to a Riemannian volume form on $\mathbf{X}(\mathbb{R})$)
+$\beta_{\chi}(x)=\beta_{\chi}.$
+Moreover, we have $\beta_{\chi}=-1/\chi(Y)=1/\chi(a_{-1})$.
+Definition 1.11.
+For any decreasing function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the set of $\psi$-Diophantine points in $\mathbf{X}(\mathbb{R})$ by
+$E_{\mathbf{X}}(\psi)=\{x\in\mathbf{X}(\mathbb{R}):\exists C>0\textup{ such that }d_{\textup{CC}}(x,v)\geq C\cdot(H(v))^{-\beta_{\chi}}\psi(H(v))\;(\forall v\in\mathbf{X}(\mathbb{Q}))\}$
+where the constant $\beta_{\chi}$ is defined as in Theorem 1.10. We denote by $E_{\mathbf{X}}(\psi)^{c}$ the complement of $E_{\mathbf{X}}(\psi)$ in $\mathbf{X}(\mathbb{R})$.
+For any function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the lower order at infinity of $\psi$ by
+$\gamma(\psi)=\liminf_{t\to\infty}\left(-\frac{\ln\psi(t)}{\ln t}\right).$
+From a correspondence theorem [17, Proposition 3.2.4] and Theorem 1.6, we may deduce the following result.
 Theorem 1.12.
-Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the standard representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ on the complex vector space $V=\mathbb{C}^{n}$ defined by
-$\rho(g)\cdot v=g\cdot v\quad(g\in\operatorname{\mathbf{SL}}_{n},v\in V)$
-via matrix multiplication.
-Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4.
-Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
-$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
-The second result generalizes the main results in [17].
+Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a decreasing function.
+Then $0\leq\gamma(\psi)\leq\infty$.
+If the highest weight $\chi$ of $\rho_{\chi}$ is a multiple of $\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}{\alpha}$, then the Hausdorff dimension of $E_{\mathbf{X}}(\psi)^{c}$ (with respect to a standard Riemannian metric on $\mathbf{X}(\mathbb{R})$) equals
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}=$ $\dim\mathbf{X}-\left(\frac{1}{\beta_{\chi}}-\frac{1}{\gamma(\psi)+\beta_{\chi}}\right)\frac{1}{\chi(a_{-1})\nu_{0}(a_{-1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}\alpha(a_{-1})$
+for any $0\leq\gamma(\psi)\leq\infty$.
+1.3.2. Connection to rational approximation to linear subspaces
+Diophantine approximation on the Grassmann variety of $l$-dimensional subspaces in $\mathbb{R}^{n}$ by $k$-dimensional $\mathbb{Q}$-subspaces $(1\leq k\leq l<n)$ is a problem suggested by Schmidt in [71].
+It has attracted increasing attention and one may refer to e.g.
+[67, 19, 13, 43, 44, 26].
+Recently, de Saxcé has made a breakthrough in [16] which answers several Schmidt’s problems in [71].
+Here we discuss a connection between Theorem 1.6 and [16].
+We mainly follow the setup in [16].
+For any $1\leq l\leq n-1$ ($l\in\mathbb{N}$), let $X_{l}(\mathbb{R})$ be the Grassmann variety of $l$-dimensional linear $\mathbb{R}$-subspaces in $\mathbb{R}^{n}$.
+Let $X_{l}(\mathbb{Q})$ be the set of $\mathbb{Q}$-rational points in $X_{l}(\mathbb{R})$, i.e.
+the set of $l$-dimensional $\mathbb{Q}$-subspaces in $\mathbb{R}^{n}$.
+We define a distance $d(\cdot,\cdot)$ between subspaces in $\mathbb{R}^{n}$ as follows.
+Let $P_{1}$ and $P_{2}$ be two linear $\mathbb{R}$-subspaces.
+If $P_{1}=\mathbb{R}\cdot v_{1}$ and $P_{2}=\mathbb{R}\cdot v_{2}$ are two lines $(v_{1},v_{2}\in\mathbb{R}^{n}\setminus\{0\})$, then the distance between $P_{1}$ and $P_{2}$ is the usual distance on the projective space $\mathbf{P}\mathbb{R}^{n-1}$, i.e.
+$d(P_{1},P_{2})=\sin(\theta(P_{1},P_{2}))=\frac{\|v_{1}\wedge v_{2}\|}{\|v_{1}\|\|v_{2}\|}$
+where $\theta(P_{1},P_{2})$ is the angle between $P_{1}$ and $P_{2}$, and the norms are the Euclidean norms on $\mathbb{R}^{n}$ and $\bigwedge^{2}\mathbb{R}^{n}$ respectively.
+If $P_{1}$ is a line, then the distance between $P_{1}$ and $P_{2}$ is defined by
+$d(P_{1},P_{2})=\min\{d(P_{1},Q):Q\textup{ is a one-dimensional linear subspace in }P_{2}\}.$
+If $P_{1}$ and $P_{2}$ are two general subspaces in $\mathbb{R}^{n}$, then we define the distance between $P_{1}$ and $P_{2}$ by
+[…]
+For $1\leq k\leq n-1$, we define a height function $H(\cdot)$ on $X_{k}(\mathbb{Q})$ as follows.
+For any $\mathbb{Q}$-subspace $P\in X_{k}(\mathbb{Q})$, we choose an integral basis $\{v_{i}\}_{1\leq i\leq k}\subset\mathbb{Z}^{n}$ for $P$ and define the height of $P$ by
+$H(P)=\|v_{1}\wedge\cdots\wedge v_{k}\|$
+where the norm is the Euclidean norm on $\bigwedge^{k}\mathbb{R}^{n}$.
+Now let $1\leq k\leq l<n$.
+With the distance $d(\cdot,\cdot)$ and the height function $H(\cdot)$ as above, for any $x\in X_{l}(\mathbb{R})$, we may define the Diophantine exponent of $x$ for approximation by $k$-dimensional rational subspaces by
+$\beta_{k}(x)=\inf\{\beta>0:d(v,x)\geq C\cdot H(v)^{-\beta}\;(\forall v\in X_{k}(\mathbb{Q}))\textup{ for some }C>0\}.$
+Theorem 1.13 ([16, Theorem 1]).
+Let $1\leq k\leq l<n$. For any $x\in X_{l}(\mathbb{R})$, we have
+$\beta_{k}(x)\geq\frac{n}{k(n-l)}$
+and the equality holds for almost every point $x\in X_{l}(\mathbb{R})$ with respect to a Riemannian volume form on $X_{l}(\mathbb{R})$.
+Definition 1.14.
+For any decreasing function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the set of $\psi$-Diophantine subspaces in $X_{l}(\mathbb{R})$ for approximation by $k$-dimensional $\mathbb{Q}$-subspaces by
+$E_{l,k}(\psi)=\{x\in X_{l}(\mathbb{R}):d(v,x)\geq C\cdot H(v)^{-\frac{n}{k(n-l)}}\cdot\psi(H(v))\;(\forall v\in X_{k}(\mathbb{Q}))\textup{ for some }C>0\}.$
+We denote by $E_{l,k}(\psi)^{c}$ the complement of $E_{l,k}(\psi)$ in $X_{l}(\mathbb{R})$.
+As in §1.3.1, for any function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the lower order at infinity of $\psi$ by
+$\gamma(\psi)=\liminf_{t\to\infty}\left(-\frac{\ln\psi(t)}{\ln t}\right).$
+From a correspondence theorem [16, Proposition 1] and Theorem 1.6 with the representation $\rho_{k}:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ $(V=\bigwedge^{k}\mathbb{C}^{n})$ which is the $k$-th exterior product of the standard representation of $\operatorname{\mathbf{SL}}_{n}$, we may deduce the following result.
 Theorem 1.15.
-Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ where $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$, $V=\mathfrak{g}=\mathfrak{sl}_{n}$ is the Lie algebra of $\operatorname{\mathbf{SL}}_{n}$ and $\rho=\operatorname{Ad}$ is the adjoint representation of $\mathbf{G}$.
-Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\mathbf{G}$, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4.
-Then for any $0\leq\tau<\nu_{0}(a_{1})$, we have
-$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}.$
-We also deduce the following
-Theorem 1.8.
-Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible representation of $\operatorname{\mathbf{SL}}_{2}$ defined over $\mathbb{Q}$.
-Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{2}$ defined as in Theorem 1.4.
-Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
-$\dim_{H}S_{\tau}^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
-Remark 1.9.
-Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be the irreducible representation in Theorem 1.8 with $\dim V=n+1$ $(n\geq 1)$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ and let $\rho_{0}$ be the standard representation of $\operatorname{\mathbf{SL}}_{n+1}$ on $\mathbb{C}^{n+1}$.
-Note that $V\cong\mathbb{C}^{n+1}$ and $\rho(\operatorname{\mathbf{SL}}_{2})\subset\operatorname{\mathbf{SL}}_{n+1}\subset\operatorname{\mathbf{GL}}(V)$.
-Then by Theorem 1.8, we have
-$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}\cap\rho(\operatorname{\mathbf{SL}}_{2}(\mathbb{R}))=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}$
-where $\beta_{0}$ is the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3. On the other hand, by Theorem 1.6
-$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}=\dim\operatorname{\mathbf{SL}}_{n+1}-\frac{(n+1)\cdot\tau}{\nu_{0}(\rho(a_{1}))}$
-where $\nu_{0}$ is the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n+1}$ with respect to $\{\rho(a_{t})\}_{t\in\mathbb{R}}$ defined as in Theorem 1.4.
-This gives an example of Jarník-Besicovitch type theorems for some algebraic subvarieties in affine spaces.
-Remark 1.10.
-One can see from the arguments in this note that all the theorems stated above also work for any open bounded subset $U$ in $\mathbf{G}(\mathbb{R})$, i.e.
-$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\cap U=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$.
-(See Remark 5.6.)
-1.3. Discussions and open problems
-[…]
-$\liminf_{t\to\infty}-\frac{1}{t}\log\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})=\tau.$
-When $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and $\rho$ is the standard representation of $\mathbf{G}$, the Hausdorff dimension and the packing dimension of $\operatorname{Sing}(\tau)$ are discussed in [43] in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ is a singular one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, and the proofs rely on a variational principle in the parametric geometry of numbers.
-This variational principle is later generalized in [39] for any one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$.
-For the lower bound and the upper bound we obtain in Theorem 1.3 and Theorem 1.4, we have computed in several cases (e.g.
-Theorems 1.61.7 and 1.8) that these two bounds are equal.
-It would be interesting to compute these two bounds in other cases and determine the representations $\rho$ for which these two bounds are equal.
+Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a decreasing function, and $1\leq k\leq l<n$.
+Then $0\leq\gamma(\psi)\leq\infty$, and the Hausdorff dimension of $E_{l,k}(\psi)^{c}$ (with respect to a standard Riemannian metric on $X_{l}(\mathbb{R})$) equals
+$\dim_{H}E_{l,k}(\psi)^{c}=(l-k)(n-l)+\frac{n}{n/(k(n-l))+\gamma(\psi)}$
+for any $0\leq\gamma(\psi)\leq\infty$.
+1.4. Strategy of the proofs and organization of the paper
+The investigation of the shrinking target problem in $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ in [25] provides a strategy to address the main problem above.
+However, in this paper, since we are dealing with a shrinking target problem in a homogeneous space of a semisimple algebraic group $\mathbf{G}$ from the perspective of representations, the studies on the structure of the group $\mathbf{G}$ and the structure of the representation $\rho$ would be fundamental in our analysis.
+Due to the complexities of the structures of $\mathbf{G}$ and $\rho$, we have to upgrade the method in [25] considerably to tackle the main problem.
+Here we may reduce the problem to computing $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+We first define a notion of rational elements in $\mathbf{G}(\mathbb{R})$ according to the representation $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$, and mainly focus on the rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+The set of rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is usually not countable, but a collection of leaves in a foliation in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+To get a lower bound of the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, we would like to choose neighborhoods around these (leaves of) rational elements and construct a Cantor-type subset.
+In order to apply Hausdorff dimension formulas, at each level of the Cantor-type subset, the neighborhoods should be disjoint.
+In our case, it may happen that the neighborhoods at each level are not disjoint if we proceed in the usual way.
+It implies that there are surplus rational elements in the selected leafs, and we have to sieve them out so that the neighborhoods around the remaining rational elements are disjoint.
+It is also required that the proportion of the remaining rational elements is not small compared to the set of all rational elements in order to avoid any loss of Hausdorff dimension.
+For this purpose, we will need the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$.
+The mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$, together with the structure of the Siegel sets in $\mathbf{G}(\mathbb{R})$, not only helps us sieve out surplus rational elements, but also gives an asymptotic estimate about the measure of the subset of the remaining rational elements.
+Then we may apply the Hausdorff dimension formula of a Cantor-type set and get a lower bound for $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+To get an upper bound of $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, we will need the Bruhat decomposition of $\mathbf{G}$, or precisely the decomposition
+$\mathbf{G}(\mathbb{Q})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
+For any $p\in S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, from the structures of the representation $\rho$ and the Siegel sets in $\mathbf{G}(\mathbb{R})$, one may deduce that there exist a Weyl element $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, a sequence of rational elements $\{q_{k}\}_{k\in\mathbb{N}}$ and a sequence of subsets $\{E_{k}\}_{k\in\mathbb{N}}$ in $\mathbf{F}_{w}(\mathbb{R})$ (defined from the preimages of certain open subsets under the morphism $\Psi_{w}$ contracted by the action of $\{a_{t}\}_{t\in\mathbb{R}}$) such that
+$p\in E_{k}\cdot w^{-1}q_{k}\;(k\in\mathbb{N}).$
+So we can divide the set $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ into finitely many subsets $\mathcal{E}_{w,\epsilon}(\psi)$ (indexed by the elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$), each of which corresponds to a type of Diophantine approximation in the sense that any $p\in\mathcal{E}_{w,\epsilon}(\psi)$ is covered by infinitely many subsets of the form $E\cdot w^{-1}q$ described as above.
+Now fix $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$.
+As mentioned before, the set of rational elements $q$ is a collection of leaves in some foliation in $\mathbf{G}(\mathbb{R})$, and by the structure of the representation $\rho$ and the morphism $\Psi_{w}$, the shape of the subset $E$ defined above may be arbitrary and there may be rational elements $q$ far away from a point $p\in\mathcal{E}_{w,\epsilon}(\psi)$ but one still has $p\in E\cdot w^{-1}q$.
+So if we construct an open cover of $\mathcal{E}_{w,\epsilon}(\psi)$ directly from cutting all the subsets of the form $E\cdot w^{-1}q$ into cubes, there would be many unnecessary open cubes counted in the open cover.
+In order to obtain an upper bound of the Hausdorff dimension of $\mathcal{E}_{w,\epsilon}(\psi)$, we project the subset $\mathcal{E}_{w,\epsilon}(\psi)$ into the compact quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma)$, and construct an open cover for the projection of $\mathcal{E}_{w,\epsilon}(\psi)$ instead.
+The projections of the subsets of the form $E\cdot w^{-1}q$ overlap in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma)$, and by a careful analysis, one can show that the union of these projections can be described in a quite economical way, from which an efficient open cover can be constructed for the projection of $\mathcal{E}_{w,\epsilon}(\psi)$.
+Then by the countable stability of Hausdorff dimension, we may obtain an upper bound of $\dim_{H}\mathcal{E}_{w,\epsilon}(\psi)$ for each $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, and consequently an upper bound for $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+Note that in this way, we are able to avoid many computations in [25], and the only important information we need is about the growth rates of the volume estimates of certain real algebraic varieties and the growth rates of the numbers of rational points in these algebraic varieties, for which we may appeal to algebraic geometry (Cf.
+Manin’s conjecture [3]).
+The paper is organized as follows:
+  1. $\bullet$
+    In §2, we first give introductions to the reduction theory of arithmetic lattices and the theory of irreducible representations of complex semisimple groups and complex Lie algebras.
+Then we prove Theorem 1.3.
+After that, we define a notion of rational elements and discuss some properties of the rational elements which will be important in the subsequent sections.
+  2. $\bullet$
+    In §3, we assume that the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on any arithmetic quotient $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing, where $\mathbf{G}(\mathbb{R})^{0}$ is the identity component of $\mathbf{G}(\mathbb{R})$ (as a Lie group) and $\Gamma$ is a subgroup of finite index in $\mathbf{G}(\mathbb{Z})$.
+We select certain rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ according to some algebraic conditions, and apply the mixing property to obtain an asymptotic estimate about the measure of the subset of the selected rational elements.
+In §4, we construct a Cantor-type subset $\mathbf{A}_{\infty}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ from neighborhoods of these selected rational elements.
+With the asymptotic estimate result established in §3 and the condition that $0\leq\tau(\psi)<\beta_{0}(a_{-1})$, we compute the Hausdorff dimension of $\mathbf{A}_{\infty}$ and prove Theorem 1.4.
+  3. $\bullet$
+    In §5, we discuss the case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ but the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on any arithmetic quotient of $\mathbf{G}(\mathbb{R})^{0}$ is not mixing.
+Then there is a proper normal $\mathbb{Q}$-subgroup $\mathbf{H}$ of $\mathbf{G}$ such that $\mathbf{H}$ contains $\{a_{t}\}_{t\in\mathbb{R}}$ and the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on any arithmetic quotient of $\mathbf{H}(\mathbb{R})^{0}$(= the identity component of $\mathbf{H}(\mathbb{R})$) is mixing.
+We show that the arguments in §3 and §4 work equally well in the group $\mathbf{H}(\mathbb{R})$ and Theorem 1.4 still holds in this case.
+At the end of this section, we discuss the case $\tau(\psi)=\beta_{0}(a_{-1})$.
+This completes the proof of Theorem 1.4.
+  4. $\bullet$
+    In §6, we divide the subset $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ into finitely many subsets $\mathcal{E}_{w,\epsilon}(\psi)$ indexed by elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, each of which corresponds to a type of Diophantine approximation.
+We project the subsets $\mathcal{E}_{w,\epsilon}(\psi)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$ and construct open covers for the projections of $\mathcal{E}_{w,\epsilon}(\psi)$.
+This gives an upper bound of the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and proves Theorem 1.5.
+  5. $\bullet$
+    In §7, we derive Theorems 1.61.7 and 1.8 from Theorems 1.4 and  1.5. The proofs of Theorem 1.12 and Theorem 1.15 are given in §8.
 2. Preliminaries
-In this section, we list some preliminaries needed in this note. Then we give a definition of rational elements in $\mathbf{G}(\mathbb{R})$ and discuss some properties of rational elements.
-We first need the reduction theory of arithmetic subgroups of $\mathbf{G}(\mathbb{R})$ [5].
+In this section, we list some preliminaries needed in this paper and prove Theorem 1.3.
+Then we give a definition of rational elements in $\mathbf{G}(\mathbb{R})$ and discuss some properties of rational elements.
+Let $\mathbf{G}$, $\mathbf{T}$, $\{a_{t}\}_{t\in\mathbb{R}}$ and the representation $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ satisfy the assumption in §1.2.
+We first need the reduction theory of arithmetic subgroups of $\mathbf{G}(\mathbb{R})$ [9, 7].
 Let $K$ be a maximal compact subgroup in $\mathbf{G}(\mathbb{R})$ and $\Gamma$ an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})$.
-We fix a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$.
-Denote by $\Phi$ the set of $\mathbb{Q}$-roots with respect to $\mathbf{T}$, $\Phi^{+}$ the set of positive $\mathbb{Q}$-roots corresponding to the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ and $\Delta$ the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$.
+We fix a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$ with the Levi subgroup $Z(\mathbf{T})$ (the centralizer of $\mathbf{T}$ in $\mathbf{G}$).
+Denote by $\Phi$ the set of $\mathbb{Q}$-roots in $\mathbf{G}$ with respect to $\mathbf{T}$, $\Phi^{+}$ the set of positive $\mathbb{Q}$-roots corresponding to the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ and $\Delta$ the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$.
 For any parabolic $\mathbb{Q}$-subgroup $\mathbf{P}$, denote by $R_{u}(\mathbf{P})$ its unipotent radical.
-Let $M$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\mathbf{G}(\mathbb{R})}(\mathbf{T}(\mathbb{R}))$ (= the centralizer of $\mathbf{T}(\mathbb{R})$ in $\mathbf{G}(\mathbb{R})$).
+We write $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}=Z(\mathbf{T})$ is the Levi subgroup, and write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$ so that $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$.
+Let $M=\mathbf{M}_{a}(\mathbb{R})^{0}$ be the identity component of the Lie group $\mathbf{M}_{a}(\mathbb{R})$.
 For $\eta>0$, denote by
 $T_{\eta}=\{a\in\mathbf{T}(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\Delta\}.$
-A Siegel set in $\mathbf{G}(\mathbb{R})$ is a subset of the form $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ for some $\eta>0$ and a relatively compact open subset $\Omega$ containing identity in $M\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R})$, and the group $\mathbf{G}(\mathbb{R})$ can be written as
+A Siegel set in $\mathbf{G}(\mathbb{R})$ is a subset of the form $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ for some $\eta>0$ and some relatively compact open subset $\Omega$ containing identity in $M\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R})$, and the group $\mathbf{G}(\mathbb{R})$ can be written as
 $\mathbf{G}(\mathbb{R})=S_{\eta,\Omega}\cdot\mathcal{K}\cdot\Gamma$
-for some Siegel set $S_{\eta,\Omega}$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$. Moreover, the finite set $\mathcal{K}$ satisfies the property that
+for some Siegel set $S_{\eta,\Omega}$ and a finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$. Moreover, the finite set $\mathcal{K}$ satisfies the property that
 $\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
-where $\mathbf{P}_{0}$ is the minimal parabolic $\mathbb{Q}$-subgroup. Denote by
+where $\mathbf{P}_{0}$ is the minimal parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$. Denote by
 $\mathcal{K}=\{x_{1},x_{2},\dots,x_{k}\}=\{x_{j}\}_{j\in J}\subset\mathbf{G}(\mathbb{Q})$
 and we may assume that $e\in\mathcal{K}$.
-In what follows, we choose $\Gamma$ to be an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})\cap\mathbf{G}(\mathbb{R})^{0}$, where $\mathbf{G}(\mathbb{R})^{0}$ denotes the connected component of identity in the Lie group $\mathbf{G}(\mathbb{R})$, so that $\rho(\Gamma)$ preserves the lattice $\mathbb{Z}^{d}$ in $V$.
-Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$ and its unstable horospherical subgroup is contained in $R_{u}(\overline{\mathbf{P}}_{0})$ where $\overline{\mathbf{P}}_{0}$ is the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ determined by $\Phi\setminus\Phi^{+}$.
-We also write $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the centralizer of $\mathbf{T}$ in $\mathbf{G}$, and write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$ so that $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$.
-Now we choose a maximal $\mathbb{Q}$-torus $\mathbf{S}$ in $\mathbf{P}_{0}$ containing $\mathbf{T}$. The Lie algebra $\mathfrak{g}$ of $\mathbf{G}$ can be written as a direct sum
-$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Psi}\mathfrak{g}_{\alpha}$
-of root spaces relative to $\mathbf{S}$ via the adjoint representation of $\mathbf{G}$, where $\Psi$ denotes the set of roots in $\mathfrak{g}$ relative to $\mathbf{S}$.
-We may determine a set of positive roots in $\Psi$, which we denote by $\Psi^{+}$, such that
+In what follows, we choose $\Gamma$ to be an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})\cap\mathbf{G}(\mathbb{R})^{0}$, where $\mathbf{G}(\mathbb{R})^{0}$ denotes the identity component of the Lie group $\mathbf{G}(\mathbb{R})$, so that $\rho(\Gamma)$ preserves the lattice $\mathbb{Z}^{d}$ in $V$.
+Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$ and the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, where $\overline{\mathbf{P}}_{0}$ is the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ determined by $\Phi\setminus\Phi^{+}$ with the same Levi subgroup $\mathbf{M}_{0}=Z(\mathbf{T})$.
+Now we choose a maximal $\mathbb{Q}$-torus $\mathbf{S}$ in $\mathbf{P}_{0}$ containing $\mathbf{T}$.
+The Lie algebra $\mathfrak{g}$ of $\mathbf{G}$ can be written as a direct sum of root spaces relative to $\mathbf{S}$ via the adjoint representation of $\mathbf{G}$
+$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Psi}\mathfrak{g}_{\alpha},$
+where $\Psi$ is the set of roots in $\mathfrak{g}$ relative to $\mathbf{S}$. We may determine a set of positive roots in $\Psi$, which we denote by $\Psi^{+}$, such that
 $\operatorname{Lie}(R_{u}(\mathbf{P}_{0}))\subset\sum_{\alpha\in\Psi^{+}}\mathfrak{g}_{\alpha}\textup{ and }\operatorname{Lie}(R_{u}(\overline{\mathbf{P}}_{0}))\subset\sum_{\alpha\in\Psi\setminus\Psi^{+}}\mathfrak{g}_{\alpha}.$
+Here $\operatorname{Lie}(R_{u}(\mathbf{P}_{0}))$ and $\operatorname{Lie}(R_{u}(\overline{\mathbf{P}}_{0}))$ denote the Lie algebras of $R_{u}(\mathbf{P}_{0})$ and $R_{u}(\overline{\mathbf{P}}_{0})$ respectively.
 The set of simple roots in $\Psi^{+}$ is denoted by $\Pi$.
-It is known that there are complete classifications of irreducible representations of complex semisimple groups and Lie algebras [21, 27].
-Let $V=\bigoplus_{\lambda}V_{\lambda}$ be the decomposition of the vector space $V$ in $\rho$ into the direct sum of weight spaces $V_{\lambda}$ relative to $\mathbf{S}$.
+It is known that there are complete classifications of irreducible representations of complex semisimple groups and Lie algebras [41, 55]. Let $V$ be the vector space in the representation $\rho$, and
+$V=\bigoplus_{\lambda}V_{\lambda}$
+the decomposition of $V$ into the direct sum of weight spaces $V_{\lambda}$ relative to $\mathbf{S}$.
 According to the theorem of highest weight, there is a unique highest weight $\lambda_{0}$ among the weights $\lambda$’s (the order is determined by $(\Psi,\Psi^{+},\Pi)$) such that
@@
 Note that $\beta_{0}$ is defined over $\mathbb{Q}$ and $\beta_{0}=\lambda_{0}|_{\mathbf{T}}$.
-In particular, $\lambda_{0}(\mathbf{T}(\mathbb{R}))\subset\mathbb{R}$ and $\beta_{0}$ is the highest weight among $\beta$’s relative to $\mathbf{T}$ (the order is determined by the root system $(\Phi,\Phi^{+},\Delta)$) (as discussed in §1).
-Since the stable subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$, we have $\alpha(a_{t})\leq 0$ $(t\geq 0)$ for any $\alpha\in\Phi^{+}$.
-From property (3) of the representation $\rho$ above and the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, one can deduce that $\beta_{0}(a_{t})=\lambda_{0}(a_{t})\leq 0$.
-If $\beta_{0}(a_{t})=\lambda_{0}(a_{t})=0$, then by the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, for any weight $\lambda$ in $\rho$ relative to $\mathbf{S}$, we have $\lambda(a_{t})=0$ and $\{a_{t}\}_{t\in\mathbb{R}}\subset\ker\rho$, which contradicts the assumption that $\dim\ker\rho=0$.
-Therefore we have $\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0$ and
+In particular, $\lambda_{0}(\mathbf{T}(\mathbb{R}))\subset\mathbb{R}$ and $\beta_{0}$ is the highest weight among $\beta$’s relative to $\mathbf{T}$ where the order is determined by the root system $(\Phi,\Phi^{+},\Delta)$.
+Since the stable subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$, we have $\alpha(a_{t})\leq 0$ $(t>0)$ for any $\alpha\in\Phi^{+}$.
+From property (3) of the representation $\rho$ above and the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, one can deduce that $\beta_{0}(a_{t})=\lambda_{0}(a_{t})\leq 0$ $(t>0)$.
+If $\beta_{0}(a_{t})=\lambda_{0}(a_{t})=0$ $(t>0)$, then by the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$ and $\lambda_{0}$ is the highest weight relative to $\mathbf{S}$, for any other weight $\lambda$ in $\rho$ relative to $\mathbf{S}$, we also have $\lambda(a_{t})=0$ $(t>0)$.
+This implies that $\{a_{t}\}_{t\in\mathbb{R}}\subset\ker\rho$, which contradicts the assumption that $\dim\ker\rho=0$.
+Therefore we have $\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0$ $(t>0)$.
+Note that
 $\rho(a_{t})\cdot v=e^{\beta_{0}(a_{t})}\cdot v\to 0\textup{ as }t\to\infty$
@@
 Therefore, $\mathbf{P}_{\beta_{0}}$ is a parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$.
-Let $\overline{\mathbf{P}}_{\beta_{0}}$ be the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$, and denote by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ the unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$.
-Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
+It is known that there exists a $\mathbb{Q}$-torus $\mathbf{T}_{\beta_{0}}$ in $\mathbf{T}$ such that $Z(\mathbf{T}_{\beta_{0}})$ (the centralizer of $\mathbf{T}_{\beta_{0}}$ in $\mathbf{G}$) is a Levi subgroup of $\mathbf{P}_{\beta_{0}}$.
+We denote by $\overline{\mathbf{P}}_{\beta_{0}}$ the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ with the same Levi subgroup $Z(\mathbf{T}_{\beta_{0}})$, and denote by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ the unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$.
+Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ in $\mathbf{G}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
 $_{\mathbb{Q}}{W}=N(\mathbf{T})/Z(\mathbf{T}).$
-Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above. Then the Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
-$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
+Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above, and $_{\mathbb{Q}}\mathcal{W}$} a set of representatives of $_{\mathbb{Q}}{W}$ in $N(\mathbf{T})(\mathbb{Q})$.
+Then the Bruhat decomposition of $\mathbf{G}$ is the following [8, Theorem 21.15]
+$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q}).$
+Note that the Weyl group $_{\mathbb{Q}}{W}$ acts transitively on the set of minimal parabolic $\mathbb{Q}$-subgroups containing $Z(\mathbf{T})$ [8, Corollary 21.3].
+So there exists an element in $_{\mathbb{Q}}{W}$ which sends $\mathbf{P}_{0}$ to $\overline{\mathbf{P}}_{0}$, and we let $\bar{w}$ be the representative of this element in $_{\mathbb{Q}}\mathcal{W}$}.
+Then we have
+$\mathbf{G}(\mathbb{Q})=\mathbf{G}(\mathbb{Q})\cdot\bar{w}^{-1}=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}\mathcal{W}}\cdot(\mathbf{P}_{0}(\mathbb{Q})\cdot\bar{w}^{-1})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}\mathcal{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q})$
 which implies that
-$\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
-If we choose a set $\{w_{i}\}_{i\in I}$ of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$, then for each $w_{i}$, define
-$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
-and we have
-$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$
-Then one can further write the Bruhat decomposition as
-$\mathbf{G}(\mathbb{Q})=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
-$=$ $\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
-where the index sets $I$ and $\bar{I}$ are defined as in §1. We denote by
-$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
-Note that the group $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}\subset\mathbf{P}_{\beta_{0}}$ is generated by unipotent subgroups whose Lie algebras are direct sums of $\mathbb{Q}$-root spaces in $\mathbf{G}$ relative to $\mathbf{T}$, and $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ by the structure of the representation $\rho$.
+$\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}\mathcal{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
+For each $w\in{{}_{\mathbb{Q}}\mathcal{W}}$, define
+$\mathbf{F}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad\mathbf{H}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}\mathbf{P}_{\beta_{0}}w$
+and by [8, Proposition 21.9], we have
+$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w}\cdot\mathbf{F}_{w}.$
+We may assume that $e\in{{}_{\mathbb{Q}}\mathcal{W}}$. It follows that
+$\mathbf{G}(\mathbb{Q})=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{Q})\cdot(w\mathbf{H}_{w}w^{-1})(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$
+$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
+Here, as discussed in §1.2, the subsets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$ $(w\in{{}_{\mathbb{Q}}\mathcal{W}})$ may overlap, and we choose any subset $_{\mathbb{Q}}\overline{\mathcal{W}}$} in $_{\mathbb{Q}}\mathcal{W}$} (as small as possible) such that
+$\mathbf{G}(\mathbb{Q})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
+Note that, by [8, Proposition 21.9], the group $w\mathbf{H}_{w}w^{-1}\subset\mathbf{P}_{\beta_{0}}$ is generated by unipotent subgroups whose Lie algebras are sums of root spaces of positive $\mathbb{Q}$-roots in $\mathfrak{g}$ relative to $\mathbf{T}$, and $\beta_{0}$ is the highest weight in $\rho$.
+Hence $w\mathbf{H}_{w}w^{-1}$ fixes every element in $V_{\beta_{0}}$ by the structure of the representation $\rho$.
+Now we prove Theorem 1.3.
+Proof of Theorem 1.3.
+Since $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is bounded, by definition
+$\tau(\psi)=\liminf_{t\to\infty}\left(-\frac{\ln(\psi(t))}{t}\right)\geq 0.$
+Note that we have already proved $\beta_{0}(a_{-1})>0$. Now suppose that $\tau=\tau(\psi)>\beta_{0}(a_{-1})$. Then for any sufficiently small $\epsilon>0$, there exists $C_{\epsilon}>0$ such that
+$\psi(t)\leq C_{\epsilon}\cdot e^{-(\tau-\epsilon)t}\quad(\forall t>0).$
+Let $g\in\mathbf{G}(\mathbb{R})$. Since $\beta_{0}(a_{-1})>0$ is the fastest contracting rate of the $\{a_{t}\}_{t\in\mathbb{R}}$-action on $V$, we have
+$\|\rho(a_{t})v\|\geq e^{-\beta_{0}(a_{-1})t}\|v\|\quad(\forall v\in\rho(g)\cdot\mathbb{Z}^{d}\setminus\{0\})$
+and
+$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq e^{-\beta_{0}(a_{-1})t}\cdot\delta(\rho(g)\cdot\mathbb{Z}^{d}).$
+Since $\tau>\beta_{0}(a_{-1})$, by choosing $\epsilon$ sufficiently small, one may find a constant $C>0$ such that
+$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq C\cdot\psi(t)\quad(t>0).$
+which shows that $g\in S_{\rho}(\psi)$ and $S_{\rho}(\psi)=\mathbf{G}(\mathbb{R})$. Note that this argument also works for $\tau=\infty$. This completes the proof of Theorem 1.3. ∎
+In the rest of the paper, we assume that $0\leq\tau(\psi)\leq\beta_{0}(a_{-1})$. In the following, we define the notion of rational elements and discuss some related properties.
 Definition 2.1.
-An element $g\in\mathbf{G}(\mathbb{R})$ is called rational if $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is Zariski dense in $V_{\beta_{0}}$.
+An element $g\in\mathbf{G}(\mathbb{R})$ is called rational if $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is Zariski dense in $V_{\beta_{0}}$, or equivalently, if $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is a lattice in the vector space of real points in $V_{\beta_{0}}$.
 Lemma 2.2.
@@
 and $\rho(\sigma(g))\cdot\Lambda_{g}$ is Zariski dense in $V_{\beta_{0}}$. Then we obtain that
-$\rho(\sigma(g)g^{-1})V_{\beta_{0}}=\rho(\sigma(g)g^{-1})\overline{\rho(g)\cdot\Lambda_{g}}=V_{\beta_{0}}$
+$\rho(\sigma(g)g^{-1})V_{\beta_{0}}=\rho(\sigma(g)g^{-1})\overline{\rho(g)\cdot\Lambda_{g}}=\overline{\rho(\sigma(g))\cdot\Lambda_{g}}=V_{\beta_{0}}$
 and $\sigma(g)g^{-1}$ is in the stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$.
-Here $\overline{\rho(g)\cdot\Lambda_{g}}$ denotes the Zariski closure of $\rho(g)\cdot\Lambda_{g}$.
+Here $\overline{\rho(g)\cdot\Lambda_{g}}$ and $\overline{\rho(\sigma(g))\cdot\Lambda_{g}}$ denote the Zariski closures of $\rho(g)\cdot\Lambda_{g}$ and $\rho(\sigma(g))\cdot\Lambda_{g}$ respectively.
 On the other hand, $\sigma(g)g^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cap\mathbf{P}_{\beta_{0}}=\{e\}$.
@@
 Corollary 2.3.
-Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if $g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$.
+Let $w\in{{}_{\mathbb{Q}}\mathcal{W}}$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+Then $w\cdot g$ is rational if and only if $g\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{F}_{w}(\mathbb{Q})$.
 Proof.
-Let $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that $g=h_{i}\cdot f_{i}$.
-Suppose that $w_{i}\cdot g$ is rational.
-Since $w_{i}\in\mathbf{G}(\mathbb{Q})$, by definition, $w_{i}\cdot h_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is also rational in $\mathbf{G}(\mathbb{R})$.
-We know that $w_{i}\cdot h_{i}\cdot w_{i}^{-1}$ preserves $V_{\beta_{0}}$, so $w_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is rational.
-By the fact that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and Lemma 2.2, we conclude that $w_{i}\cdot f_{i}\cdot w_{i}^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$.
-Conversely, suppose that $g=h_{i}\cdot f_{i}$ where $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Then
-$w_{i}\cdot g=(w_{i}h_{i}w_{i}^{-1})\cdot(w_{i}f_{i}w_{i}^{-1})\cdot w_{i}.$
-Note that $w_{i}h_{i}w_{i}^{-1}\in\mathbf{P}_{\beta_{0}}$ and $w_{i}\in\mathbf{G}(\mathbb{Q})$. By definition, we know that $w_{i}\cdot g$ is rational. This completes the proof of the lemma. ∎
+Let $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{R})$ such that $g=h\cdot f$.
+Suppose that $w\cdot g$ is rational.
+Since $w\in\mathbf{G}(\mathbb{Q})$, by definition, $w\cdot g\cdot w^{-1}=w\cdot h\cdot f\cdot w^{-1}$ is also rational in $\mathbf{G}(\mathbb{R})$.
+Since $w\cdot h\cdot w^{-1}$ preserves $V_{\beta_{0}}$, we get that $w\cdot f\cdot w^{-1}$ is rational.
+By the fact that $w\mathbf{F}_{w}w^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and Lemma 2.2, we conclude that $w\cdot f\cdot w^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ and $f\in\mathbf{F}_{w}(\mathbb{Q})$.
+Conversely, suppose that $g=h\cdot f$ where $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{Q})$. Then
+$w\cdot g=(whw^{-1})\cdot(wfw^{-1})\cdot w.$
+Note that $whw^{-1}\in\mathbf{P}_{\beta_{0}}$ and $w\in\mathbf{G}(\mathbb{Q})$. By definition, we know that $w\cdot g$ is rational. This completes the proof of the corollary. ∎
 By properties of the subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ in the reduction theory, we can give another characterization of rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
 Lemma 2.4.
-Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if
-$g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma).$
+Let $w\in{{}_{\mathbb{Q}}\mathcal{W}}$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w\cdot g$ is rational if and only if
+$g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma).$
 Proof.
-If $w_{i}\cdot g$ is rational, then by Corollary 2.3
-$g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{G}(\mathbb{Q})=\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
-Suppose that $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma)$.
-Then there exist $h\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
-$g=h\cdot f,\quad f\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
+If $w\cdot g$ is rational, then by Corollary 2.3, we have
+$g\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{F}_{w}(\mathbb{Q})\subset\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{G}(\mathbb{Q})=\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma\subset\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
+Now suppose that $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma)$.
+Then there exist $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{R})$ such that
+$g=h\cdot f,\quad f\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
 We write
-$f=h_{i}\cdot p\cdot x\cdot\gamma$
-for some $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$. Let $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Then we have
-$\sigma(f)=\sigma(h_{i})\cdot\sigma(p)\cdot x\cdot\gamma.$
-This implies that $x\cdot\gamma=p^{-1}\cdot h_{i}^{-1}\cdot f=\sigma(p)^{-1}\cdot\sigma(h_{i})^{-1}\cdot\sigma(f)$. Since the product map
-$\mathbf{P}_{0}\times\mathbf{H}_{w_{i}}\times\mathbf{F}_{w_{i}}\to\mathbf{G}$
-is injective, we have $\sigma(f)=f$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$ and hence $f\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. By Corollary 2.3, $w_{i}\cdot g$ is rational. ∎
+$f=h\cdot p\cdot x\cdot\gamma$
+for some $h\in\mathbf{H}_{w}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$. Let $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Then we have
+$\sigma(f)=\sigma(h)\cdot\sigma(p)\cdot x\cdot\gamma.$
+This implies that $x\cdot\gamma=p^{-1}\cdot h^{-1}\cdot f=\sigma(p)^{-1}\cdot\sigma(h_{i})^{-1}\cdot\sigma(f)$. Since the product map
+$\mathbf{P}_{0}\times\mathbf{H}_{w}\times\mathbf{F}_{w}\to\mathbf{G}$
+is injective, we have $\sigma(f)=f$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$ and hence $f\in\mathbf{F}_{w}(\mathbb{Q})$. By Corollary 2.3, $w\cdot g$ is rational. ∎
 Definition 2.5.
 Let $g$ be a rational element in $\mathbf{G}(\mathbb{R})$.
-The denominator (or the height) $d(g)$ of $g$ is defined to be the co-volume of the lattice $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$.
-In the following lemma, we discuss the relation between the shortest vector and the co-volume of a discrete subgroup of the form $\rho(g)\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}$.
+The height $0pt(g)$ of $g$ is defined to be the co-volume of the lattice $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in the vector space of real points in $V_{\beta_{0}}$.
+In the following lemma, we discuss the relation between the length of the shortest non-zero vector and the co-volume of the discrete subgroup $\rho(g)\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}$ for $g\in\mathbf{G}(\mathbb{R})$.
+For convenience, we write $A\lesssim B$ $(A\gtrsim B)$ if there exists a constant $c>0$ such that
+$A\leq c\cdot B\quad(A\geq c\cdot B).$
+If $A\lesssim B$ and $A\gtrsim B$, then we write $A\asymp B$. We will specify the implicit constants in the contexts if necessary.
 Lemma 2.6.
-Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Suppose that $w_{i}\cdot g$ is rational. Then
-$d(w_{i}\cdot g)\sim\delta(\rho(w_{i}\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})^{\dim V_{\beta_{0}}}.$
+Let $w\in{{}_{\mathbb{Q}}\mathcal{W}}$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Suppose that $w\cdot g$ is rational. Then
+$0pt(w\cdot g)\asymp\delta(\rho(w\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})^{\dim V_{\beta_{0}}}.$
 Here the implicit constant depends only on $\mathbf{G}$ and $\Gamma$.
 Proof.
-By Corollary 2.3, there exist $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$ such that $g=h_{i}\cdot f_{i}$. Note that
-$f_{i}\in\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
-and there exist $p\in\mathbf{P}_{0}(\mathbb{Q})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that $f_{i}=w_{i}^{-1}\cdot p\cdot x\cdot\gamma$.
+By Corollary 2.3, there exist $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{Q})$ such that $g=h\cdot f$. Note that
+$f\in\mathbf{G}(\mathbb{Q})=w^{-1}\cdot\mathbf{G}(\mathbb{Q})=w^{-1}\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
+and there exist $p\in\mathbf{P}_{0}(\mathbb{Q})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that $f=w^{-1}\cdot p\cdot x\cdot\gamma$.
 We know that $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the Levi factor of $\mathbf{P}_{0}$.
@@
 Then we have
-$w_{i}\cdot g=(w_{i}\cdot h_{i}\cdot w_{i}^{-1})\cdot u\cdot p_{1}\cdot p_{2}\cdot p_{3}\cdot x\cdot\gamma.$
-Since $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ and $R_{u}(\mathbf{P}_{0})$ fix every element in $V_{\beta_{0}}$, and $\mathbf{T}$ and $\mathbf{M}_{a}$ preserve the weight space $V_{\beta_{0}}$, we obtain
-$\rho(w_{i}\cdot g)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2}\cdot p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2})(\rho(p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}).$
-The lemma then follows from the facts that $p_{1}\in\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$, $p_{2}$ is in a fixed compact subset in $\mathbf{M}_{a}(\mathbb{R})$ and $\rho(p_{3}\cdot x)\mathbb{Z}^{d}$ is commensurable with $\mathbb{Z}^{d}$ by $\Gamma$ and $\mathcal{K}$.
+$w\cdot g=(w\cdot h\cdot w^{-1})\cdot u\cdot p_{1}\cdot p_{2}\cdot p_{3}\cdot x\cdot\gamma.$
+Since $w\mathbf{H}_{w}w^{-1}$ and $R_{u}(\mathbf{P}_{0})$ fix every element in $V_{\beta_{0}}$, and $\mathbf{T}$ and $\mathbf{M}_{a}$ preserve the weight space $V_{\beta_{0}}$, we obtain
+$\rho(w\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2}\cdot p_{3}\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2})(\rho(p_{3}\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}).$
+The lemma then follows from the facts that $p_{1}\in\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$, $p_{2}$ is in a fixed compact subset in $\mathbf{M}_{a}(\mathbb{R})$ and $\rho(p_{3}\cdot x)\cdot\mathbb{Z}^{d}$ is commensurable with $\mathbb{Z}^{d}$ by $\Gamma$ and $\mathcal{K}$.
 ∎
-Now we consider the case that $w_{i}=e$ is the identity element in ${}_{\mathbb{Q}}W$.
+Now we consider the case where $w=e$ is the identity element in $_{\mathbb{Q}}\mathcal{W}$.
+Then $\mathbf{F}_{w}=\mathbf{F}_{e}$ and $\mathbf{H}_{w}=\mathbf{H}_{e}$.
 Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-Then $\mathbf{F}_{w_{i}}=\mathbf{F}_{e}$ and $\mathbf{H}_{w_{i}}=\mathbf{H}_{e}$.
 By Lemma 2.4, there exist $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that
@@
 $p=a\cdot m\cdot u$
-where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$. Then we can compute the denominator $d(g)$ of $g$ as follows:
+where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$. Then we can compute the height $0pt(g)$ of $g$ as follows:
 $(\rho(g)\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}=\rho(a\cdot m\cdot u\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(a\cdot m\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(a\cdot m)(\rho(x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})$
 and
-$d(g)=c_{x}\cdot e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}$
+$0pt(g)=c_{x}\cdot e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}$
 for some constant $c_{x}>0$ depending only on $x$.
@@
 for some $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x_{j}\in\mathcal{K}$ and $\gamma\in\Gamma$.
-3. Counting rational points
-In this section, we consider the problem of counting rational points in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-To estimate the Hausdorff dimension, one usually constructs boxes centered at certain rational points and obtains a tree-like subset.
-At each level of the tree-like subset, the boxes are disjoint.
-In our case, it may happen that the boxes at each level are not disjoint if we proceed in the usual way.
-It implies that the boxes constructed are too many, and we have to sieve out some proportion of the rational points so that the boxes centered at the remaining rational points are disjoint.
-It is also required that the number of the remaining rational points is not small compared to the total number of rational points in order to avoid any loss of Hausdorff dimension.
-For our purpose, in the following, we will first sieve some proportion of rational points according to several algebraic conditions, and then count the remaining rational points by the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$.
-The disjointness property of the boxes centered at the remaining rational points will follow from the transversal structure of some submanifolds in the homogeneous space $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
+Furthermore, if $p=a\cdot m\cdot u$ where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, then $a$ is called the central coordinate of $g$ which we denote by $a_{g}$.
+3. Counting rational elements
+In this section, we consider the problem of counting rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+As discussed in §1, to get a lower bound of the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, we need to select certain rational elements and construct a Cantor type subset from disjoint neighborhoods of these rational elements.
+For our purpose, in the following, we will select rational elements according to several algebraic conditions, and then count these rational elements by the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
+The disjointness of the neighborhoods around these rational elements will follow from the transversal structure of some submanifolds in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
 For any $\mathbb{Q}$-algebraic group $\mathbf{L}$ in $\mathbf{G}$ and $\delta>0$, we write $B_{\mathbf{L}}(\delta)$ for the open ball of radius $\delta>0$ around the identity in $\mathbf{L}(\mathbb{R})$.
@@
 $B_{\mathbf{L}}(\delta,t):=a_{-t}\cdot B_{\mathbf{L}}(\delta)\cdot a_{t}.$
-The connected component of identity of $\mathbf{L}$ in the Zariski topology is denoted by $\mathbf{L}^{0}$, and the connected component of identity in the Lie group $\mathbf{L}(\mathbb{R})$ is denoted by $\mathbf{L}(\mathbb{R})^{0}$.
-We write $A\lesssim B$ $(A\gtrsim B)$ if there exists a constant $c>0$ such that
-$A\leq c\cdot B\quad(A\geq c\cdot B).$
-If $A\lesssim B$ and $A\gtrsim B$, then we write $A\sim B$. We will specify the implicit constants in the contexts if necessary.
-In this section and next section, we assume that the action of the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
-This holds true when the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into any $\mathbb{Q}$-simple factor of $\mathbf{G}$.
-Later in §5, we will explain how to establish the results of counting rational points and Hausdorff dimension estimates in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ appears only in some $\mathbb{Q}$-simple factors of $\mathbf{G}$.
-Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. For any $0<A<B$, define
-$S(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
-We know that the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ has the following decomposition
+The identity component of $\mathbf{L}$ with respect to the Zariski topology is denoted by $\mathbf{L}^{0}$, and the identity component of $\mathbf{L}(\mathbb{R})$ (as a Lie group) is denoted by $\mathbf{L}(\mathbb{R})^{0}$.
+In this section and §4, we assume that the action of the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
+This assumption holds when $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into any $\mathbb{Q}$-simple factor of $\mathbf{G}$.
+Later in §5, we will explain how to establish the results about counting rational elements and estimating Hausdorff dimensions in the case where $\{a_{t}\}_{t\in\mathbb{R}}$ projects trivially into some of the $\mathbb{Q}$-simple factors of $\mathbf{G}$.
+Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which projects injectively into $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$.
+For any $0<A<B$, define
+$S(U,A,B)=\{q\in U:q\textup{ rational and }A\leq 0pt(q)\leq B\}.$
+The Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ has the following direct sum decomposition
 $\mathfrak{a}=\operatorname{Lie}(a_{t})\oplus\ker(\beta_{0})$
 where $\operatorname{Lie}(a_{t})$ is the Lie algebra of the one parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ and we have
-$\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0.$
-Denote by $\pi_{\ker(\beta_{0})}$ the projection from $\mathfrak{a}$ to $\ker(\beta_{0})$ according to this decomposition.
-We also write $\pi_{\ker(\beta_{0})}(a)$ for $\pi_{\ker(\beta_{0})}(\log(a))$ whenever $a\in\exp(\mathfrak{a})$.
-Here $\log$ is the inverse of the exponential map $\exp$.
+$\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0\;(t>0).$
+Denote by $\pi_{\ker(\beta_{0})}$ the projection of $\mathfrak{a}$ onto $\ker(\beta_{0})$ along the linear subspace $\operatorname{Lie}(a_{t})$.
+For convenience, we will write $\pi_{\ker(\beta_{0})}(a)$ for $\pi_{\ker(\beta_{0})}(\log(a))$ whenever $a\in\exp(\mathfrak{a})$ (here $\log$ is the inverse of the exponential map $\exp$).
 For any compact subset $L$ in $\operatorname{Lie}(a_{t})$ and any compact subset $K$ in $\ker(\beta_{0})$, we denote by
 $\mathfrak{a}_{L,K}=\{x\in\mathfrak{a}:x=y_{1}+y_{2},y_{1}\in L,y_{2}\in K\}.$
-For any $x_{j}\in\mathcal{K}$, any compact subset $K_{1}$ in $\ker(\beta_{0})$, any compact subset $K_{2}$ in $\mathbf{H}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$, we define
+For any $x_{j}\in\mathcal{K}$ ($j\in J$), any compact subset $K_{1}$ in $\ker(\beta_{0})$, any compact subset $K_{2}$ in $\mathbf{H}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$, we define
 $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$
-to be the set of all rational points $q$ in $U$ such that
+to be the set of all rational elements $q$ in $U$ such that
   1. (1)
-    $A\leq d(q)\leq B$ and $q$ is $j$-rational for $x_{j}\in\mathcal{K}$;
+    $A\leq 0pt(q)\leq B$ and $q$ is $j$-rational for $x_{j}\in\mathcal{K}$;
   2. (2)
@@
 then $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$ (and such elements exist as $e\in\mathcal{K}$).
-The elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ are the rational points we are interested in when we construct a tree-like subset in §4.
+The elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ are the rational elements we select when we construct a Cantor type subset in §4.
 We denote by
 $S_{K_{1},K_{2},K_{3},K_{4}}(U,A,B):=\bigcup_{j\in J}S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B).$
-To count the rational points in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$, we need results about limiting distributions of translates of unipotent orbits pushed by the one-parameter semisimple flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
-The following is a direct consequence of the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
+To count the rational elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ $(j\in J)$, we need the following result about limiting distributions of translates of unipotent orbits pushed by the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$, which is a direct consequence of the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
 Proposition 3.1.
@@
 Let $\chi_{W}$ denote the characteristic function of $W$.
-Then for any bounded open subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})$ we have
-[…]
+Then for any bounded open subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ we have
+$\lim_{t\to\infty}\frac{1}{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)}\int_{U}\chi_{W}(a_{t}\cdot nx)d\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(n)=\int_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}\chi_{W}d\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}$
 where $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}$ is the Haar measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
 Remark 3.2.
-Note that $R_{u}(\overline{\mathbf{P}}_{0})$ is not necessarily the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$.
-Here we can still apply the mixing property to obtain Proposition 3.1 as long as $R_{u}(\overline{\mathbf{P}}_{0})$ is contained in the group generated by the unstable subgroup and the central subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$.
-(See [25, §2].)
-Now we can follow the same arguments as in [17, 48] and measure the size of the subset $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ for a fixed index $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$, and for some sufficiently large numbers $B>A>0$.
+Note that $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is not necessarily the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}(\mathbb{R})$.
+Here we can still apply the mixing property to obtain Proposition 3.1 as long as $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is contained in the group generated by the unstable subgroup and the centralizer of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}(\mathbb{R})$.
+(See [49, §2].)
+Now we define a measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, which will be used to study the size of the subset $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ for some sufficiently large numbers $0<A<B$.
 Recall that
@@
 where $\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\cap\mathbf{P}_{\beta_{0}}$.
-We fix a Haar measure $\mu_{\mathbf{H}_{e}}$ on $\mathbf{H}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\mathbf{F}_{e}}$ on $\mathbf{F}_{e}(\mathbb{R})$ such that the product maps
-$\mathbf{F}_{e}\times\mathbf{H}_{e}\to\mathbf{F}_{e}\cdot\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\textup{ and }\mathbf{H}_{e}\times\mathbf{F}_{e}\to\mathbf{H}_{e}\cdot\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{0})$
-induce Haar measures on $R_{u}(\overline{\mathbf{P}}_{0}).$ Then for any $q\in\mathbf{F}_{e}(\mathbb{Q})$, we define $m_{\mathbf{H}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which is supported on $\mathbf{H}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\mathbf{H}_{e}}$ via the product map
-$\mathbf{H}_{e}\times\{q\}\to\mathbf{H}_{e}\cdot q\subset R_{u}(\overline{\mathbf{P}}_{0}).$
-We define
+By [55, Proposition 5.26], we fix a Haar measure $\mu_{\mathbf{H}_{e}}$ on $\mathbf{H}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\mathbf{F}_{e}}$ on $\mathbf{F}_{e}(\mathbb{R})$ such that the product maps
+[…]
+induce Haar measures on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ For any $q\in\mathbf{F}_{e}(\mathbb{Q})$, we define $m_{\mathbf{H}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which is supported on $\mathbf{H}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\mathbf{H}_{e}}$ via the product map
+$\mathbf{H}_{e}(\mathbb{R})\times\{q\}\to\mathbf{H}_{e}(\mathbb{R})\cdot q\subset R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$
+Then we define
 $m_{\mathbf{H}_{e}}:=\sum_{q\in\mathbf{F}_{e}(\mathbb{Q})}m_{\mathbf{H}_{e}q}.$
-Note that $m_{\mathbf{H}_{e}}$ is not a locally finite measure on $R_{u}(\overline{\mathbf{P}}_{0})$, and it is defined by the leaf-wise measures on the countable leaves through the rational elements in $\mathbf{F}_{e}(\mathbb{Q})$ in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ induced by the group action of $\mathbf{H}_{e}$ on $R_{u}(\overline{\mathbf{P}}_{0})$.
+Note that $m_{\mathbf{H}_{e}}$ is not a locally finite measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, and it is defined by the leaf-wise measures on the countable leaves $\mathbf{H}_{e}(\mathbb{R})\cdot q$ $(q\in\mathbf{F}_{e}(\mathbb{Q}))$ in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ induced by the group action of $\mathbf{H}_{e}(\mathbb{R})$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
 Note also that by Corollary 2.3, we have
 $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\subset\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\mathbf{H}_{e}(\mathbb{R})\cdot q=\mathbf{H}_{e}(\mathbb{R})\cdot\mathbf{F}_{e}(\mathbb{Q}).$
-In the following, we estimate the size of the subset
+In the following, we use the arguments in [25, 82] and estimate the $m_{\mathbf{H}_{e}}$-measure of the subset
 $S_{K_{1},K_{2},K_{3},K_{4},j}(U,(l/2)^{\dim V_{\beta_{0}}},l^{\dim V_{\beta_{0}}})$
-with respect to the measure $m_{\mathbf{H}_{e}}$ for sufficiently large $l>0$. For convenience, we write
+for $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}\cap\mathcal{K}$ and for sufficiently large $l>0$. For convenience, we write
 $A_{l}=(l/2)^{\dim V_{\beta_{0}}}\textup{ and }B_{l}=l^{\dim V_{\beta_{0}}}\quad(\forall l>0).$
@@
 $\beta_{0}(a_{T})=-\ln l.$
-Let $q$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, we may write
+Let $q$ be a rational element in $U\subset R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, we may write
 $q=a\cdot h\cdot m\cdot u\cdot x_{k}\cdot\gamma\in\mathbf{H}_{e}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot x_{k}\cdot\Gamma$
-for some $a\in\mathbf{T}(\mathbb{R})$, $h\in\mathbf{H}_{e}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$, $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, $x_{k}\in C$ and $\gamma\in\Gamma$. Then
+for some $a\in\mathbf{T}(\mathbb{R})$, $h\in\mathbf{H}_{e}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$, $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, $x_{k}\in\mathcal{K}$ and $\gamma\in\Gamma$. Then
 $q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
-$\iff$ $q\in U,\;A_{l}\leq d(q)\leq B_{l},\textrm{ and }a\in\exp(\mathfrak{a}),\pi_{\ker(\beta_{0})}(a)\in K_{1},h\in K_{2},m\in K_{3},u\in K_{4},k=j.$
+$\iff$ $q\textup{ rational in }U,\;A_{l}\leq 0pt(q)\leq B_{l},\textrm{ and}$
+$\quad a\in\exp(\mathfrak{a}),\pi_{\ker(\beta_{0})}(a)\in K_{1},h\in K_{2},m\in K_{3},u\in K_{4},k=j.$
 $\iff$ $a_{T}\cdot q\Gamma\in a_{T}\cdot U\Gamma/\Gamma\textrm{ and }a_{T}\cdot q\Gamma\in\exp(\mathfrak{a}_{I_{0},K_{1}})\cdot K_{2}\cdot K_{3}\cdot K_{4}\cdot x_{j}\cdot\Gamma/\Gamma$
@@
 This implies that
-$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
-Since $\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$ is a compact subset in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$, there exist $\delta_{0}>0$ and a small neighborhood of identity $B_{\mathbf{F}_{e}}(\delta_{0})$ in $\mathbf{F}_{e}(\mathbb{R})$ such that
+$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
+Since $\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}$ is a compact subset in $\mathbf{P}_{\beta_{0}}(\mathbb{R})$ and
+$\mathbf{F}_{e}(\mathbb{R})\cap\mathbf{P}_{\beta_{0}}(\mathbb{R})=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})\cap\mathbf{P}_{\beta_{0}}(\mathbb{R})=\{e\},$
+there exists a small neighborhood of identity $B_{\mathbf{F}_{e}}(\delta_{0})$ in $\mathbf{F}_{e}(\mathbb{R})$ (for some $\delta_{0}>0$) such that
 $B_{\mathbf{F}_{e}}(\delta_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
-is a homeomorphism. We conclude that the following product map
-$B_{\mathbf{F}_{e}}(\delta_{0})\times a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$
-gives a transversal structure near the set $a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$ in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
+is a homeomorphism, and hence the following product map
+(1) $B_{\mathbf{F}_{e}}(\delta_{0})\times a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma$
+is bijective.
 Consequently, for any $p,q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, the subsets $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot p\Gamma$ and $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot q\Gamma$ are disjoint, where
 $B_{\mathbf{F}_{e}}(\delta_{0},T):=a_{-T}\cdot B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}.$
-Now we estimate the size of $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ with respect to the measure $m_{\mathbf{H}_{e}}$.
+Now we estimate the $m_{\mathbf{H}_{e}}$-measure of $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, following the arguments in [25, §4].
 First we prove an upper bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$.
@@
 $B_{\mathbf{F}_{e}}(\epsilon,T)\subset B_{\mathbf{F}_{e}}(\epsilon)\textup{ and }B_{\mathbf{F}_{e}}(\epsilon,T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\subset B_{\mathbf{F}_{e}}(\epsilon)\cdot U\Gamma/\Gamma.$
-Since $B_{\mathbf{F}_{e}}(\delta_{0},T)$ and $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ form a transversal structure near $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, by Proposition 3.1, we deduce that
-$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))$
-[…]
-$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\cdot\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}(B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma)$
-$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))$
+Since the map (1) is bijective, the following product map
+$B_{\mathbf{F}_{e}}(\delta_{0},T)\times S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma$
+is also bijective.
+By Proposition 3.1 and the fact that $U$ projects injectively into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$, we deduce that
+$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))$
+$=$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon,T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$
+$\leq$ $\int_{B_{\mathbf{F}_{e}}(\epsilon)\cdot U}\chi_{B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma}(a_{T}u\Gamma)d\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(u)$
+$\asymp$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\cdot\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}(B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma)$
+$\asymp$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))$
 as $T\to\infty$.
@@
 where the implicit constant depends only on $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $I_{0}$, $\mathbf{G}$ and $\Gamma$.
-Similarly, using the arguments in [17, §4], we can prove a lower bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$ and for any sufficiently large $l>0$ we have
+Similarly, using the arguments in [25, §4], we can prove a lower bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$ and for any sufficiently large $l>0$ we have
 $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\gtrsim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
@@
 Proposition 3.3.
-Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})$. Then for any $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
-$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
-and
-$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
+Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which projects injectively into $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$.
+Then for any $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$, we have
+$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
+as $l\to\infty$. Consequently, we have
+$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
 Here the implicit constants depend only on the compact subsets $K_{i}$’s, $\mathbf{G}$ and $\Gamma$.
-Remark 3.4.
-For our purpose, in the following, we fix some compact subsets $K_{1}\subset\ker\beta_{0}$, $K_{2}\subset\mathbf{H}_{e}(\mathbb{R})$, $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$ discussed above.
-4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
-In this section, we prove Theorem 1.3 under the assumption that the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
+4. Proof of Theorem 1.4: The case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the flow $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing
+In this section, we prove Theorem 1.4 under the assumption that $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
+In the following, we fix some compact subsets $K_{1}\subset\ker\beta_{0}$, $K_{2}\subset\mathbf{H}_{e}(\mathbb{R})$, $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$.
 Lemma 4.1.
-Fix $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and
+Fix $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$. Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and
 $\beta_{0}(a_{T})=-\ln l$
@@
 By the discussion in §3, we know that
-$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
+$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
 Choose $\theta_{1},\theta_{2}>0$ sufficiently small so that for any $a\in\exp(\mathfrak{a}_{I_{0},K_{1}})$ we have
@@
 Then
-$a_{T}\cdot B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p=B_{\mathbf{H}_{e}}(\theta_{1})\cdot a\cdot h\cdot m\cdot u\cdot\gamma\subset\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma$
+$a_{T}\cdot B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p=B_{\mathbf{H}_{e}}(\theta_{1})\cdot a\cdot h\cdot m\cdot u\cdot x_{j}\cdot\gamma\subset\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma$
 where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. By definition, we have
@@
 This completes the proof of the lemma. ∎
-Lemma 4.2.
-We have
-$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}>0.$
-Proof.
-Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{e})$, $\alpha(a_{1})=0$.
-Let $\Omega_{\mathbf{F}_{e}}$ and $\Omega_{\mathbf{H}_{e}}$ be small open neighborhoods of identity in $\mathbf{F}_{e}(\mathbb{R})$ and $\mathbf{H}_{e}(\mathbb{R})$ respectively, and
-$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{e}}\cdot\Omega_{\mathbf{H}_{e}}.$
-Note that $\mathbf{H}_{e}(\mathbb{R})$ stabilizes every element in $V_{\beta_{0}}$.
-Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{e}}$ and $h\in\Omega_{\mathbf{H}_{e}}$, we have
-$a_{t}\cdot p=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}=f\cdot(a_{t}ha_{-t})\cdot a_{t}.$
-The element $a_{t}$ is rational and
-$\delta(\rho(a_{t})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(a_{t})}.$
-This implies that
-$\delta(\rho(a_{t}\cdot p)\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(a_{-1})t}$
-for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{e}}$.
-Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$.
-Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
-This contradicts the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
-∎
 Let $X$ be a Riemannian manifold, $m$ a volume form on $X$ and $E$ a compact subset of $X$.
 Denote by $\operatorname{diam}(S)$ the diameter of a set $S\subset X$.
-A collection $\mathcal{A}$ of compact subsets of $E$ is said to be tree-like if $\mathcal{A}$ is the union of finite sub-collections $\mathcal{A}_{k}$ such that
+A collection $\mathcal{A}$ of compact subsets of $E$ is said to be tree-like if $\mathcal{A}$ is the union of finite sub-collections $\mathcal{A}_{k}$ $(k\in\mathbb{N})$ such that
   1. (1)
@@
   2. (2)
-    For any $k$ and $S_{1},S_{2}\in\mathcal{A}_{j}$, either $S_{1}=S_{2}$ or $S_{1}\cap S_{2}=\emptyset$;
+    For any $k$ and $S_{1},S_{2}\in\mathcal{A}_{k}$, either $S_{1}=S_{2}$ or $S_{1}\cap S_{2}=\emptyset$;
   3. (3)
@@
     $d_{k}(\mathcal{A}):=\sup_{S\in\mathcal{A}_{k}}\operatorname{diam}(S)\to 0$ as $k\to\infty$.
-We write $\mathbf{A}_{k}:=\bigcup_{A\in\mathcal{A}_{k}}A$ and $\mathbf{A}_{\infty}:=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. We also define
+We write $\mathbf{A}_{k}:=\bigcup_{A\in\mathcal{A}_{k}}A$ and define a Cantor-type subset $\mathbf{A}_{\infty}:=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. We also define
 $\Delta_{k}(\mathcal{A}):=\inf_{S\in\mathcal{A}_{k}}\frac{m(\mathbf{A}_{k+1}\cap S)}{m(S)}.$
+We will use the following result to compute the lower bound of $\dim_{H}S_{\rho}(\psi)^{c}$ in Theorem 1.4.
 Theorem 4.2 ([49, 61, 80]).
@@
 $\dim_{H}(\mathbf{A}_{\infty})\geq\dim_{H}X-\limsup_{k\to\infty}\frac{\sum_{i=0}^{k}\log(\Delta_{i}(\mathcal{A}))}{\log(d_{k+1}(\mathcal{A}))}.$
-Proof of Theorem 1.3.
-We fix the compact subsets $K_{1}\subset\ker(\beta_{0}),K_{2}\subset\mathbf{H}_{e}(\mathbb{R}),K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0},K_{4}\subset R_{u}(\mathbf{P}_{0})$ as in Remark 3.4, and also fix $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$.
-Let $\epsilon>0$ be a sufficiently small number.
-Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
+Proof of Theorem 1.4 assuming that $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the flow $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing.
+We first get a lower bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+We fix some compact subsets $K_{1}\subset\ker(\beta_{0}),K_{2}\subset\mathbf{H}_{e}(\mathbb{R}),K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$ as in the begining of this section, and also fix some index $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$.
+Assume that $0\leq\tau=\tau(\psi)<\beta_{0}(a_{-1})$.
+Let $\epsilon>0$ be a sufficiently small number such that $\tau+\epsilon<\beta_{0}(a_{-1})$.
+Let $\nu_{0}$ be any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
 $\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
-Then by Lemma 4.2, $\nu_{0}(a_{1})>0$.
-We start with a small open bounded box $U_{0}$ in $R_{u}(\overline{\mathbf{P}}_{0})$.
+Since the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, we have $\nu_{0}(a_{1})>0$.
+In the following, we construct a tree-like collection $\mathcal{A}=\{\mathcal{A}_{k}\}_{k\in\mathbb{N}}$ of compact subsets in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+We start with a small open bounded box $U_{0}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which projects injectively into $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma)$.
 For $k=0$, we set $\mathcal{A}_{0}=\left\{U_{0}\right\}$.
@@
 By Proposition 3.3, we know that
-$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
+$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
 and
-$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
-where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for some $\theta_{2}>0$ as defined in Lemma 4.1.
+$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
+where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for some $\theta_{2}>0$ defined in Lemma 4.1.
 Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
 $\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset,$
-we devide the region $\mathcal{F}_{q}\cap U_{0}$ into small cubes of side length
+we divide the region $\mathcal{F}_{q}\cap U_{0}$ into small cubes of side length
 $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$
 where $\theta_{1}>0$ is the constant defined in Lemma 4.1.
-Then we collect those cubes $R$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$, and denote the corresponding collection by
-$\mathcal{G}_{1,q}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset\}.$
+Then we collect those cubes $R\subset\mathcal{F}_{q}\cap U_{0}$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$, and denote the corresponding collection by
+$\mathcal{G}_{1,q}:=\{R:R\subset\mathcal{F}_{q}\cap U_{0},\;R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset\}.$
 Note that $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{1}}))/10$ is smaller than the minimum side length of the rectangle $B_{\mathbf{H}_{e}}(\theta_{1},T_{1})$, so by Lemma 4.1, we know that each cube in $\mathcal{G}_{1,q}$ is contained in
 $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$
-where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
-$\mathcal{H}_{1}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R.$
+where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$.
+For any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ $(q\in\mathbf{F}_{e}(\mathbb{Q}))$ in $\mathcal{F}_{\mathbf{H}_{e}}$ such that
+$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})=\emptyset,$
+we set $\mathcal{G}_{1,q}=\emptyset$.
+Now let
+$\mathcal{H}_{1}=\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{1,q}}R.$
 Then we have
 $l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
-$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
-$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R)$
+$\asymp m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
+$\leq m_{\mathbf{H}_{e}}(\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{1,q}}R)$
 $\leq m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
-$\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
-and
-$m_{\mathbf{H}_{e}}(\mathcal{H}_{1})\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0}).$
+$\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
+and hence
+$m_{\mathbf{H}_{e}}(\mathcal{H}_{1})\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0}).$
 Note that each cube $R$ in $\mathcal{H}_{1}$ is contained in $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$. Let
@@
 By the computations in §3, we can choose a sufficiently small number $\tilde{\delta}_{0}>0$ such that
-$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
+$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
 is a homeomorphism, and the subsets in the collection
@@
 are disjoint by this homeomorphism as
-$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{1}).$
+$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{1})\textup{ and }a_{T_{1}}\cdot q\Gamma\in\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\;(\forall q\in\mathcal{H}_{1}).$
 We write
@@
 $a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}$
-may be larger than the diameter of $U_{0}$, so some cubes we obtain here from dividing the subset $\mathcal{P}_{1}$ may be outside the set $U_{0}$.
+may be larger than the diameter of $U_{0}$, so some cubes we obtain here (by dividing the subset $\mathcal{P}_{1}$) may be outside the set $U_{0}$.
 For our purpose, we collect only those cubes which are inside the subset $U_{0}$.
-In this manner, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{1}(U_{0})$ inside $U_{0}$, which we denote by $\mathcal{A}_{1}$.
-We remark here that if $\Phi^{0}(\mathbf{F}_{e})=\emptyset$, then all the cubes we obtain from dividing $\mathcal{P}_{1}$ are inside the subset $U_{0}$ if $T_{1}>0$ and $l_{1}>0$ are chosen to be sufficiently large (and also if we shrink or enlarge $U_{0}$ slightly to avoid the complexity caused by the boundary of $U_{0}$).
-We will see later that in the computations there is little difference between the case $\Phi^{0}(\mathbf{F}_{e})\neq\emptyset$ and the case $\Phi^{0}(\mathbf{F}_{e})=\emptyset$.
-Indeed, when we apply Theorem 4.3 and calculate $\Delta_{i}(\mathcal{A})$ and $d_{i}(\mathcal{A})$, the difference between these two cases may affect the value of the formula in finite steps, but eventually when we take the limit, this difference will disappear since we keep choosing sufficiently large numbers $T_{1},l_{1}$ and later $T_{2},l_{2},T_{3},l_{3},\dots$ to offset the effects by the difference at early stages.
-Similarly, we can construct $\mathcal{A}_{k}$ inductively for any $k\in\mathbb{N}$.
-For $k>1$ we choose sufficiently large numbers $l_{k}>0$ and $T_{k}>0$ such that $\beta_{0}(a_{\tau_{k}})=-\ln l_{k}$.
-For each cube $S\in\mathcal{A}_{k-1}$, by the proposition above, we know that
-$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
+In this manner, we obtain a family of disjoint cubes inside $U_{0}$ constructed from $\mathcal{P}_{1}$, which we denote by $\mathcal{A}_{1}$.
+We remark here that if $\Phi^{0}(\mathbf{F}_{e})=\emptyset$, then all the cubes we obtain by dividing $\mathcal{P}_{1}$ are inside the subset $U_{0}$ if $T_{1}>0$ and $l_{1}>0$ are chosen to be sufficiently large (and also if we enlarge $U_{0}$ slightly to avoid the complexity caused by the boundary of $U_{0}$).
+We will see later that in the computations there are no significant differences between the case $\Phi^{0}(\mathbf{F}_{e})\neq\emptyset$ and the case $\Phi^{0}(\mathbf{F}_{e})=\emptyset$.
+Indeed, when we apply Theorem 4.2 and calculate $\Delta_{i}(\mathcal{A})$ and $d_{i}(\mathcal{A})$, the differences between these two cases may affect the value of the formula in finite steps, but eventually when we take the limit, these differences will disappear since we keep choosing sufficiently large numbers $T_{1},l_{1},T_{2},l_{2},T_{3},l_{3},\dots$ to offset the effects by the differences in the early stages.
+In a similar way, we can construct $\mathcal{A}_{k}$ for any $k\in\mathbb{N}$ inductively.
+For $k>1$ we choose sufficiently large numbers $l_{k}>0$ and $T_{k}>0$ such that $\beta_{0}(a_{T_{k}})=-\ln l_{k}$.
+For each cube $S\in\mathcal{A}_{k-1}$, by Proposition 3.3, we know that
+$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
 and
-$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
+$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
 where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for $\theta_{2}>0$ as defined in Lemma 4.1.
@@
 $\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset,$
-we devide the region $\mathcal{F}_{q}\cap S$ into small cubes of side length
+we divide the region $\mathcal{F}_{q}\cap S$ into small cubes of side length
 $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$
 where $\theta_{1}>0$ is the constant defined in Lemma 4.1.
-Then we collect those cubes which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$, and denote the corresponding collection by
-$\mathcal{G}_{k,q,S}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset\}.$
+Then we collect those cubes $R\subset\mathcal{F}_{q}\cap S$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$, and denote the corresponding collection by
+$\mathcal{G}_{k,q,S}=\{R:R\subset\mathcal{F}_{q}\cap S,\;R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset\}.$
 Note that $\theta_{1}\cdot\exp(-\nu_{0}(a_{T_{k}}))/10$ is smaller than the minimum side length of the rectangle $B_{\mathbf{H}_{e}}(\theta_{1},T_{k})$, so by Lemma 4.1, we know that each cube in $\mathcal{G}_{k,q,S}$ is contained in
 $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$
-where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
-$\mathcal{H}_{k,S}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R.$
+where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$.
+For any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ $(q\in\mathbf{F}_{e}(\mathbb{Q}))$ in $\mathcal{F}_{\mathbf{H}_{e}}$ such that
+$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})=\emptyset,$
+we set $\mathcal{G}_{k,q,S}=\emptyset$.
+Let
+$\mathcal{H}_{k,S}=\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{k,q,S}}R.$
 Then we have
 $l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
-$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
-$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R)$
+$\asymp m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
+$\leq m_{\mathbf{H}_{e}}(\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{k,q,S}}R)$
 $\leq m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
-$\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
-and
-$m_{\mathbf{H}_{e}}(\mathcal{H}_{k,S})\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S).$
+$\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
+and hence
+$m_{\mathbf{H}_{e}}(\mathcal{H}_{k,S})\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S).$
 Note that each cube in $\mathcal{H}_{k,S}$ is contained in $S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$. Let
@@
 By the computations in §3, we can choose a sufficiently small number $\tilde{\delta}_{0}>0$ such that
-$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
+$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
 is a homeomorphism, and the subsets in the collection
@@
 are disjoint by this homeomorphism as
-$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{k}).$
+$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{k})\textup{ and }a_{T_{k}}\cdot q\Gamma\in\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\;(\forall q\in\mathcal{H}_{k,S}).$
 We denote by
@@
 if $T_{k}>0$ and $l_{k}>0$ are sufficiently large.
-Note that $\Phi^{0}(\mathbf{F}_{e})$ may not be empty, and some of these cubes may be outside $S$, and here we collect only those cubes which are inside $S$.
-Thus, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{k}(S)$ inside $S$.
-We denote by $\mathcal{A}_{k}$ the collection of all these disjoint cubes constructed from $\mathcal{F}_{k}(S)$ inside $S$ where $S$ ranges over all elements in $\mathcal{A}_{k-1}$.
-In this manner, we obtain a sequence $\left\{l_{k}\right\}_{k\in\mathbb{N}}$ of sufficiently large numbers $l_{k}$ with
-$l_{k+1}\gg l_{k}\;(\forall k\in\mathbb{N})$
-and a tree-like structure $\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ of finite collections of cubes. Using the notation in Theorem 4.3, we have
-$d_{k}(\mathcal{A})\sim l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
+Note that $\Phi^{0}(\mathbf{F}_{e})$ may not be empty, and some of the cubes in $\mathcal{P}_{k,S}$ may be outside $S$, and here we collect only those cubes in $\mathcal{P}_{k,S}$ which are inside $S$.
+Thus, we obtain a family of disjoint cubes inside $S$ constructed from $\mathcal{P}_{k,S}$.
+We denote by $\mathcal{A}_{k}$ the collection of all the disjoint cubes inside $S$ constructed from $\mathcal{P}_{k,S}$ where $S$ ranges over all elements in $\mathcal{A}_{k-1}$.
+In this manner, we obtain an increasing sequence of sufficiently large numbers $\left\{l_{k}\right\}_{k\in\mathbb{N}}$ and a tree-like collection $\mathcal{A}=\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ of cubes in $U_{0}$.
+Using the notation in Theorem 4.2, we have
+$d_{k}(\mathcal{A})\asymp l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
 where $d_{k}(\mathcal{A})$ is the diameter of the family $\mathcal{A}_{k}$. Moreover, one can compute that
-$\begin{cases}\Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})=\emptyset\\ \Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}\cdot\prod_{\alpha\in\Phi^{0}(\mathbf{F}_{e})}l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})\neq\emptyset.\end{cases}$
-Now let $\mathbf{A}_{\infty}=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. By applying Theorem 4.3, we can compute (assuming that $l_{k+1}$ is much larger than $l_{k}$ for any $k\in\mathbb{N}$) that
+$\begin{cases}\Delta_{k}(\mathcal{A})\asymp l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})=\emptyset\\ \Delta_{k}(\mathcal{A})\asymp l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}\cdot\prod_{\alpha\in\Phi^{0}(\mathbf{F}_{e})}l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})\neq\emptyset.\end{cases}$
+Now let $\mathbf{A}_{\infty}=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. By Theorem 4.2, we can compute (assuming that $l_{k+1}$ is much larger than $l_{k}$ for any $k\in\mathbb{N}$) that
 $\dim_{H}(\mathbf{A}_{\infty})\geq$ $\dim_{H}X-\limsup_{k\to\infty}\frac{\sum_{i=0}^{k}\log(\Delta_{i}(\mathcal{A}))}{\log(d_{k+1}(\mathcal{A}))}$
 $=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
-$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
+$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})(\tau+\epsilon)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
+By the definition of $\tau=\tau(\psi)$, there exists a divergent sequence $\{s_{k}\}_{k\in\mathbb{N}}\subset\mathbb{R}_{+}$ such that
+$\tau=\tau(\psi)=\lim_{k\to\infty}\left(-\frac{\ln\psi(s_{k})}{s_{k}}\right).$
+Now we choose the sequence $\{l_{k}\}_{k\in\mathbb{N}}$ in the construction of $\mathbf{A}_{\infty}$ so that the sequence $\{t_{k}\}_{k\in\mathbb{N}}$ equals the sequence $\{s_{k}\}_{k\in\mathbb{N}}$ defined as above.
+Then we can prove the following
 Lemma 4.3.
-We have $\mathbf{A}_{\infty}\subset S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+We have $\mathbf{A}_{\infty}\subset S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
 Proof.
-By the construction, for any $p\in\mathbf{A}_{\infty}$, there exists a sequence of rational points
+By the construction, for any $p\in\mathbf{A}_{\infty}$, there exists a sequence of rational elements
 $q_{k}\in S_{K_{1},K_{2},K_{3},K_{4},j}(S_{k-1},A_{l_{k}},B_{l_{k}}),\quad S_{k-1}\in\mathcal{A}_{k-1}$
@@
 $a_{t_{k}}\cdot p\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot(a_{t_{k}}\cdot q_{k}).$
-Note that by definition, the denominator of the rational element $a_{t_{k}}\cdot q_{k}$ is equal to the co-volume of the lattice $a_{t_{k}}\cdot q_{k}\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$ and
-$d(a_{t_{k}}\cdot q_{k})=e^{\beta_{0}(a_{t_{k}})\cdot\dim V_{\beta_{0}}}\cdot d(q_{k}).$
-By Lemma 2.6, we deduce that
-$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\sim\delta(\rho(a_{t_{k}}\cdot q_{k})\mathbb{Z}^{d})\lesssim d(a_{t_{k}}\cdot q_{k})^{\frac{1}{\dim V_{\beta_{0}}}}$
-$\lesssim e^{\beta_{0}(a_{t_{k}})}\cdot l_{k}\leq e^{-(\tau+\epsilon)t_{k}}$
-and $p\in S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. This completes the proof of the lemma. ∎
+Note that by definition, the height of the rational element $a_{t_{k}}\cdot q_{k}$ is equal to the co-volume of the lattice $\rho(a_{t_{k}}\cdot q_{k})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$ and
+$0pt(a_{t_{k}}\cdot q_{k})=e^{\beta_{0}(a_{t_{k}})\cdot\dim V_{\beta_{0}}}\cdot d(q_{k}).$
+By Lemma 2.6, we deduce that for any $k\in\mathbb{N}$
+$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\asymp\delta(\rho(a_{t_{k}}\cdot q_{k})\mathbb{Z}^{d})\lesssim 0pt(a_{t_{k}}\cdot q_{k})^{\frac{1}{\dim V_{\beta_{0}}}}$
+$\lesssim e^{\beta_{0}(a_{t_{k}})}\cdot l_{k}\leq e^{-(\tau+\epsilon)t_{k}}.$
+Note that by our choice of $\{t_{k}\}_{k\in\mathbb{N}}$, for $\epsilon>0$ and for sufficiently large $k\in\mathbb{N}$,
+$\frac{\ln\psi(t_{k})}{t_{k}}\leq\tau+\frac{\epsilon}{2}\implies\psi(t_{k})\geq e^{-(\tau+\epsilon/2)t_{k}}.$
+So we have
+$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\lesssim e^{-(\tau+\epsilon)t_{k}}\leq\psi(t_{k})\cdot e^{-\frac{\epsilon}{2}t_{k}}$
+and by definition $p\in S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. This completes the proof of the lemma. ∎
 By Lemma 4.3 and the computation for $\dim_{H}\mathbf{A}_{\infty}$, we have
-$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})(\tau+\epsilon)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
 By taking $\epsilon\to 0$, we obtain
-$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
-Using the same argument as in [17, Section 10], we conclude that
-$\dim_{H}(S_{\tau}^{c})\geq$ $\dim\mathbf{G}(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
-Note that by Lemma 4.2, we have $\nu_{0}(a_{1})>0$ and the argument here works for all $0\leq\tau<\beta_{0}(a_{-1})$.
-This completes the proof of Theorem 1.3 (provided that $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$).
-∎
-5. Proof of Theorem 1.3: $\{a_{t}\}_{t\in\mathbb{R}}$ in a proper normal $\mathbb{Q}$-subgroup of $\mathbf{G}$
-In this section, we discuss the case when the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, and explain how to modify the arguments in §3 and §4 and give a proof of Theorem 1.3.
-We first discuss the ergodic properties of group actions on homogeneous spaces, and one can refer to e.g.
-[29, 35, 37] for details.
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
+By definition, if an element $g\in\mathbf{G}(\mathbb{R})$ belongs to $S_{\rho}(\psi)^{c}$, then for any $h\in\mathbf{P}_{0}(\mathbb{R})$, $h\cdot g$ also belongs to $S_{\rho}(\psi)^{c}$.
+Using the same argument as in [25, §10], we can conclude that
+$\dim_{H}(S_{\rho}(\psi)^{c})\geq$ $\dim\mathbf{P}_{0}(\mathbb{R})+\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$\geq$ $\dim\mathbf{G}(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
+This completes the proof of Theorem 1.4 in the case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. ∎
+5. Proof of Theorem 1.4: The other cases
+In this section, we discuss the case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, and explain how to modify the arguments in §3 and §4 to prove Theorem 1.4.
+We study the case $\tau(\psi)=\beta_{0}(a_{-1})$ at the end of this section.
+We first discuss the ergodic properties of group actions on homogeneous spaces, for which one may refer to [59, 66, 69].
 Let $\mathbf{G}_{i}$ $(1\leq i\leq k)$ be the $\mathbb{Q}$-simple factors of $\mathbf{G}$.
@@
 By Godement compactness criterion (Cf.
-[7, Theorem 11.6]) and [8, 6.21], every simple factor of $\mathbf{G}_{i}(\mathbb{R})^{0}$ is not compact.
+[9, Theorem 11.6]) and [10, 6.21], every simple factor of $\mathbf{G}_{i}(\mathbb{R})^{0}$ is not compact.
 Let $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ be the projection of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$.
@@
 $\Gamma\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}=\prod_{i=1}^{s}\Gamma_{i}.$
-In the follwing, we fix these lattices $\Gamma_{i}$’s and $\Gamma$ and denote by
+In the follwing, we fix the lattices $\Gamma_{i}$’s and $\Gamma$ and denote by
 $\tilde{\Gamma}=\Gamma\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}.$
-Therefore, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ is mixing.
+So the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ is mixing.
 Now all the arguments in §2, §3 and §4 can be carried over to $\tilde{\mathbf{G}}$ and the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ almost verbatim. Indeed, denote by
@@
 which is a minimal parabolic $\mathbb{Q}$-subgroup of $\tilde{\mathbf{G}}$, and let $\overline{\tilde{\mathbf{P}}}_{0}=\overline{\mathbf{P}}_{0}\cap\tilde{\mathbf{G}}$.
-Let $U$ be an open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and define
+Let $U$ be a small open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and define
 $\tilde{S}(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
@@
 $\tilde{\mathbf{H}}_{e}=\mathbf{H}_{e}\cap\tilde{\mathbf{G}},\quad\tilde{\mathbf{F}}_{e}=\mathbf{F}_{e}\cap\tilde{\mathbf{G}}$.
-In the following, we need results in the reduction theory about $\tilde{\mathbf{G}}(\mathbb{R})/\tilde{\Gamma}$.
-Let $\tilde{K}$ be a maximal compact subgroup in $\tilde{\mathbf{G}}(\mathbb{R})$.
-Let $\tilde{M}$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\tilde{\mathbf{G}}(\mathbb{R})}(\tilde{\mathbf{T}}(\mathbb{R}))$ (= the centralizer of $\tilde{\mathbf{T}}(\mathbb{R})$ in $\tilde{\mathbf{G}}(\mathbb{R})$).
-Denote by
+[…]
 $\tilde{T}_{\eta}=\{a\in\tilde{\mathbf{T}}(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\tilde{\Delta}\}$
-where $\tilde{\Delta}$ is the set of positive $\mathbb{Q}$-simple roots in $\tilde{\mathbf{G}}$.
-A Siegel set in $\tilde{\mathbf{G}}(\mathbb{R})$ is a subset of the form $\tilde{S}_{\eta,\Omega}=\tilde{K}\cdot\tilde{T}_{\eta}\cdot\tilde{\Omega}$ for some $\eta\in\mathbb{R}$ and a relatively compact open subset $\tilde{\Omega}$ containing identity in $\tilde{M}\cdot R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, and the group $\tilde{\mathbf{G}}(\mathbb{R})$ can be written as
+[…]
 $\tilde{\mathbf{G}}(\mathbb{R})=\tilde{S}_{\eta,\Omega}\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}$
-for some Siegel set $\tilde{S}_{\eta,\Omega}$ and some finite subset $\tilde{\mathcal{K}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$.
-Moreover, the finite set $\tilde{\mathcal{K}}$ satisfies the property that
+[…]
 $\tilde{\mathbf{G}}(\mathbb{Q})=\tilde{\mathbf{P}}_{0}(\mathbb{Q})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}$.
-Denote by $\tilde{\mathcal{K}}=\{\tilde{x}_{j}\}_{j\in\tilde{J}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$ and we may assume that $e\in\tilde{\mathcal{K}}$.
+[…]
+The proof of the following lemma is similar to those of Corollary 2.3 and Lemma 2.4.
 Lemma 5.1.
@@
     An element $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ is rational if and only if
-    $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})\cap(\tilde{\mathbf{H}}_{1}(\mathbb{R})\cdot\tilde{\mathbf{P}}_{0}(\mathbb{R})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}).$
+[…]
 Definition 5.2.
@@
 for some $h\in\tilde{\mathbf{H}}_{e}(\mathbb{R})$, $p\in\tilde{\mathbf{P}}_{0}(\mathbb{R})$ and $\gamma\in\tilde{\Gamma}$.
-Let $\tilde{\mathfrak{a}}$ be the Lie algebra of $\tilde{\mathbf{T}}(\mathbb{R})$.
-For any $\tilde{x}_{j}\in\tilde{\mathcal{K}}$, any compact subset $K_{1}$ in $\ker(\beta_{0})\cap\tilde{\mathfrak{a}}$, any compact subset $K_{2}$ in $\tilde{\mathbf{H}}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, we define
+[…]
 $\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$
-to be the set of all rational points $q$ in $U$ such that
+to be the set of all rational elements $q$ in $U$ such that
   1. (1)
@@
   2. (2)
-    $q=a\cdot h\cdot m\cdot u\cdot\tilde{x}_{j}\cdot\gamma$ for some $a\in\exp(\tilde{\mathfrak{a}})$, $\pi_{\ker(\beta_{0})}(a)\in K_{1}$, $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\tilde{\Gamma}$.
+[…]
 Note that $\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)\neq\emptyset$ implies that $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$.
@@
 Then for any $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$, we define $m_{\tilde{\mathbf{H}}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ which is supported on $\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\tilde{\mathbf{H}_{e}}}$ via the product map
-$\tilde{\mathbf{H}}_{e}\times\{q\}\to\tilde{\mathbf{H}}_{e}\cdot q\subset R_{u}(\overline{\tilde{\mathbf{P}}}_{0}).$
+$\tilde{\mathbf{H}}_{e}(\mathbb{R})\times\{q\}\to\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q\subset R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R}).$
 We define
@@
 Proposition 5.3.
-Let $U$ be a small open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$.
-Then for any $j\in\tilde{J}$ with $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
-$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U)$
+[…]
+$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U)$
 and
-$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U).$
+$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U).$
 Here the implicit constants depend only on the compact subsets $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
-The analogues of Lemma 4.1 and Lemma 4.2 hold as well in $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
+The analogue of Lemma 4.1 holds as well in $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
 Lemma 5.4.
@@
 $\beta_{0}(a_{T})=-\ln l$
-for some $T>0$ and $l>1$. Let ~$\tilde{\mathcal{F}}_{q}=\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$ such that
-~$\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$
+for some $T>0$ and $l>1$. Let $\tilde{\mathcal{F}}_{q}=\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$ such that
+$\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$
 Then there exist $\theta_{1}>0$ and $\theta_{2}>0$ such that for any $p\in\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
 $B_{\tilde{\mathbf{H}}_{e}}(\theta_{1},T)\cdot p\cap U\subset\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$
-where $\tilde{K}_{2}=B_{\tilde{\mathbf{H}}_{1}}(\theta_{2})\cdot K_{2}$ and $B_{\tilde{\mathbf{H}}_{1}}(\theta_{1},T)=a_{-T}\cdot B_{\tilde{\mathbf{H}}_{e}}(\theta_{1})\cdot a_{T}$.
+where $\tilde{K}_{2}=B_{\tilde{\mathbf{H}}_{e}}(\theta_{2})\cdot K_{2}$ and $B_{\tilde{\mathbf{H}}_{e}}(\theta_{1},T)=a_{-T}\cdot B_{\tilde{\mathbf{H}}_{e}}(\theta_{1})\cdot a_{T}$.
 Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
-Lemma 5.5.
-We have
-$\max\{\alpha(a_{1}):\alpha\in\Phi(\tilde{\mathbf{F}}_{e})\}>0.$
-Consequently, using the same argument as in §4, we can conclude that
-$\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))\geq\dim\tilde{\mathbf{G}}-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\tilde{\nu}_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
-where $\tilde{\nu}_{0}$ is the $\mathbb{Q}$-root in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$ such that
+Consequently, using the same argument as in §4, we can conclude that for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))\geq\dim\tilde{\mathbf{G}}-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\beta_{0}(a_{-1})\tilde{\nu}_{0}(a_{1})}$
+where $\tilde{\nu}_{0}$ is any $\mathbb{Q}$-root in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$ such that
 $\tilde{\nu}_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\tilde{\mathbf{P}}}_{0}))\}.$
-Note that $\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$ all commute with $\{a_{t}\}_{t\in\mathbb{R}}$, and an element $g\in S_{\tau}^{c}$ if and only if $h\cdot g\in S_{\tau}^{c}$ for any $h\in\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$.
+Note that $\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$ commutes with $\{a_{t}\}_{t\in\mathbb{R}}$, and an element $g\in S_{\rho}(\psi)^{c}$ if and only if $h\cdot g\in S_{\rho}(\psi)^{c}$ for any $h\in\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$.
 Therefore, we have
-$\dim_{H}(S_{\tau}^{c})=$ $\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))+\sum_{i=s+1}^{k}\dim\mathbf{G}_{i}(\mathbb{R})$
-$\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
+$\dim_{H}(S_{\rho}(\psi)^{c})=$ $\dim_{H}(S_{\rho}(\psi)^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))+\sum_{i=s+1}^{k}\dim\mathbf{G}_{i}(\mathbb{R})$
+$\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\nu_{0}(a_{1})\beta_{0}(a_{-1})}$
 where $\nu_{0}(a_{1}):=\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}$. Here we use the fact that $\nu_{0}(a_{1})=\tilde{\nu}_{0}(a_{1})$ and
@@
 One can see from the arguments in §4 and §5 that we actually prove that for any open bounded subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$
-$\dim_{H}(S_{\gamma}^{c}\cap U)\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
-Note that $\mathbf{P}_{0}\cdot R_{u}(\overline{\mathbf{P}}_{0})$ is a Zariski open dense subset in $\mathbf{G}$, and if $x\in S_{\gamma}^{c}\cap U$, then for any $g\in\mathbf{P}_{0}(\mathbb{R})$, $g\cdot x$ is also an element in $S_{\gamma}^{c}$.
-This implies that for any open bounded subset $\tilde{U}\subset\mathbf{G}(\mathbb{R})$
-$\dim_{H}(S_{\gamma}^{c}\cap\tilde{U})\geq\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
-6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
-In this section, we prove Theorem 1.4. We will compute an upper bound of $\dim_{H}S_{\gamma}^{c}$ by constructing open covers of the subset $S_{\gamma}^{c}$. Recall that
-$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap U)\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\nu_{0}(a_{1})\beta_{0}(a_{-1})}.$
+Note that $\mathbf{P}_{0}\cdot R_{u}(\overline{\mathbf{P}}_{0})$ is a Zariski open dense subset in $\mathbf{G}$, and for any open bounded subset $\tilde{U}\subset\mathbf{G}(\mathbb{R})$, one can find open subsets $U_{1}$ and $U_{2}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and $\mathbf{P}_{0}(\mathbb{R})$ respectively such that
+$U_{2}\cdot U_{1}\subset\tilde{U}$.
+Moreover, by definition, if $x\in S_{\rho}(\psi)^{c}$, then for any $g\in\mathbf{P}_{0}(\mathbb{R})$, $g\cdot x$ is also an element in $S_{\rho}(\psi)^{c}$. This implies that
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap\tilde{U})\geq\dim U_{2}+\dim_{H}(S_{\rho}(\psi)^{c}\cap U_{1})\geq\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
+Finally, we study the case $\tau(\psi)=\beta_{0}(a_{-1})$ and conclude the proof of Theorem 1.4.
+Proposition 5.6.
+Let $\mathbf{G}$ be a connected semisimple algebraic group, $\rho$ a $\mathbb{Q}$-rational irreducible representation of $\mathbf{G}$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$ and $\beta_{0}$ the highest weight defined as in §2.
+Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $\tau(\psi)=\beta_{0}(a_{-1})$.
+Then
+  1. (1)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
+  2. (2)
+    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
+    $\dim_{H}S_{\rho}(\psi)^{c}\geq\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
+Proof.
+Suppose that $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded. Then there exists $C>0$ such that
+$\psi(t)\leq C\cdot e^{-\beta_{0}(a_{-1})t}\quad(t>0).$
+Let $g\in\mathbf{G}(\mathbb{R})$. Since $\beta_{0}(a_{-1})>0$ is the fastest contracting rate of the $\{a_{t}\}_{t\in\mathbb{R}}$-action on $V$, we have
+$\|\rho(a_{t})v\|\geq e^{-\beta_{0}(a_{-1})t}\|v\|\quad(\forall v\in\rho(g)\cdot\mathbb{Z}^{d}\setminus\{0\})$
+and
+$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq e^{-\beta_{0}(a_{-1})t}\cdot\delta(\rho(g)\cdot\mathbb{Z}^{d}).$
+One can find a constant $C_{g}>0$ depending on $g$ such that
+$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq C_{g}\cdot\psi(t)\quad(t>0)$
+which shows that $g\in S_{\rho}(\psi)$ and $S_{\rho}(\psi)=\mathbf{G}(\mathbb{R})$. This proves the first claim.
+Now suppose that $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded. We first consider the case where $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. We may write
+$\psi(t)\cdot e^{\beta_{0}(a_{-1})t}=e^{h(t)}$
+where $h(t)$ is not bounded from above. Since $\tau(\psi)=\beta_{0}(a_{-1})$, we have
+$\limsup_{t\to\infty}\frac{h(t)}{t}=0.$
+Let $\{s_{k}\}_{k\in\mathbb{N}}$ be a divergent sequence in $\mathbb{R}_{+}$ such that $\lim_{k\to\infty}h(s_{k})=\infty.$ Then
+$\lim_{k\to\infty}\frac{h(s_{k})}{s_{k}}=0.$
+In §4, we construct a Cantor-type subset $\mathbf{A}_{\infty}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ with parameters $l_{k}$ and $t_{k}$ $(k\in\mathbb{N})$.
+Here to prove the proposition with $\tau(\psi)=\beta_{0}(a_{-1})$, we may take $l_{k}=e^{h(s_{k})/2}$, $t_{k}=s_{k}$ and $\tau_{k}=\beta_{0}(a_{-1})-h(s_{k})/(2s_{k})$ $(k\in\mathbb{N})$.
+Then
+$t_{k}=\frac{\ln l_{k}}{\beta_{0}(a_{-1})-\tau_{k}}.$
+Using the same argument in Lemma 4.3, one can prove that for any $p\in\mathbf{A}_{\infty}$
+$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\lesssim e^{-\tau_{k}\cdot t_{k}}=e^{-\beta_{0}(a_{-1})t_{k}+h(t_{k})/2}=\psi(t_{k})\cdot e^{-h(t_{k})/2}\quad(k\in\mathbb{N})$
+and $\mathbf{A}_{\infty}\subset S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Theorem 4.2, we may estimate the Hausdorff dimension of $\mathbf{A}_{\infty}$ and obtain
+$\dim_{H}S_{\rho}(\psi)^{c}\geq\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
+If the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, then we may repeat the arguments in this section and reduce the problem in the homogeneous space $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
+This completes the proof of the proposition.
+∎
+6. An upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}$
+In this section, we prove Theorem 1.5.
+We will first compute an upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by constructing open covers of the subset $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, and then derive Theorem 1.5 from the upper bound of $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+We denote by
+$\mathcal{C}:=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}w\mathbf{F}_{w}(\mathbb{R}).$
 Proposition 6.1.
@@
 $\delta(\rho(g)\cdot\mathbb{Z}^{d})\leq r$
-for some $g\in\mathbf{G}(\mathbb{R})$ and $r>0$. Then there exist $\tilde{g}\in\mathcal{C}_{\bar{I}}$ and a discrete primitive subgroup
+for some $g\in\mathbf{G}(\mathbb{R})$ and $r>0$. Then there exist $\tilde{g}\in\mathcal{C}$ and a discrete primitive subgroup of rank $\dim V_{\beta_{0}}$
 $\Lambda_{g}\subset\rho(g)\cdot\mathbb{Z}^{d}$
@@
 $\mathbf{G}(\mathbb{R})=S_{\eta,\Omega}\cdot\mathcal{K}\cdot\Gamma$
-for some Siegel set $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ where
+for some Siegel set $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, where
 $T_{\eta}=\{a\in T(\mathbb{R}):\lambda(a)\leq\eta,\;\lambda\textup{ a simple root in }\Delta\}$
@@
 $\{aua^{-1}:a\in T_{\eta}\textup{ and }u\in\Omega\}$
-is compact. So we can write
+is relatively compact. So we can write
 $g=\tilde{k}\cdot a\cdot x\cdot\gamma$
-for some $\tilde{k}$ in a fixed compact subset in $\mathbf{G}(\mathbb{R})$.
+for $\tilde{k}=k\cdot(a\cdot u\cdot a^{-1})$ in a fixed compact subset in $\mathbf{G}(\mathbb{R})$.
 Now since $\delta(\rho(g)\cdot\mathbb{Z}^{d})\leq r$, there exists $y\in\mathbb{Z}^{d}\setminus\{0\}$ such that $\rho(\tilde{k}\cdot a\cdot x)\cdot y$ is a shortest vector in $\rho(g)\cdot\mathbb{Z}^{d}$ and
-$\|\rho(a\cdot x)\cdot y\|\sim\|\rho(\tilde{k}\cdot a\cdot x)\cdot y\|\leq r.$
+$\|\rho(a\cdot x)\cdot y\|\asymp\|\rho(\tilde{k}\cdot a\cdot x)\cdot y\|\leq r.$
 We write
 $\rho(x)\cdot y=\sum_{\text{weight }\beta}y_{\beta}$
-according to the decomposition of the complex vector space $V=\oplus_{\beta}V_{\beta}$ in $\rho$ into weight spaces $V_{\beta}$ relative to $\mathbf{T}$, where $y_{\beta}\in V_{\beta}$.
-Let $\tilde{\beta}$ be a weight among the weights $\beta$’s in the summation such that $y_{\tilde{\beta}}\neq 0$.
-Then by the structure of the representation $\rho$, we compute that
+according to the decomposition of $V=\oplus_{\beta}V_{\beta}$ in $\rho$ into the weight spaces $V_{\beta}$ relative to $\mathbf{T}$, where $y_{\beta}\in V_{\beta}$.
+Let $\tilde{\beta}$ be a weight among the weights $\beta$’s in the summation above such that $y_{\tilde{\beta}}\neq 0$.
+Then
 (2) $\|\rho(a\cdot x)\cdot y\|=\|e^{\tilde{\beta}(a)}y_{\tilde{\beta}}+\cdots\|\lesssim r.$
-Note that $y\in\mathbb{Z}^{d}$ and $x\in\mathcal{K}$, and there is a positive lower bound for $\|y_{\tilde{\beta}}\|$ depending only on $\rho$ and $\mathcal{K}$.
+Note that $y\in\mathbb{Z}^{d}$ and $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, and there is a positive lower bound for $\|y_{\tilde{\beta}}\|$ depending only on $\rho$ and $\mathcal{K}$.
+So by equation (2), we have
+$e^{\tilde{\beta}(a)}\lesssim r.$
 On the other hand, there exists a unique discrete primitive subgroup $\tilde{\Lambda}\subset\mathbb{Z}^{d}$ such that $\rho(x)\cdot\tilde{\Lambda}$ is a discrete and Zariski-dense subgroup in $V_{\beta_{0}}$, and for any $w\in\tilde{\Lambda}\setminus\{0\}$ we have
-(2) $\|\rho(\tilde{k}\cdot a\cdot x)\cdot w\|\sim\|\rho(a)\cdot\rho(x)\cdot w\|=e^{\beta_{0}(a)}\|\rho(x)\cdot w\|.$
-Comparing equations (1) and (2) and using the fact that $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, $a\in T_{\eta}$ and the relation between $\tilde{\beta}$ and $\beta_{0}$ according to the structure of the representation $\rho$, one can deduce that
-$e^{\beta_{0}(a)}\lesssim_{\eta}e^{\tilde{\beta}(a)}\lesssim r$
-and for any $w\in\tilde{\Lambda}\setminus\{0\}$
+(3) $\|\rho(\tilde{k}\cdot a\cdot x)\cdot w\|\asymp\|\rho(a)\cdot\rho(x)\cdot w\|=e^{\beta_{0}(a)}\|\rho(x)\cdot w\|.$
+By the condition $a\in T_{\eta}$ and the relation between $\tilde{\beta}$ and $\beta_{0}$ according to the structure of the representation $\rho$, one can compute that
+$e^{\beta_{0}(a)}\lesssim_{\eta}e^{\tilde{\beta}(a)}\lesssim r.$
+Consequently, from equation (3) one can deduce that for any $w\in\tilde{\Lambda}\setminus\{0\}$
 $\|\rho(\tilde{k}\cdot a\cdot x)w\|\lesssim r\cdot\|\rho(x)w\|.$
-Let $\Lambda_{g}=\rho(g)(\rho(\gamma^{-1})\tilde{\Lambda})$. Then we have
+[…]
 $\Lambda_{g}\subset\rho(g)\cdot\mathbb{Z}^{d},\quad\operatorname{Vol}(\Lambda_{g})\lesssim r^{\operatorname{rank}\Lambda_{g}}=r^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{k}^{-1})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
-By the Bruhat decomposition, we know that
-$\tilde{k}^{-1}\in\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
-By the fact that $\mathbf{P}_{\beta_{0}}$ stabilizes $V_{\beta_{0}}$, we conclude that there exists $\tilde{g}\in\mathcal{C}_{\bar{I}}$ such that
-$\tilde{k}^{-1}\in\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot\tilde{g}\textup{ and }\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
-This completes the proof of the proposition. ∎
-Let $\epsilon_{0}>0$ which we will determine later.
-Let $g$ be an element in $S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-Then by definition, there exists a sequence $\{t_{k}\}\subset\mathbb{N}$ such that
-$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq\epsilon_{0}\cdot e^{-\tau t_{k}}.$
-By Proposition 6.1, for each $k\in\mathbb{N}$, there exist $\tilde{g}_{k}\in\mathcal{C}_{\bar{I}}$ and a discrete subgroup $\Lambda_{k}\subset\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d}$ such that
-$\operatorname{Vol}(\Lambda_{k})\lesssim(\epsilon_{0}e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g}_{k})\cdot\Lambda_{k}\textup{ is Zariski dense in }V_{\beta_{0}}.$
-Let $\tilde{g}_{k}=w_{i,k}\cdot f_{i,k}\in\mathcal{C}_{\bar{I}}$ where $w_{i,k}\in\{w_{i}\}_{i\in\bar{I}}$ and $f_{i,k}\in\mathbf{F}_{w_{i,k}}(\mathbb{R})$. Then it implies that
-$\tilde{q}_{k}:=w_{i,k}\cdot f_{i,k}\cdot a_{t_{k}}\cdot g=w_{i,k}\cdot a_{t_{k}}\cdot(a_{-t_{k}}f_{i,k}a_{t_{k}})\cdot g$
-$q_{k}:=(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\tilde{q}_{k}=w_{i,k}(a_{-t_{k}}f_{i,k}a_{t_{k}})g$
-are rational elements in $\mathbf{G}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, we can compute that
-$d(q_{k})=e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}\cdot d(\tilde{q}_{k})$
-$d(\tilde{q}_{k})\cdot e_{V_{\beta_{0}}}=\rho_{\beta_{0}}(\tilde{g}_{k})\cdot\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right),\quad\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}=\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right)/d(\tilde{q}_{k})$
-$\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|=$ $\operatorname{Vol}(\Lambda_{k})/d(\tilde{q}_{k})$
-$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
-where $e_{V_{\beta_{0}}}=\bigwedge^{\dim V_{\beta_{0}}}V_{\beta_{0}}$ is the unit vector which represents the vector space $V_{\beta_{0}}$ in the $\dim V_{\beta_{0}}$-exterior product space of $V$, and $\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}$ is the vector constructed from a $\mathbb{Z}$-basis of $\Lambda_{k}$ and represents the lattice $\Lambda_{k}$.
-The implicit constants here depend only on $\mathbf{G}$, $\Gamma$ and $\rho$.
+We know that
+$\tilde{k}^{-1}\in\mathbf{G}(\mathbb{R})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w\cdot\mathbf{F}_{w}(\mathbb{R}).$
+So there exists $\tilde{g}\in\mathcal{C}$ such that
+$\tilde{k}^{-1}\in\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot\tilde{g}$.
+By the fact that $\mathbf{P}_{\beta_{0}}$ stabilizes $V_{\beta_{0}}$, we conclude that
+$\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
+Note that all the implicit constants above depend only on $\mathbf{G}$, $\Gamma$ and $\rho$. This completes the proof of the proposition. ∎
+Let $g$ be an element in $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+Then from Definition 1.1, one can deduce that there exists a divergent sequence $\{t_{k}\}\subset\mathbb{R}_{+}$ such that
+$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq\psi(t_{k}).$
+Let $\epsilon>0$ be a sufficiently small number. By the definition of $\tau=\tau(\psi)$, we have
+$\psi(t)\leq C_{\epsilon}\cdot e^{-(\tau-\epsilon)t}\quad(\forall t>0)$
+for some $C_{\epsilon}>0$ depending only on $\epsilon$. Hence
+$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq C_{\epsilon}\cdot e^{-(\tau-\epsilon)t_{k}}.$
+Here by adding a multiplicative constant in the inequality above, we may assume that $t_{k}\in\mathbb{N}$.
+By Proposition 6.1, for each $k\in\mathbb{N}$, there exist $\tilde{g}_{k}\in\mathcal{C}$ and a discrete subgroup $\Lambda_{k}\subset\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d}$ of rank $\dim V_{\beta_{0}}$ such that
+$\operatorname{Vol}(\Lambda_{k})\lesssim(C_{\epsilon}\cdot e^{-(\tau-\epsilon)t_{k}})^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g}_{k})\cdot\Lambda_{k}\textup{ is Zariski dense in }V_{\beta_{0}}.$
+Since $_{\mathbb{Q}}\overline{\mathcal{W}}$} is finite, by passing to a subsequence, we may write $\tilde{g}_{k}=w\cdot f_{k}\in\mathcal{C}$ where $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ is fixed for all $k\in\mathbb{N}$ and $f_{k}\in\mathbf{F}_{w}(\mathbb{R})$.
+Then by definition
+$\tilde{q}_{k}:=\tilde{g}_{k}\cdot a_{t_{k}}\cdot g=w\cdot f_{k}\cdot a_{t_{k}}\cdot g=w\cdot a_{t_{k}}\cdot(a_{-t_{k}}f_{k}a_{t_{k}})\cdot g$
+$q_{k}:=(wa_{-t_{k}}w^{-1})\cdot\tilde{q}_{k}=w\cdot(a_{-t_{k}}f_{k}a_{t_{k}})\cdot g$
+are rational elements in $\mathbf{G}(\mathbb{R})$.
+In the $\dim V_{\beta_{0}}$-exterior product space of $V$, let $e_{V_{\beta_{0}}}=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}$ be the unit vector which represents the vector space $V_{\beta_{0}}$ (see §1), and $\wedge^{\dim V_{\beta_{0}}}\Lambda_{k}$ the vector in $\bigwedge^{\dim V_{\beta_{0}}}V$ which is constructed from a $\mathbb{Z}$-basis of $\Lambda_{k}$ and represents the lattice $\Lambda_{k}$.
+Then we can compute that
+$0pt(q_{k})=e^{\beta_{0}(wa_{-t_{k}}w^{-1})\cdot\dim V_{\beta_{0}}}\cdot 0pt(\tilde{q}_{k})$
+$0pt(\tilde{q}_{k})\cdot e_{V_{\beta_{0}}}=\rho_{\beta_{0}}(\tilde{g}_{k})\cdot\left(\wedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right),\quad\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}=\left(\wedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right)/0pt(\tilde{q}_{k})$
+$\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|=$ $\operatorname{Vol}(\Lambda_{k})/0pt(\tilde{q}_{k})$
+$\lesssim$ $(C_{\epsilon}\cdot e^{-(\tau-\epsilon)t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(wa_{-t_{k}}w^{-1})\cdot\dim V_{\beta_{0}}}/0pt(q_{k}).$
+The implicit constants here depend only on $\mathbf{G}$, $\Gamma$ and $\rho$. For convenience, we denote by
+$\lambda_{w}(\epsilon):=(\beta_{0}(wa_{-1}w^{-1})-(\tau-\epsilon))\cdot\dim V_{\beta_{0}}.$
 Definition 6.2.
-Fix an arbitrary $i\in\bar{I}$. For any $R>0$, define
-$E_{w_{i}}(R):=\{f\in\mathbf{F}_{w_{i}}(\mathbb{R}):\|\rho_{\beta_{0}}(w_{i}fw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq R\}.$
+Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. For any $R>0$, define
+$E_{w}(R):=\{f\in\mathbf{F}_{w}(\mathbb{R}):\|\rho_{\beta_{0}}(wfw^{-1})\cdot e_{V_{\beta_{0}}}\|\leq R\}.$
 We also define the following morphism
-$\Psi_{w_{i}}:\mathbf{F}_{w_{i}}(\mathbb{R})\to\bigwedge_{j=1}^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(w_{i}xw_{i}^{-1})\cdot e_{V_{\beta_{0}}}.$
-Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$. Hence $\Psi_{w_{i}}$ is an isomorphism onto its image.
+$\Psi_{w}:\mathbf{F}_{w}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w}(x)=\rho_{\beta_{0}}(wxw^{-1})\cdot e_{V_{\beta_{0}}}.$
+Note that $w\mathbf{F}_{w}w^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and the stabilizer of $\mathbb{C}\cdot e_{V_{\beta_{0}}}$ is equal to $\mathbf{P}_{\beta_{0}}$.
+Hence $\Psi_{w}$ is an isomorphism onto its image.
 From the discussion above, we know that
-$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim$ $\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|$
-$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
-and by definition of $E_{w_{i,k}}(R)$, we have
-$g=$ $(a_{-t_{k}}f_{i,k}^{-1}a_{t_{k}})\cdot w_{i,k}^{-1}\cdot q_{k}$
-$\in$ $\left(a_{-t_{k}}\cdot E_{w_{i,k}}\left(C_{1}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k})\right)\cdot a_{t_{k}}\right)\cdot w_{i,k}^{-1}\cdot q_{k}$
-for some constant $C_{1}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Moreover, by the structure of the representation $\rho$, we know that
-$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\Psi_{w_{i,k}}(f_{i,k}^{-1})\|\gtrsim\|e_{V_{\beta_{0}}}\|=1$
+$\|\rho_{\beta_{0}}(wf_{k}^{-1}w^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim$ $\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k})$
+and by definition of $E_{w}(R)$, we have
+$g=$ $(a_{-t_{k}}f_{k}^{-1}a_{t_{k}})\cdot w^{-1}\cdot q_{k}$
+$\in$ $\left(a_{-t_{k}}\cdot E_{w}\left(C_{1}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k})\right)\cdot a_{t_{k}}\right)\cdot w^{-1}\cdot q_{k}$
+for some constant $C_{1}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Moreover, by the structure of the representation $\rho_{\beta_{0}}$, we know that
+$\|\rho_{\beta_{0}}(wf_{k}^{-1}w^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\Psi_{w}(f_{k}^{-1})\|\gtrsim\|e_{V_{\beta_{0}}}\|=1$
 so
-$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
+$0pt(q_{k})\leq C_{2}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
 for some constant $C_{2}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Therefore, $q_{k}$ is a rational element in $\mathbf{G}(\mathbb{R})$ with
-$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
-and $w_{i,k}^{-1}\cdot q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-By passing to a subsequence, we may assume that $w_{i,k}$ is a fixed element in $\{w_{i}\}_{i\in\bar{I}}$ independent of $k$.
-In the following, we take $C_{0}=\max\{C_{1},C_{2}\}$.
+$0pt(q_{k})\leq C_{2}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
+and $w^{-1}\cdot q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. In the following, we take $C_{0}=\max\{C_{1},C_{2}\}$.
 Definition 6.3.
-For any $i\in\bar{I}$, define $\mathcal{E}_{i}(\tau)$ to be the subset of elements $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
-$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
-$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
-and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+For any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and any $\epsilon>0$, define $\mathcal{E}_{w,\epsilon}(\psi)$ to be the subset of elements $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
+$g\in\left(a_{-t_{k}}E_{w}(C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k}))a_{t_{k}}\right)w^{-1}q_{k}$
+$0pt(q_{k})\leq C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
+and $w^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
 Lemma 6.4.
-Let $i\in\bar{I}$.
-Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-Then there exists a constant $\theta_{i}>0$ depending only on $i$ such that $d(q)\geq\theta_{i}$.
+Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$.
+Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ with $w^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+Then there exists a constant $\theta_{w}>0$ depending only on $w$ such that $d(q)\geq\theta_{w}$.
 Proof.
-By definition, we know that $R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$ By Corollary 2.3, let
-$w_{i}^{-1}\cdot q=u\cdot v$
-for some $u\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$.
-Let $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ be a bounded fundamental domain of $\mathbf{F}_{w_{i}}(\mathbb{R})/(\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R}))$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$.
+By definition, we know that $R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w}\cdot\mathbf{F}_{w}.$ By Corollary 2.3, let
+$w^{-1}\cdot q=u\cdot v$
+for some $u\in\mathbf{H}_{w}(\mathbb{R})$ and $v\in\mathbf{F}_{w}(\mathbb{Q})$.
+Let $\tilde{\Omega}_{\mathbf{F}_{w}}$ be a bounded fundamental domain of $\mathbf{F}_{w}(\mathbb{R})/(\Gamma\cap\mathbf{F}_{w}(\mathbb{R}))$ in $\mathbf{F}_{w}(\mathbb{R})$.
 We may write
 $v=\tilde{v}\cdot\gamma$
-for some $\tilde{v}\in\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R})$. Then
-$q\cdot\mathbb{Z}^{d}=(w_{i}\cdot u\cdot v)\cdot\mathbb{Z}^{d}=(w_{i}uw_{i}^{-1})\cdot(w_{i}\tilde{v}\cdot\mathbb{Z}^{d}).$
-By the fact that $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ and $\tilde{v}$ is in the bounded subset $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, we can conclude that the co-volume of $(w_{i}\tilde{v}\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}$ has a lower bound depending only on $w_{i}$ and $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, and so does $q\cdot\mathbb{Z}^{d}$.
+for some $\tilde{v}\in\tilde{\Omega}_{\mathbf{F}_{w}}$ and $\gamma\in\Gamma\cap\mathbf{F}_{w}(\mathbb{R})$. Then
+$q\cdot\mathbb{Z}^{d}=(w\cdot u\cdot v)\cdot\mathbb{Z}^{d}=(wuw^{-1})\cdot(w\tilde{v}\cdot\mathbb{Z}^{d}).$
+By the fact that $w\mathbf{H}_{w}w^{-1}$ fixes every element in $V_{\beta_{0}}$ and $\tilde{v}$ is in the bounded subset $\tilde{\Omega}_{\mathbf{F}_{w}}$, we can conclude that the co-volume of $(w\tilde{v}\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}$ has a lower bound depending only on $w$ and $\tilde{\Omega}_{\mathbf{F}_{w}}$, and so does $q\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$.
 This completes the proof of the lemma.
 ∎
-Now we choose $\epsilon_{0}>0$ so that
-$\min_{i\in\bar{I}}\{\theta_{i}\}>C_{0}\cdot\epsilon_{0}^{\dim V_{\beta_{0}}}.$
 Lemma 6.5.
-If $\mathcal{E}_{i}(\tau)\neq\emptyset$ for some $i\in\bar{I}$, then $0\leq\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$.
+Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and $\epsilon>0$. If $\mathcal{E}_{w,\epsilon}(\psi)\neq\emptyset$, then $\tau-\epsilon\leq\beta_{0}(wa_{-1}w^{-1})$.
 Proof.
-Let $g\in\mathcal{E}_{i}(\tau)$. By definition, there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
-$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
-$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
-and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 6.4, we have
-$\theta_{i}\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
-If $\tau\geq\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$, then by taking $t_{k}\to\infty$, we obtain that
-$\theta_{i}\leq C_{0}\cdot(\epsilon_{0})^{\dim V_{\beta_{0}}}$
-which contradicts the choice of $\epsilon_{0}$. This completes the proof of the lemma. ∎
-From the discussion above, we obtain the following
+Let $g\in\mathcal{E}_{w,\epsilon}(\psi)$.
+By definition, there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
+$g\in\left(a_{-t_{k}}E_{w}(C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k}))a_{t_{k}}\right)w^{-1}q_{k}$
+$0pt(q_{k})\leq C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
+and $w^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 6.4, we have
+$\theta_{w}\leq C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}.$
+Then by taking $t_{k}\to\infty$, we obtain that $\tau-\epsilon\leq\beta_{0}(wa_{-1}w^{-1})$. This completes the proof of the lemma. ∎
 Proposition 6.6.
-We have
-$S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{i\in\bar{I}}\mathcal{E}_{i}(\tau).$
-In the following, for each $i\in\bar{I}$, we will consider the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ $(\mathcal{E}_{i}(\tau)\neq\emptyset)$ and compute an upper bound for the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$.
-These upper bounds will give an upper bound for $\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ by Proposition 6.6.
-For notational convenience, we may set $\epsilon_{0}=1$ as it does not affect the computations in the rest of this section.
-Now fix $i\in\bar{I}$.
-To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we first need to estimate the volume of the subset $E_{w_{i}}(R)$ in ${\mathbf{F}}_{w_{i}}(\mathbb{R})$, which can be written as
-$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))$
-where $B_{R}$ is the ball of radius $R$ centered at 00 in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ and $\mu_{\mathbf{F}_{w_{i}}}$ is the Haar measure on ${\mathbf{F}}_{w_{i}}(\mathbb{R})$.
+For any $\epsilon>0$, we have
+$S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathcal{E}_{w,\epsilon}(\psi).$
+Moreover, for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, $\beta_{0}(wa_{-1}w^{-1})\leq\beta_{0}(a_{-1})$.
+Proof.
+The inclusion
+$S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathcal{E}_{w,\epsilon}(\psi)$
+follows from the discussion above.
+Note that for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, $\beta_{w}(a):=\beta_{0}(waw^{-1})$ $(\forall a\in\mathbf{T})$ is a weight in the representation $\rho$, and $\beta_{0}$ is the highest weight.
+It implies that $\beta_{w}(a_{-1})\leq\beta_{0}(a_{-1})$.
+This completes the proof of the proposition.
+∎
+In the following, we will consider the subsets $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$ $(w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}})$ and compute upper bounds for the Hausdorff dimensions of these subsets.
+These upper bounds will give an upper bound for $\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ by Proposition 6.6.
+Note that we are only interested in the non-empty subsets $\mathcal{E}_{w,\epsilon}(\psi)\neq\emptyset$.
+So here we will only consider the elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ which satisfy $\tau\leq\beta_{0}(wa_{-1}w^{-1}).$ For the elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ with $\tau>\beta_{0}(wa_{-1}w^{-1})$, we may choose $\epsilon>0$ sufficiently small so that $\tau-\epsilon>\beta_{0}(wa_{-1}w^{-1})$, and then by Lemma 6.5, $\mathcal{E}_{w,\epsilon}(\psi)=\emptyset$.
+For notational convenience, we may set $C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}=1$ as it does not affect the computations in the rest of this section.
+Now let $\epsilon>0$ be a sufficiently small number and fix $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ with $\tau\leq\beta_{0}(wa_{-1}w^{-1})$.
+To compute the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$, we first need to estimate the volume of the subset $E_{w}(R)$ in ${\mathbf{F}}_{w}(\mathbb{R})$, which can be written as
+$\operatorname{Vol}(E_{w}(R))=\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))$
+where $B_{R}$ is the ball of radius $R$ centered at 0 in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ and $\mu_{\mathbf{F}_{w}}$ is the Haar measure on ${\mathbf{F}}_{w}(\mathbb{R})$.
 We need the following result about the asymptotic volume estimates of algebraic varieties.
-Theorem 6.7 ([3, Corollary 16.3]).
-[…]
+Theorem 6.7 ([4, Corollary 16.3]).
+Let $\mathcal{O}$ be a closed orbit of a group $\mathbf{H}(\mathbb{R})$ of real points of an algebraic group $\mathbf{H}$ in an $\mathbb{R}$-vector space $V$, $\mu$ an $\mathbf{H}(\mathbb{R})$-invariant measure on $\mathcal{O}$ and $\|\cdot\|$ a Euclidean norm on $V$.
+Let $B_{R}=\{v\in V:\|v\|\leq R\}$.
+Then
 $\mu(B_{R})\sim cR^{a}(\log R)^{b}\;(\textup{as }R\to\infty)$
 for some $a\in\mathbb{Q}_{\geq 0}$, $b\in\mathbb{Z}_{\geq 0}$ and $c>0$.
-We apply Theorem 6.7 to the morphism $\Psi_{w_{i}}$ where the closed orbit $\mathcal{O}=\Psi_{w_{i}}(\mathbf{F}_{w_{i}}(\mathbb{R}))$ and the measuer $\mu$ is the push-forward $\Psi_{w_{i}}^{*}(\mu_{\mathbf{F}_{w_{i}}})$ of the Haar measure on $\mathbf{F}_{w_{i}}(\mathbb{R})$, and obtain the following
+We apply Theorem 6.7 to the morphism $\Psi_{w}$ where the closed orbit $\mathcal{O}=\Psi_{w}(\mathbf{F}_{w}(\mathbb{R}))$ and the measuer $\mu$ is the push-forward $\Psi_{w}^{*}(\mu_{\mathbf{F}_{w}})$ of the Haar measure on $\mathbf{F}_{w}(\mathbb{R})$, and obtain the following
 Corollary 6.8.
-There exist constants $a_{w_{i}}\in\mathbb{Q}_{\geq 0}$, $b_{w_{i}}\in\mathbb{Z}_{\geq 0}$ and $c_{w_{i}}>0$ such that
-$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))\sim c_{w_{i}}R^{a_{w_{i}}}(\log R)^{b_{w_{i}}}\;(\textup{as }R\to\infty).$
-To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we also need an upper bound for the asymptotic number of rational elements in
-$(w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q}).$
+There exist constants $a_{w}\in\mathbb{Q}_{\geq 0}$, $b_{w}\in\mathbb{Z}_{\geq 0}$ and $c_{w}>0$ such that
+$\operatorname{Vol}(E_{w}(R))=\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))\sim c_{w}R^{a_{w}}(\log R)^{b_{w}}\;(\textup{as }R\to\infty).$
+To compute the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$, we also need an upper bound for the asymptotic estimate of the number of rational elements in
+$(w\mathbf{F}_{w}w^{-1})(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q}).$
 Consider the morphism
-$\tilde{\Psi}_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V,\quad\tilde{\Psi}_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
-Let $g\in w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in a bounded set $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$.
-By definition, $d(g)$ is the co-volume of $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}(\mathbb{R})$.
-In the $\dim V_{\beta_{0}}$-exterior product space of $V$, this implies that there exists a constant $C>0$ depending only on $\tilde{U}$, such that the length of the primitive integral vector in the line spanned by $\rho(g^{-1})\cdot e_{V_{\beta_{0}}}$ is less than $C\cdot d(g)$.
-So the number of rational points $g$ in $\tilde{U}$ whose denominators are less than $l>0$ is less than or equal to the total number of primitive integral points in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ whose lengths are less than $C\cdot l$.
-This leads us to the results about the Manin’s conjecture [2] (for relevant results about the Manin’s conjecture, one may refer to [11, 16, 18, 19, 32, 36, 38, 40] and the references therein.) In particular, there exists a constant $A_{w_{i}}>0$ such that for any $\epsilon>0$
-($\ast$) #|$\#\left|\{g\in\tilde{U}\cap w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q}):d(g)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{A_{w_{i}}+\epsilon}$
-where the implicit constant depends only on $\epsilon$, $\tilde{U}$, $\mathbf{G}$ and $\rho$. In the following, we will fix the constants $a_{w_{i}},b_{w_{i}},c_{w_{i}}$ and $A_{w_{i}}>0$.
-Lemma 6.9.
-Let $i\in\bar{I}$ such that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0.$ Then
-$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{w_{i}})\}>0.$
-Proof.
-The proof is similar to those of Lemma 4.2 and Lemma 5.5.
-Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{w_{i}})$, $\alpha(a_{1})=0$.
-Let $\Omega_{\mathbf{F}_{w_{i}}}$ and $\Omega_{\mathbf{H}_{w_{i}}}$ be small open neighborhoods of identity in $\mathbf{F}_{w_{i}}(\mathbb{R})$ and $\mathbf{H}_{w_{i}}(\mathbb{R})$ respectively, and
-$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{w_{i}}}\cdot\Omega_{\mathbf{H}_{w_{i}}}.$
-Note that $w_{i}\mathbf{H}_{w_{i}}(\mathbb{R})w_{i}^{-1}$ stabilizes every element in $V_{\beta_{0}}$.
-Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{w_{i}}}$ and $h\in\Omega_{\mathbf{H}_{w_{i}}}$, we have
-$a_{t}\cdot p\cdot w_{i}^{-1}=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}w_{i}^{-1}=fw_{i}^{-1}\cdot(w_{i}a_{t}ha_{-t}w_{i}^{-1})\cdot w_{i}a_{t}w_{i}^{-1}.$
-The element $w_{i}a_{t}w_{i}^{-1}$ is rational and
-$\delta(\rho(w_{i}a_{t}w_{i}^{-1})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(w_{i}a_{t}w_{i}^{-1})}.$
-This implies that
-$\delta(\rho(a_{t}\cdot p\cdot w_{i}^{-1})\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(w_{i}a_{-1}w_{i}^{-1})t}$
-for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{w_{i}}}$ and $w_{i}$.
-Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$.
-Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot w_{i}^{-1}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})/\Gamma$.
-Set $\Gamma_{w_{i}}=\Gamma\cap w_{i}^{-1}\Gamma w_{i}$.
-Then $\Gamma_{w_{i}}$ is commensurable with $\Gamma$ and $w_{i}^{-1}\Gamma w_{i}$, and we can conclude that for any $p\Gamma_{w_{i}}\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot\Gamma_{w_{i}}$, the orbit $a_{t}\cdot p\Gamma_{w_{i}}$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma_{w_{i}}$.
-Similiar to Lemma 4.2 and Lemma 5.5, this contradicts the ergodic property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\Gamma_{w_{i}}\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$, where $\tilde{\mathbf{G}}$ is the product of some $\mathbb{Q}$-simple factors $\mathbf{G}_{i}$ $(1\leq i\leq s)$ of $\mathbf{G}$ such that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})$ $(1\leq i\leq s)$.
-This completes the proof of the lemma.
-∎
-Now we compute an upper bound for the Hausdorff dimension of the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$.
-In the computation, we project the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by the natural projection map
+$\Psi_{w}:\mathbf{F}_{w}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w}(x)=\rho_{\beta_{0}}(wxw^{-1})\cdot e_{V_{\beta_{0}}}.$
+Let $\Omega_{\beta_{0}}$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ which contains a fundamental domain of $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})\cap\Gamma)$.
+Let $g\in w\mathbf{F}_{w}w^{-1}(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in $\Omega_{\beta_{0}}$.
+By definition, $0pt(g)$ is the co-volume of $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}(\mathbb{R})$.
+In the $\dim V_{\beta_{0}}$-exterior product space of $V$, this implies that there exists a constant $C_{\beta_{0}}>0$ depending only on $\Omega_{\beta_{0}}$, such that the length of the primitive integral vector in the line spanned by $\rho(g^{-1})\cdot e_{V_{\beta_{0}}}$ is less than $C_{\beta_{0}}\cdot 0pt(g)$.
+So the number of rational elements $g$ in $\Omega_{\beta_{0}}\cap(w\mathbf{F}_{w}w^{-1})(\mathbb{Q})$ whose heights are less than $l>0$ is bounded above by the total number of primitive integral points in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ whose lengths are less than $C_{\beta_{0}}\cdot l$.
+This leads to the results about Manin’s conjecture [3], i.e.
+the asymptotic estimate of the number of rational points in an algebraic variety (See e.g.
+[12, 24, 33, 34, 63, 68, 72, 76, 73] and the references therein.) For the general situation, we define
+$A_{w}=A_{w}(\Omega_{\beta_{0}}):=\limsup_{l\to\infty}\frac{\log|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq l\}|}{\log l}$
+and from the discussion above, we know that $A_{w}$ is a finite number which is bounded above by the growth rate of the number of the primitive integral points in $\bigwedge^{\dim V_{\beta_{0}}}V$.
+We also have that for any $\epsilon>0$
+$|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq l\}|\lesssim_{\epsilon,\Omega_{\beta_{0}}}l^{A_{w}+\epsilon}$
+where the implicit constant depends only on $\epsilon,\Omega_{\beta_{0}},\mathbf{G}$ and $\rho$.
+In the following, we will fix the constants $a_{w},b_{w},c_{w}$ in Corollary 6.8 and $A_{w}$ for the bounded open subset $\Omega_{\beta_{0}}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$.
+Now we compute an upper bound for the Hausdorff dimension of the subset $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$ $(\tau\leq\beta_{0}(wa_{-1}w^{-1}))$.
+In the computation, we project the subset $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by the natural projection map
 $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}:R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\to R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
-and compute the Hausdorff dimension of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$.
-Since $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}$ is a local isometry, by the countable stability of Hausdorff dimension, we then obtain the upper bounds for $\dim_{H}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}^{-1}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))))$ and $\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$.
+and compute the Hausdorff dimension of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$.
+Since $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}$ is a local isometry, by the countable stability of Hausdorff dimension, we then obtain upper bounds for $\dim_{H}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}^{-1}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))))$ and $\dim_{H}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$.
 Let $\tilde{\Omega}$ be a bounded fundamental domain of $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-We choose a bounded subset $\Omega_{\mathbf{H}_{w_{i}}}$ in $\mathbf{H}_{w_{i}}(\mathbb{R})$ and a bounded subset $\Omega_{\mathbf{F}_{w_{i}}}$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
-$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w_{i}}}\cdot\Omega_{\mathbf{F}_{w_{i}}}.$
-Now suppose that $q\in\mathbf{G}(\mathbb{R})$ is a rational element with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-Let $\tilde{q}\in\mathbf{G}(\mathbb{R})$ such that $w_{i}^{-1}\tilde{q}$ is a representative of $w_{i}^{-1}q$ in $\tilde{\Omega}$ with
-$w_{i}^{-1}\tilde{q}=w_{i}^{-1}q\gamma$
+We choose a bounded subset $\Omega_{\mathbf{H}_{w}}$ in $\mathbf{H}_{w}(\mathbb{R})$ and a bounded subset $\Omega_{\mathbf{F}_{w}}$ in $\mathbf{F}_{w}(\mathbb{R})$ such that
+$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w}}\cdot\Omega_{\mathbf{F}_{w}}.$
+Without loss of generality, we may assume that
+$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w}}\cdot\Omega_{\mathbf{F}_{w}}\subset\Omega_{\beta_{0}}.$
+Now suppose that $q\in\mathbf{G}(\mathbb{R})$ is a rational element with $w^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+Let $\tilde{q}\in\mathbf{G}(\mathbb{R})$ such that $w^{-1}\tilde{q}$ is a representative of $w^{-1}q$ in $\tilde{\Omega}$ with
+$w^{-1}\tilde{q}=w^{-1}q\gamma$
 for some $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-Then one can check that $\tilde{q}$ is still a rational element in $\mathbf{G}(\mathbb{R})$ and $d(\tilde{q})=d(q)$ since $\Gamma$ preserves the lattice $\mathbb{Z}^{d}$.
-So by Corollary 2.3, we still have
-$w_{i}^{-1}\tilde{q}\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
-Since $w_{i}^{-1}\tilde{q}$ is an element in the bounded fundamental domain $\tilde{\Omega}$, it implies that
-$w_{i}^{-1}\tilde{q}\in\Omega_{\mathbf{H}_{w_{i}}}\cdot(\mathbf{F}_{w_{i}}(\mathbb{Q})\cap\Omega_{\mathbf{F}_{w_{i}}}).$
+Then one can check that $\tilde{q}$ is still a rational element in $\mathbf{G}(\mathbb{R})$ and $0pt(\tilde{q})=0pt(q)$ since $\Gamma$ preserves the lattice $\mathbb{Z}^{d}$.
+So by Corollary 2.3, we have
+$w^{-1}\tilde{q}\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{F}_{w}(\mathbb{Q}).$
+Since $w^{-1}\tilde{q}$ is an element in the bounded fundamental domain $\tilde{\Omega}$, it implies that
+$w^{-1}\tilde{q}\in\Omega_{\mathbf{H}_{w}}\cdot(\mathbf{F}_{w}(\mathbb{Q})\cap\Omega_{\mathbf{F}_{w}}).$
 We then obtain the following
 Proposition 6.9.
-The set $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ consists of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
-$g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
-$d(q_{k})\leq C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
-and $w_{i}^{-1}q_{k}\in\tilde{\Omega}$.
+The set $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$ consists of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
+$g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in\left(a_{-t_{k}}E_{w}(e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k}))a_{t_{k}}\right)\cdot w^{-1}q_{k}(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$0pt(q_{k})\leq e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
+and $w^{-1}q_{k}\in\tilde{\Omega}$.
 Lemma 6.10.
 Let $\delta>0$ and $B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ be the small open ball of radius $\delta$ centered at identity in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
-Then there exists a constant $C_{\delta}>0$ depending only on $\delta$ such that for any $R>0$
-$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}(R)\subset E_{w_{i}}(C_{\delta}\cdot R)\cdot\mathbf{H}_{w_{i}}(\mathbb{R}).$
+Then there exists a constant $C_{\delta}>0$ depending only on $\delta$ and $\rho$ such that for any $R>0$
+$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w}(R)\subset E_{w}(C_{\delta}\cdot R)\cdot\mathbf{H}_{w}(\mathbb{R}).$
 Proof.
-Let $x\in B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ and $y\in E_{w_{i}}(R)$. Then $x\cdot y\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We write
+Let $x\in B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ and $y\in E_{w}(R)$. Then $x\cdot y\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and we write
 $x\cdot y=u\cdot v$
-where $u\in\mathbf{F}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{H}_{w_{i}}(\mathbb{R})$. Then we have
-$\|\Psi_{w_{i}}(u)\|=$ $\|\rho_{\beta_{0}}(w_{i}uw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\rho(w_{i}uvw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|$
-$=$ $\|\rho(w_{i}xyw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq C_{\delta}\cdot R$
-for some constant $C_{\delta}>0$ depending only on $\delta>0$. This completes the proof of the lemma. ∎
-Note that by definition, for any rational element $g\in\mathbf{G}(\mathbb{R})$, every element $x$ in the $\Gamma$-coset $g\Gamma$ of $g$ is also rational, and $d(x)=d(g)$.
+where $u\in\mathbf{F}_{w}(\mathbb{R})$ and $v\in\mathbf{H}_{w}(\mathbb{R})$. Then we have
+$\|\Psi_{w}(u)\|=$ $\|\rho_{\beta_{0}}(wuw^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\rho_{\beta_{0}}(wuvw^{-1})\cdot e_{V_{\beta_{0}}}\|$
+$=$ $\|\rho_{\beta_{0}}(wxyw^{-1})\cdot e_{V_{\beta_{0}}}\|\leq C_{\delta}\cdot R$
+for some constant $C_{\delta}>0$ depending only on $\delta>0$ and $\rho$. This completes the proof of the lemma. ∎
+Note that by definition, for any rational element $g\in\mathbf{G}(\mathbb{R})$, every element $x$ in the $\Gamma$-coset $g\Gamma$ of $g$ is also rational, and $0pt(x)=0pt(g)$.
 We will use this fact in the following lemma.
 Lemma 6.11.
-Let $i\in\bar{I}$ and $L>0$.
-Then the set of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$ is equal to
-$\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}.$
+Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and $L>0$.
+Then the set of elements $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ with $w\cdot g$ rational and $0pt(w\cdot g)=L$ is equal to
+$\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w}},y\in\Omega_{\mathbf{F}_{w}}\cap\mathbf{F}_{w}(\mathbb{Q}),0pt(w\cdot y)=L\}.$
 Proof.
-Let $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in R_{u}(\overline{\mathbf{P}}_{0})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$.
+Let $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in R_{u}(\overline{\mathbf{P}}_{0})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0}))$ with $w\cdot g$ rational and $0pt(w\cdot g)=L$.
 We write
 $g=(u\cdot v)\cdot\gamma$
-for some $u\in\Omega_{\mathbf{H}_{w_{i}}}$, $v\in\Omega_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})$. Then we have
-$w_{i}\cdot g=(w_{i}\cdot u\cdot w_{i}^{-1})\cdot w_{i}\cdot v\cdot\gamma$
-and $w_{i}\cdot v$ is rational with
-$d(w_{i}\cdot g)=d(w_{i}\cdot v)=L.$
-By Corollary 2.3, we know that $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. This completes the proof of the lemma. ∎
-Remark 6.13.
+for some $u\in\Omega_{\mathbf{H}_{w}}$, $v\in\Omega_{\mathbf{F}_{w}}$ and $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})$. Then we have
+$w\cdot g=(w\cdot u\cdot w^{-1})\cdot w\cdot v\cdot\gamma$
+and $w\cdot v$ is rational with
+$0pt(w\cdot g)=0pt(w\cdot v)=L.$
+By Corollary 2.3, we know that $v\in\mathbf{F}_{w}(\mathbb{Q})$. This completes the proof of the lemma. ∎
+Definition 6.12.
 For any $L>0$, we define
-$\mathcal{S}(L):=\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}$
-$\mathcal{F}(L):=\{y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}):d(w_{i}\cdot y)=L\}.$
-Now we construct open covers of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ with arbitrarily small diameters.
-Here we use the right-invariant metric $d_{R_{u}(\overline{\mathbf{P}}_{0})}$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which induces a metric $d_{R_{u}(\overline{\mathbf{P}}_{0})/\Gamma}$ on the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$.
+$\mathcal{S}(L):=\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w}},y\in\Omega_{\mathbf{F}_{w}}\cap\mathbf{F}_{w}(\mathbb{Q}),0pt(w\cdot y)=L\}$
+$\mathcal{F}(L):=\{y\in\Omega_{\mathbf{F}_{w}}\cap\mathbf{F}_{w}(\mathbb{Q}):0pt(w\cdot y)=L\}.$
+Now we construct open covers of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$ with arbitrarily small diameters.
+Here we use the right-invariant metric $d_{R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})}$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which induces a metric on the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$.
 Note that by Corollary 6.8, for any $\epsilon>0$ and $R>0$, we have
-$\operatorname{Vol}(E_{w_{i}}(R))\lesssim_{\epsilon}R^{a_{w_{i}}+\epsilon}$
-where the implicit constant depends only on $\epsilon$ and $i\in\bar{I}$. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
+$\operatorname{Vol}(E_{w}(R))\lesssim_{\epsilon}R^{a_{w}+\epsilon}$
+where the implicit constant depends only on $\epsilon$ and $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. Let $\nu_{0}$ be any $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
 $\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
-By Lemma 6.9, we have $\nu_{0}(a_{1})>0$.
-Fix $\delta>0$. Let $L>0$ and $t\in\mathbb{N}$ such that
-$L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
-Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ such that $d(q)=L$ and $w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ The following subset
-$\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+Since the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, we have $\nu_{0}(a_{1})>0$.
+Fix a small number $\delta>0$. Let $L>0$ and $t\in\mathbb{N}$ such that
+$L\leq e^{\lambda_{w}(\epsilon)\cdot t}.$
+Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ such that $0pt(q)=L$ and $w^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ The following subset
+$\left(a_{-t}\cdot E_{w}\left(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q)\right)\cdot a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
 can be covered by disjoint boxes of diameter at most $\delta\cdot\exp(\nu_{0}(a_{-t}))$, and by Lemma 6.10 and Lemma 6.11, we know that these boxes are contained in
-$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
-$\subset$ $\left(a_{-t}\cdot B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
-$\subset$ $\left(a_{-t}\cdot E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)\cdot a_{t}\right)\cdot\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L).$
-Since every point $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)$ satisfies the property that $w_{i}\cdot g$ is rational and $d(w_{i}\cdot g)=L$, it follows again from Lemma 6.12 that
-$\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)=\mathcal{S}(L).$
+$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))\left(a_{-t}\cdot E_{w}\left(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q)\right)\cdot a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$\subset$ $\left(a_{-t}\cdot B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w}\left(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q)\right)\cdot a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$\subset$ $\left(a_{-t}\cdot E_{w}\left(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L\right)\cdot a_{t}\right)\cdot\mathbf{H}_{w}(\mathbb{R})\cdot\mathcal{S}(L).$
+Since every point $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $\mathbf{H}_{w}(\mathbb{R})\cdot\mathcal{S}(L)$ satisfies the property that $w\cdot g$ is rational and $0pt(w\cdot g)=L$, it follows again from Lemma 6.11 that
+$\mathbf{H}_{w}(\mathbb{R})\cdot\mathcal{S}(L)=\mathcal{S}(L).$
 We conclude that the subset
-$\left(a_{-t}E_{w_{i}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q))a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$\left(a_{-t}E_{w}(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q))a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
 can be covered by disjoint boxes with diameter at most $\delta\cdot\exp(\nu_{0}(a_{-t}))$ and these boxes are contained in
-$a_{-t}E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)a_{t}\cdot\mathcal{S}(L).$
+$a_{-t}E_{w}\left(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L\right)a_{t}\cdot\mathcal{S}(L).$
 We collect all these boxes constructed above in a family which is denoted by $\mathcal{P}_{t,L}(q)$, and define
-$\mathcal{P}_{t,L}=\bigcup_{d(q)=L\textup{ and }w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})}\mathcal{P}_{t,L}(q).$
+$\mathcal{P}_{t,L}=\bigcup_{\begin{subarray}{c}0pt(q)=L\\ w^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\end{subarray}}\mathcal{P}_{t,L}(q).$
 Now since every box in $\mathcal{P}_{t,L}$ is contained in the same bounded set
-$a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t}\cdot\mathcal{S}(L)$
+$a_{-t}E_{w}(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L)a_{t}\cdot\mathcal{S}(L)$
 we can choose a maximal finite sub-collection $\mathcal{Q}_{t,L}$ of disjoint boxes in $\mathcal{P}_{t,L}$. Then by the maximality of $\mathcal{Q}_{t,L}$ we have
-$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\mathcal{Q}_{t,L}\supset\bigcup_{S\in\mathcal{P}_{t,L}}S.$
-Meanwhile, the number of boxes in $\mathcal{Q}_{t,L}$ is at most
-$\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}.$
+$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))^{2}\cdot\bigcup_{S\in\mathcal{Q}_{t,L}}S\supset\bigcup_{S\in\mathcal{P}_{t,L}}S.$
+Moreover, the number of boxes in $\mathcal{Q}_{t,L}$ is at most (up to a constant)
+$\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}((a_{-t}E_{w}(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L)a_{t})\cdot\mathcal{S}(L))}{\delta^{\dim R_{u}(\overline{\mathbf{P}}_{0})}\cdot\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}.$
 Now we define $\mathcal{G}_{t}$ to be the collection of boxes in
@@
 where
-$\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
-Note that $\operatorname{diam}\mathcal{G}_{t}\sim e^{\nu_{0}(a_{-t})}$ and $\nu_{0}(a_{-t})\neq 0$ by Lemma 6.9. Define
+$\theta_{w}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot t}.$
+Note that $\operatorname{diam}\mathcal{G}_{t}\asymp e^{\nu_{0}(a_{-t})}$ and $\nu_{0}(a_{-1})\neq 0$. Define
 $\mathcal{G}^{k}=\bigcup_{t\geq k}\mathcal{G}_{t}.$
-Then by definition and Proposition 6.10, the subset $\mathcal{G}^{k}$ is a cover of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ for any $k\in\mathbb{N}$ and $\operatorname{diam}\mathcal{G}^{k}\sim e^{\nu_{0}(a_{-k})}$.
+Then by definition and Proposition 6.9, the subset $\mathcal{G}^{k}$ is a cover of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$ for any $k\in\mathbb{N}$ and $\operatorname{diam}\mathcal{G}^{k}\asymp e^{\nu_{0}(a_{-k})}$.
 Now according to the construction of the open covers $\mathcal{G}^{k}$ $(k\in\mathbb{N})$ above, we consider the following series with respect to the parameter $s$
 $\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}\leq\sum_{t\in\mathbb{N}}\sum_{B\in\mathcal{G}_{t}}\operatorname{diam}(B)^{s}$
-$\lesssim$ $\sum_{t\in\mathbb{N}}\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w_{i}}(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}\cdot e^{s\nu_{0}(a_{-t})}.$
+$\lesssim$ $\sum_{t\in\mathbb{N}}\frac{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(a_{-t}E_{w}(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L)a_{t})\cdot\mathcal{S}(L))}{\exp(\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t}))}\cdot e^{s\nu_{0}(a_{-t})}.$
 By Corollary 6.8, for any $\epsilon>0$, we have
 $\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
-$\lesssim$ $\sum_{t\in\mathbb{N}}\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L)^{a_{w_{i}}+\epsilon}\cdot|\mathcal{F}(L)|$
-$e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}.$
-Fix $t\in\mathbb{N}$. Note that $\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$. Choose $t_{0}\in\mathbb{Z}$ such that
-$C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot t_{0}})^{\dim V_{\beta_{0}}}\leq\theta_{i}.$
-Then by equation ($\ast$6) before Lemma 6.9, we have
-$\sum_{\theta_{i}\leq L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
-$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\sum_{C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}}\leq L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|/L^{a_{w_{i}}+\epsilon}$
-$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}\sum_{L\leq C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}}}|\mathcal{F}(L)|$
-$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot(l+1)})^{\dim V_{\beta_{0}}})^{A_{w_{i}}+\epsilon}$
-$\sim$ $\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
-Combining all these equations, we obtain that
-$\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
-$\lesssim$ $\sum_{t\in\mathbb{N}}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)t})^{\dim V_{\beta_{0}}})^{a_{w_{i}}+\epsilon}\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}$
-$\cdot\sum_{t_{0}\leq l\leq t-1}(C_{0}(e^{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)\cdot l})^{\dim V_{\beta_{0}}})^{A_{w_{i}}-a_{w_{i}}}.$
-By computing the series above in the cases $A_{w_{i}}<a_{w_{i}}$, $A_{w_{i}}=a_{w_{i}}$ and $A_{w_{i}}>a_{w_{i}}$, one can conclude that the series above converges if
-$s>\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
-This implies that for any $\epsilon>0$
-$\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))\leq\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w_{i}},a_{w_{i}}\}+\epsilon)\cdot\dim V_{\beta_{0}}.$
+$\lesssim$ $\sum_{t\in\mathbb{N}}\sum_{\theta_{w}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot t}}(e^{\lambda_{w}(\epsilon)\cdot t}/L)^{a_{w}+\epsilon}\cdot|\mathcal{F}(L)|\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}.$
+Fix $t\in\mathbb{N}$.
+Note that $\tau\leq\beta_{0}(wa_{-1}w^{-1})$.
+Choose $t_{0}\in\mathbb{Z}$ such that $e^{\lambda_{w}(\epsilon)\cdot t_{0}}\leq\theta_{w}.$ Then by the definition of $A_{w}$, we have
+$\sum_{\theta_{w}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot t}}|\mathcal{F}(L)|/L^{a_{w}+\epsilon}$
+$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\sum_{e^{\lambda_{w}(\epsilon)\cdot l}\leq L\leq e^{\lambda_{w}(\epsilon)\cdot(l+1)}}|\mathcal{F}(L)|/L^{a_{w}+\epsilon}$
+$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(e^{\lambda_{w}(\epsilon)\cdot l})^{a_{w}+\epsilon}}\sum_{L\leq e^{\lambda_{w}(\epsilon)\cdot(l+1)}}|\mathcal{F}(L)|$
+$\lesssim$ $\sum_{t_{0}\leq l\leq t-1}\frac{1}{(e^{\lambda_{w}(\epsilon)\cdot l})^{a_{w}+\epsilon}}(e^{\lambda_{w}(\epsilon)\cdot(l+1)})^{A_{w}+\epsilon}$
+$\asymp$ $\sum_{t_{0}\leq l\leq t-1}(e^{\lambda_{w}(\epsilon)\cdot l})^{A_{w}-a_{w}}.$
+Combining all the inequalities above, we obtain
+(4) $\sum_{B\in\mathcal{G}^{k}}\operatorname{diam}(B)^{s}$
+$\lesssim$ $\sum_{t\in\mathbb{N}}(e^{\lambda_{w}(\epsilon)\cdot t})^{a_{w}+\epsilon}\cdot e^{\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\alpha(a_{-t})}e^{-\dim R_{u}(\overline{\mathbf{P}}_{0})\cdot\nu_{0}(a_{-t})}\cdot e^{s\nu_{0}(a_{-t})}\cdot\sum_{t_{0}\leq l\leq t-1}(e^{\lambda_{w}(\epsilon)\cdot l})^{A_{w}-a_{w}}.$
+Computing the series (4) in the cases $A_{w}<a_{w}$, $A_{w}=a_{w}$ and $A_{w}>a_{w}$, one can conclude that the series (4) converges if
+$s>\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{\lambda_{w}(\epsilon)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w},a_{w}\}+\epsilon).$
+This implies that for sufficiently small $\epsilon>0$
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$
+$\leq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{\lambda_{w}(\epsilon)}{\nu_{0}(a_{1})}\cdot(\max\{A_{w},a_{w}\}+\epsilon).$
 By taking $\epsilon\to 0$ and Proposition 6.6, we conclude that
-$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
-$\leq\max_{i\in\bar{I}}\left\{\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}$
+$\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$\leq\max_{\begin{subarray}{c}w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}\\ \beta_{0}(wa_{-1}w^{-1})\\ \geq\tau\end{subarray}}\left\{\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}\right\}.$
+Proof of Theorem 1.5.
+By the Bruhat decomposition of $\mathbf{G}$, we have
+$\mathbf{G}(\mathbb{R})=\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot w\cdot\mathbf{P}_{0}(\mathbb{R})=\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot w\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R}).$
+For any $w\in{{}_{\mathbb{Q}}\mathcal{W}}$, define
+$\mathbf{U}_{w}=(w^{-1}\cdot R_{u}(\overline{\mathbf{P}}_{0})\cdot w)\cap R_{u}(\mathbf{P}_{0})\textup{ and }\mathbf{Q}_{w}=(w^{-1}\cdot\mathbf{P}_{0}\cdot w)\cap R_{u}(\mathbf{P}_{0}).$
+Then $R_{u}(\mathbf{P}_{0})=\mathbf{Q}_{w}\cdot\mathbf{U}_{w}$ and we have
+$\mathbf{G}(\mathbb{R})=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot w\cdot(\mathbf{Q}_{w}(\mathbb{R})\cdot\mathbf{U}_{w}(\mathbb{R}))$
+$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{R})\cdot(w\cdot\mathbf{U}_{w}(\mathbb{R})\cdot w^{-1})\cdot w$
+$=$ $\mathbf{P}_{0}(\mathbb{R})\cdot R_{u}(\overline{\mathbf{P}}_{0})\cdot{{}_{\mathbb{Q}}\mathcal{W}}.$
+By definition, if an element $g\in\mathbf{G}(\mathbb{R})$ belongs to $S_{\rho}(\psi)$, then for any $h\in\mathbf{P}_{0}(\mathbb{R})$ and any $w\in\mathbf{G}(\mathbb{Q})$, $h\cdot g\cdot w$ also belongs to $S_{\rho}(\psi)$.
+Therefore, we have
+$S_{\rho}(\psi)^{c}=\mathbf{P}_{0}(\mathbb{R})\cdot(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\cdot{{}_{\mathbb{Q}}\mathcal{W}}$
 and
-$\dim_{H}S_{\tau}^{c}\leq\max_{i\in\bar{I}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}\right\}.$
-This completes the proof of Theorem 1.4.
+$\dim_{H}S_{\rho}(\psi)^{c}=\dim\mathbf{P}_{0}+\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
+$\leq$ $\max_{\begin{subarray}{c}w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}\\ \beta_{0}(wa_{-1}w^{-1})\geq\tau\end{subarray}}\left\{\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}\right\}.$
+This completes the proof of Theorem 1.5. ∎
 7. Proofs of Theorems 1.61.7 and 1.8
-In this section, we prove Theorems 1.61.7 and 1.8.
+In this section, we derive Theorems 1.6 from Theorems 1.4 and 1.5, and then prove Theorems 1.7 and 1.8.
+Let $\mathbf{G}$ be a connected $\mathbb{Q}$-simple group.
+We assume that the highest weight $\beta_{0}$ in $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ is a multiple of $\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}$ and write
+$\beta_{0}=\kappa\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}$
+for some constant $\kappa<0$,
+For any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, define
+$\mathbf{F}^{w}=R_{u}(\mathbf{P}_{0})\cap w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad\mathbf{H}^{w}=R_{u}(\mathbf{P}_{0})\cap w^{-1}\mathbf{P}_{\beta_{0}}w.$
+Since $R_{u}(\overline{\mathbf{P}}_{0})$ and $R_{u}(\mathbf{P}_{0})$ contains all the $\mathbb{Q}$-roots of $\mathbf{G}$ (with multiplicities), by [8, Proposition 21.9], we have
+$\mathbf{F}^{w}\cdot\mathbf{F}_{w}=w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad(w\mathbf{F}^{w}w^{-1})\cdot(w\mathbf{F}_{w}w^{-1})=R_{u}(\overline{\mathbf{P}}_{\beta_{0}}).$
+So by [55, Proposition 5.26], the Haar measure $\mu_{R_{u}(\overline{\mathbf{P}}_{\beta_{0}})}$ on $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ can be written as a product of the Haar measure $\mu_{w\mathbf{F}^{w}w^{-1}}$ on $w\mathbf{F}^{w}(\mathbb{R})w^{-1}$ and the Haar measure $\mu_{w\mathbf{F}_{w}w^{-1}}$ on $w\mathbf{F}_{w}(\mathbb{R})w^{-1}$.
+Note that in the representation $\rho_{\beta_{0}}:\mathbf{G}\to\operatorname{\mathbf{GL}}(\bigwedge^{\dim V_{\beta_{0}}}V)$ (the $\dim V_{\beta_{0}}$-exterior product of $\rho$), the linear space spanned by
+$e_{V_{\beta_{0}}}:=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}\in\bigwedge^{\dim V_{\beta_{0}}}V$
+is the unique weight space of highest weight $\dim V_{\beta_{0}}\cdot\beta_{0}$ relative to $\mathbf{T}$.
+Let $W_{\beta_{0}}\subset\bigwedge^{\dim V_{\beta_{0}}}V$ be the irreducible sub-representation of $\mathbf{G}$ in $\rho_{\beta_{0}}$ such that $W_{\beta_{0}}$ contains $\mathbb{C}\cdot e_{V_{\beta_{0}}}$.
+Then $\Psi_{w}$ and $a_{w},b_{w},c_{w},A_{w}$ ($w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$), which are defined in §6, can be considered as morphisms and quantities defined in the sub-representation $W_{\beta_{0}}$.
+Lemma 7.1.
+For any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, we have $a_{w}\leq a_{e}$ and $A_{w}\leq A_{e}$.
+Proof.
+Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. The inequality $A_{w}\leq A_{e}$ follows from the fact that $w\mathbf{F}_{w}w^{-1}\subseteq R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\mathbf{F}_{e}$.
+Now let $B_{w\mathbf{F}^{w}w^{-1}}(1)$ be the open ball of radius one around the identity in $w\mathbf{F}^{w}(\mathbb{R})w^{-1}$. For any $R>0$, consider the subset
+$B_{w\mathbf{F}^{w}w^{-1}}(1)\cdot(w\cdot\Psi_{w}^{-1}(B_{R})\cdot w^{-1})$
+in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})=\mathbf{F}_{e}(\mathbb{R})$. We know that there exists a constant $C>0$ such that for any $f\in B_{w\mathbf{F}^{w}w^{-1}}(1)$
+$\|\rho_{\beta_{0}}(f)\cdot v\|\leq C\|v\|\quad\left(\forall v\in\bigwedge^{\dim\beta_{0}}V\right).$
+By the definitions of the morphisms $\Psi_{w}$ and $\Psi_{e}$, it implies that for any $R>0$
+$B_{w\mathbf{F}^{w}w^{-1}}(1)\cdot(w\cdot\Psi_{w}^{-1}(B_{R})\cdot w^{-1})\subset\Psi_{e}^{-1}(B_{C\cdot R})$
+and we have
+$\mu_{w\mathbf{F}^{w}w^{-1}}(B_{w\mathbf{F}^{w}w^{-1}}(1))\cdot\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))$
+$=$ $\mu_{w\mathbf{F}^{w}w^{-1}}(B_{w\mathbf{F}^{w}w^{-1}}(1))\cdot\mu_{w\mathbf{F}_{w}w^{-1}}(w\cdot\Psi_{w}^{-1}(B_{R})\cdot w^{-1})$
+$\leq$ $\mu_{\mathbf{F}_{e}}(\Psi_{e}^{-1}(B_{C\cdot R})).$
+Note that for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$
+$\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))\sim c_{w}R^{a_{w}}(\log R)^{b_{w}}\;(\textup{as }R\to\infty).$
+So $a_{w}\leq a_{e}$. This completes the proof of the lemma. ∎
+By the irreducible representation $\rho_{\beta_{0}}:\mathbf{G}\to\operatorname{\mathbf{GL}}(W_{\beta_{0}})$, the flag variety $\mathbf{G}/\mathbf{P}_{\beta_{0}}$ can be considered as a subvariety in $\mathbf{P}(W_{\beta_{0}})$ via
+$g\mathbf{P}_{\beta_{0}}\mapsto\rho_{\beta_{0}}(g)(\mathbb{C}\cdot e_{V_{\beta_{0}}}).$
+For any rational point $x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q})\subset\mathbf{P}(W_{\beta_{0}})(\mathbb{Q})$, define the height of $x$ by
+$H_{\beta_{0}}(x):=\|v\|$
+where $v$ is the unique primitive integral vector (up to sign) in the rational line representing $x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q})$ in $W_{\beta_{0}}$.
+Then we may count the number of rational points in $(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q})$ with the height function $\|\cdot\|_{\beta_{0}}$.
+By [63, Theorem 4], there exist $a_{0},b_{0},c_{0}>0$ such that
+$\left|\{x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q}):H_{\beta_{0}}(x)\leq R\}\right|\sim c_{0}\cdot R^{a_{0}}(\log R)^{b_{0}}.$
+Moreover, we may compute that $a_{0}=1/(|\kappa|\cdot\dim V_{\beta_{0}})$.
+Lemma 7.2.
+For $e\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, we have
+$a_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}),\quad A_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}).$
+Proof.
+For any $R>0$, let $B_{R}$ denote the ball of radius $R>0$ around the origin in $W_{\beta_{0}}$, and $\chi_{R}$ the characteristic function of $B_{R}$.
+Let $\Omega_{e}$ be a bounded fundamental domain of $\mathbf{F}_{e}(\mathbb{R})/\mathbf{F}_{e}(\mathbb{Z})$ in $\mathbf{F}_{e}(\mathbb{R})$.
+Then $\Omega_{e}$ is relatively compact.
+This implies that there exists $C>0$ such that for any $g\in\Omega_{e}$
+$\|\rho_{\beta_{0}}(g)\cdot v\|\leq C\|v\|\quad(\forall v\in W_{\beta_{0}}).$
+So we have
+$\mu_{\mathbf{F}_{e}}(\Psi_{e}^{-1}(B_{R}))=$ $\int_{\mathbf{F}_{e}(\mathbb{R})}\chi_{R}(u\cdot e_{V_{\beta_{0}}})d\mu_{\mathbf{F}_{e}}(u)$
+$=$ $\sum_{\gamma\in\mathbf{F}_{e}(\mathbb{Z})}\int_{\Omega_{e}}\chi_{R}(u\cdot\gamma\cdot e_{V_{\beta_{0}}})d\mu_{\mathbf{F}_{e}}(u)$
+$\leq$ $\sum_{\gamma\in\mathbf{F}_{e}(\mathbb{Z})}\int_{\Omega_{e}}\chi_{C\cdot R}(\gamma\cdot e_{V_{\beta_{0}}})d\mu_{\mathbf{F}_{e}}(u)$
+$=$ $\mu_{\mathbf{F}_{e}}(\Omega_{e})\cdot\sum_{\gamma\in\mathbf{F}_{e}(\mathbb{Z})}\chi_{C\cdot R}(\gamma\cdot e_{V_{\beta_{0}}})$
+$=$ $\mu_{\mathbf{F}_{e}}(\Omega_{e})\cdot|\{\gamma\in\mathbf{F}_{e}(\mathbb{Z}):\|\gamma\cdot e_{V_{\beta_{0}}}\|\leq C\cdot R\}|$
+$\leq$ $\mu_{\mathbf{F}_{e}}(\Omega_{e})\cdot\left|\{x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q}):H_{\beta_{0}}(x)\leq C\cdot R\}\right|.$
+Here we use the injectivity of the morphism $\Psi_{e}$. Note that for $e\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ we have
+$\mu_{\mathbf{F}_{e}}(\Psi_{e}^{-1}(B_{R}))\sim c_{e}R^{a_{e}}(\log R)^{b_{e}}\;(\textup{as }R\to\infty).$
+It implies that
+$a_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}).$
+Now by the discussion in §6, there exists a constant $C_{\beta_{0}}>0$ depending on $\Omega_{\beta_{0}}$ such that
+$|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq l\}|\leq\left|\{x\in(\mathbf{G}/\mathbf{P}_{\beta_{0}})(\mathbb{Q}):H_{\beta_{0}}(x)\leq C_{\beta_{0}}\cdot l\}\right|.$
+By taking the limit as $l\to\infty$, we obtain
+$A_{e}\leq 1/(|\kappa|\cdot\dim V_{\beta_{0}}).$
+This completes the proof of the lemma. ∎
 Proof of Theorem 1.6.
+Let $\kappa<0$ such that
+$\beta_{0}=\kappa\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}.$
+By Lemma 7.2, one can compute that
+$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(a_{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{e},A_{e}\}\cdot\dim V_{\beta_{0}}$
+$\leq$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
+Now to derive Theorem 1.6 from Theorems 1.4 and 1.5, it suffices to prove that for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and $\tau(\psi)\leq\beta_{0}(wa_{-1}w^{-1})$
+$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{w},A_{w}\}\cdot\dim V_{\beta_{0}}$
+$\leq$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
+Note that
+$\beta_{0}=$ $\kappa\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha$
+$=$ $\kappa\cdot\left(\sum_{\alpha\in\Phi(w\mathbf{F}^{w}w^{-1})}\alpha+\sum_{\alpha\in\Phi(w\mathbf{F}_{w}w^{-1})}\alpha\right)$
+and
+$\beta_{0}(wa_{-1}w^{-1})=\kappa\cdot\left(\sum_{\alpha\in\Phi(\mathbf{F}^{w})}\alpha(a_{-1})+\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\alpha(a_{-1})\right).$
+So by Lemma 7.1 and Lemma 7.2, we have
+$-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{w},A_{w}\}\cdot\dim V_{\beta_{0}}$
+$\leq$ $-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot a_{0}\cdot\dim V_{\beta_{0}}$
+$=$ $-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{\beta_{0}(wa_{-1}w^{-1})}{\nu_{0}(a_{1})|\kappa|}-\frac{\tau(\psi)}{\nu_{0}(a_{1})|\kappa|}$
+$=$ $\sum_{\alpha\in\Phi(\mathbf{F}^{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}-\frac{\tau(\psi)}{\nu_{0}(a_{1})|\kappa|}$
+$\leq$ $-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
+Here we use the fact that all $\mathbb{Q}$-roots in $\Phi(\mathbf{F}^{w})$ are positive and $\alpha(a_{1})\leq 0$ for any $\alpha\in\Phi(\mathbf{F}^{w})$. This completes the proof of the theorem. ∎
+Proofs of Theorems 1.7 and 1.8.
 Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Without loss of generality, we may write
@@
 $\alpha_{i,j}(a)=t_{i}-t_{j}\quad(1\leq i\neq j\leq n)$
-where $a=\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}})$. Let $\{e_{1},e_{2},...,e_{n}\}$ be the standard basis in $V$ where
+where $a=\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}})$.
+Let $\rho$ be the standard representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Let $\{e_{1},e_{2},...,e_{n}\}$ be the standard basis in $V$ where
 $e_{1}=\left(\begin{array}[]{c}1\\ 0\\ 0\\ \vdots\\ 0\end{array}\right)\quad e_{2}=\left(\begin{array}[]{c}0\\ 1\\ 0\\ \vdots\\ 0\end{array}\right)\quad\cdots\cdots\quad e_{n}=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\\ 1\end{array}\right).$
@@
 $\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{cccc}*&*&\cdots&0\\ &*&\cdots&0\\ \vdots&\vdots&\cdots&\vdots\\ &*&\cdots&*\end{array}\right)\textup{ and }R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\left(\begin{array}[]{cc}I_{n-1}&*\\ 0&1\end{array}\right).$
-Suppose that $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ is a rational element in $\mathbf{G}(\mathbb{R})$. Let
-$x=\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)$
-where $p_{i}\in\mathbb{Z},q\in\mathbb{N}$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$.
-Then $x\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}\neq\{0\}$ implies that there exists a primitive integral point $(k_{1},k_{2},\dots,k_{n-1},l)^{T}\in\mathbb{Z}^{n}\setminus\{0\}$ such that
-$\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)\cdot\left(\begin{array}[]{c}k_{1}\\ k_{2}\\ \vdots\\ l\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ \end{array}\right).$
-Then we have
-$k_{i}+\frac{p_{i}}{q}\cdot l=0,\quad 1\leq i\leq n-1.$
-Hence $|l|=q$ and the denominator $d(x)$ of $x$ is equal to $q$.
-One can then deduce from the formula $d(x)=q$ that for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and any sufficiently large $l>0$, the number of rational elements in $\tilde{U}$ whose denominators are less than $l$ is bounded by $\mu_{R_{u}(\overline{\mathbf{P}}_{\beta_{0}})}(\tilde{U})\cdot l^{n}.$ Therefore, we may choose $A_{e}=n$ in Equation ($\ast$6) in §6.
-One can also compute that $a_{e}=n-1$.
-Now consider the upper bound of the Hausdorff dimension of $S_{\tau}^{c}$. In the computation, we actually use the Bruhat decomposition as
-$\mathbf{G}(\mathbb{R})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
-In our case, we may choose the index set $\bar{I}$ as small as possible (as the double cosets in the original decomposition with respect to Weyl group ${}_{\mathbb{Q}}W$ overlap).
-Note that ${}_{\mathbb{Q}}W$ is isomorphic to the symmetric group of permutations of $\{1,2,\dots,n\}$ and many elements of ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}(\mathbb{R})$.
-From this observation, one can deduce that if $w_{i}=(i,n)$ $(1\leq i\leq n)$ is the permutation of $i$ and $n$, then
-$\mathbf{G}(\mathbb{R})=\bigcup_{i=1}^{n}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
-Then by definition, one may compute that $A_{w_{i}}\leq A_{e}$, $a_{w_{i}}\leq a_{e}$ for all $1\leq i\leq n$, and
-$\mathbf{F}_{w_{i}}=\left(\begin{array}[]{cccccc}1&0&\cdots&*&\cdots&0\\ 0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&\cdots&1\end{array}\right)$
-where the only non-zero entries are in the $i$-th column.
-Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$.
-Fix $w_{i}$ $(1\leq i\leq n)$.
-Note that the roots in $\mathbf{F}_{w_{i}}$ are $\alpha_{1,i},\alpha_{2,i},\dots,\alpha_{i-1,i}$.
-According to Theorem 1.4 and the discussion in §6, we may assume that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0$, i.e.
-$b_{i}<0$; otherwise there is nothing to compute.
-Then $\nu_{0}=\alpha_{1,n}$ and by the fact that $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$ we have
-$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}$
-$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+\frac{-b_{i}-\tau}{b_{1}-b_{n}}\cdot\max\{A_{w_{i}},a_{w_{i}}\}$
-$\leq$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+n\cdot\frac{-b_{i}-\tau}{b_{1}-b_{n}}$
-$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})+nb_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
-$=$ $\dim\mathbf{G}-\frac{-b_{i+1}\cdots-b_{n}+(n-i)b_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
-$=$ $\dim\mathbf{G}-\frac{(b_{i}-b_{i+1})+\cdots(b_{i}-b_{n})}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
-$\leq$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}.$
-Therefore, by Theorem 1.4, we have
-$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
-On the other hand, for the lower bound of $\dim_{H}S_{\tau}^{c}$, by Theorem 1.3, we have
-$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
-$=$ $\dim\mathbf{G}-((b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n}))\cdot\frac{\tau}{(-b_{n})(b_{1}-b_{n})}$
-$=$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
-We conclude that
-$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
-This completes the proof of Theorem 1.6. ∎
-Proof of Theorem 1.7.
-Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and we may write
-$\{a_{t}\}_{t\in\mathbb{R}}=\left\{\operatorname{diag}(e^{b_{1}t},e^{b_{2}t},\cdots,e^{b_{n}t}):t\in\mathbb{R}\right\}$
-where $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$. We write
-$\mathbf{T}(\mathbb{R})^{0}=\{\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}}):t_{1}+t_{2}+\cdots+t_{n}=0\}$
-where $\mathbf{T}(\mathbb{R})^{0}$ is the connected component of identity in $\mathbf{T}(\mathbb{R})$.
-Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$ and $R_{u}(\overline{\mathbf{P}}_{0})$ is the upper triangular unipotent subgroup in $\mathbf{G}$.
-Let $V_{\beta_{0}}$ be the highest weight space in $\mathfrak{sl}_{n}$
+One may verify that $\beta_{0}=(-\frac{1}{n})\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha$. Theorem 1.7 follows from Theorem 1.6.
+Now consider the adjoint representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Let $V_{\beta_{0}}$ be the highest weight space in $\mathfrak{sl}_{n}$
 $V_{\beta_{0}}=\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ &0&\cdots&0\end{array}\right).$
-Note that here $\nu_{0}=-\beta_{0}$ is the highest root in $\mathfrak{sl}_{n}$ and $\beta_{0}(a_{t})=(b_{n}-b_{1})t<0$.
+Note that here $\nu_{0}=-\beta_{0}$ is the highest root in $\mathfrak{sl}_{n}$ and $\beta_{0}(a_{t})=(b_{n}-b_{1})t<0$ $(t>0)$.
 The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are
 $\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{ccccc}*&0&\cdots&0&0\\ &*&\cdots&*&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ &*&\cdots&*&0\\ &*&\cdots&*&*\end{array}\right)\textup{ and }R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=\left(\begin{array}[]{ccccc}1&*&\cdots&*&*\\ 0&1&\cdots&0&*\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&*\\ 0&0&\cdots&0&1\end{array}\right).$
-Let $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in $\mathbf{G}(\mathbb{R})$. We write
-$x=\left(\begin{array}[]{ccccc}1&\frac{a_{1}}{b}&\cdots&\frac{a_{n-2}}{b}&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&0&\cdots&0&\frac{p_{1}}{q}\\ 0&1&\cdots&0&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&\frac{p_{n-1}}{q}\\ 0&0&\cdots&0&1\end{array}\right)$
-where $(a_{1},a_{2},\dots,a_{n-2},b)=1$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then by definition,
-$x\cdot\mathfrak{sl}_{n}(\mathbb{Z})\cdot x^{-1}\cap V_{\beta_{0}}\neq\{0\}.$
-It implies that there exists $l\in\mathbb{R}$ such that
-$x^{-1}\cdot\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ l&0&\cdots&0\end{array}\right)\cdot x\in\mathfrak{sl}_{n}(\mathbb{Z}).$
-One can compute that
-$\left(\begin{array}[]{ccccc}-\frac{p_{1}}{q}\cdot l&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ -\frac{p_{n-1}}{q}\cdot l&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ l&l\cdot\frac{a_{1}}{b}&\cdots&l\cdot\frac{a_{n-2}}{b}&l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\end{array}\right)\in\mathfrak{sl}_{n}(\mathbb{Z})$
-and consequently
-$\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q},\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{a_{i}}{b}(1\leq i\leq n-2),\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{d}{bq}\in\mathbb{Z}$
-$l,\;l\cdot\frac{a_{1}}{b},\dots,l\cdot\frac{a_{n-2}}{b},\;l\cdot\frac{d}{bq}\in\mathbb{Z}$
-where $d=bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}$. Then one can deduce from the formulas above that
-$l=s\cdot\frac{bq^{2}}{\gcd(d,q)}\;(s\in\mathbb{Z}).$
-Therefore, the denominator $d(x)$ of $x$ is equal to
-$d(x)=\frac{bq^{2}}{\gcd(d,q).}$
-It follows from the same argument as in [17, Lemma 8.6] that
-|$\left|\{x\in\tilde{U}:d(x)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{n-1+\epsilon}$
-for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$, and hence $A_{e}=n-1$.
-Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$ using Theorem 1.4.
-We choose elements $\{w_{i}\}_{i\in\bar{I}}$ in the Weyl group ${}_{\mathbb{Q}}W$ such that the Bruhat decomposition stated in Theorem 1.4 holds
-$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
-Note that many Weyl elements in ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}$.
-Here we choose $\{w_{i}\}_{i\in\bar{I}}$ as follows.
-We identify ${}_{\mathbb{Q}}W$ with the symmetric group of permutations on $\{1,2,...,n\}$.
-Then there are three types of elements in $\{w_{i}\}_{i\in\bar{I}}$:
-Type I:
-$(1,2,\dots,n-1,n)$, $(n,2,\dots,n-1,1)$
-Type II:
-$(1,2\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$, $(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$, $(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1)
-Type III:
-$(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1).
-The index set $\bar{I}$ has cardinality $n(n-1)$, and one may check that $\{w_{i}\}_{i\in\bar{I}}$ meets our requirement. Note that $A_{w_{i}}\leq A_{e}$ for any $i\in\bar{I}$.
-For a Weyl element $w$ of type I in $\{w_{i}\}_{i\in\bar{I}}$, if $w=(1,2,\dots,n)$ is the identity element, then the number $a_{w}$ in Corollary 6.8 comes from the morphism $\Psi_{w}$ defined by
-$\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}0&0&\cdots&0&0\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ 1&0&\cdots&0&0\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)^{-1}$
-$=$ $\left(\begin{array}[]{ccccc}z&-x_{1}\cdot z&\cdots&-x_{n-2}\cdot z&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)z\\ y_{1}&-x_{1}\cdot y_{1}&\cdots&-x_{n-2}\cdot y_{1}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ y_{n-2}&-x_{1}\cdot y_{n-2}&\cdots&-x_{n-2}\cdot y_{n-2}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{n-2}\\ 1&-x_{1}&\cdots&-x_{n-2}&x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z\end{array}\right)$
-where $x_{1},x_{2},\dots,x_{n-2},y_{1},y_{2},\dots,y_{n-2},z$ are the variables of $\Psi_{w}$.
-One can estimate $\operatorname{Vol}(E_{w}(R))$ and compute that $a_{w}\leq n-1$.
-Then according to Theorem 1.4, for $w$, we have
-$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
-$=$ $\dim\mathbf{G}-(n-1)+\frac{(b_{1}-b_{n})-\tau}{b_{1}-b_{n}}\cdot(n-1)=\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
-Note that for $w=(n,2,3,\cdots,n-1,1)$, $\beta_{0}(wa_{-1}w^{-1})\leq 0$ and there is nothing to compute in this case.
-Now let us check a Weyl element $w$ of type II in $\{w_{i}\}_{i\in\bar{I}}$.
-If $w=(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$ or $w=(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1), then $\beta_{0}(w\cdot a_{-1}\cdot w^{-1})\leq 0$ and there is nothing to compute.
-If $w=(1,2,\dots,j-1,n,j+1,\dots,n-1,j)$ $(2\leq j\leq n-1)$, one can compute that
-$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccc}1&*&*&\cdots&*&\cdots&*\\ 0&1&0&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&0&\cdots&1\end{array}\right).$
-One can then deduce that $a_{w}\leq n-1$, which is similar to the type I case. Then under the condition that $(\beta_{0}(wa_{-1}w^{-1})>\tau$ we have
-$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
-$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{1}-b_{2})+(b_{1}-b_{3})+\cdots+(b_{1}-b_{n})}{b_{1}-b_{n}}+\frac{(b_{2}-b_{j})+(b_{3}-b_{j})+\cdots+(b_{j-1}-b_{j})}{b_{1}-b_{n}}\right)$
-$\hskip 28.45274pt+\frac{(b_{1}-b_{j})-\tau}{b_{1}-b_{n}}\cdot(n-1)$
-$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{j+1})+(b_{1}-b_{j+2})+\cdots+(b_{1}-b_{n})-(n-j)(b_{1}-b_{j})}{b_{1}-b_{n}}-\frac{\tau}{b_{1}-b_{n}}\cdot(n-1)$
-$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
-Similarly, for $w=(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$ ($12\leq j\leq n-$ 1),
-$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccc}1&0&\cdots&0&0&\cdots&*\\ 0&1&\cdots&0&0&\cdots&*\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&*&\cdots&*\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&1\end{array}\right)$
-and $a_{w}\leq n-1$. So under the condition that $(\beta_{0}(wa_{-1}w^{-1})>\tau$ we have
-$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
-$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n})}{b_{1}-b_{n}}+\frac{(b_{j}-b_{j+1})+(b_{j}-b_{j+2})+\cdots+(b_{j}-b_{n-1})}{b_{1}-b_{n}}\right)$
-$\hskip 28.45274pt+\frac{(b_{j}-b_{n})-\tau}{b_{1}-b_{n}}\cdot(n-1)$
-$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{j}-b_{n})-j(b_{j}-b_{n})}{b_{1}-b_{n}}-\frac{\tau}{b_{1}-b_{n}}\cdot(n-1)$
-$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
-Let us check a Weyl element $w$ of type III. If $w=(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1), then
-$\beta_{0}(w\cdot a_{-1}\cdot w^{-1})\leq 0$
-and there is nothing to prove. If $w=(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, then one can compute that
-$\mathbf{F}_{w}=\left(\begin{array}[]{ccccccccc}1&0&\cdots&0&0&\cdots&*&\cdots&0\\ 0&1&\cdots&0&0&\cdots&*&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1&*&\cdots&*&\cdots&*\\ 0&0&\cdots&0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0&\cdots&0&\cdots&1\end{array}\right)$
-and $a_{w}\leq n-1$. Then under the condition that $\beta_{0}(wa_{-1}w^{-1})>\tau$, we have
-$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w},a_{w}\}\cdot\dim V_{\beta_{0}}$
-$\leq$ $\dim\mathbf{G}-\left(\frac{(b_{i}-b_{i+1})+(b_{i}-b_{i+2})+\cdots+(b_{i}-b_{n})}{\nu_{0}(a_{1})}+\frac{(b_{1}-b_{j})+(b_{2}-b_{j})+\cdots+(b_{j-1}-b_{j})}{\nu_{0}(a_{1})}-\frac{b_{i}-b_{j}}{\nu_{0}(a_{1})}\right)$
-$\hskip 28.45274pt+\frac{(b_{i}-b_{j})-\tau}{\nu_{0}(a_{1})}\cdot(n-1)$
-$=$ $\dim\mathbf{G}-\frac{(n-i)b_{i}+b_{1}+\cdots+b_{i}-b_{j}-\cdots-b_{n}-(j-1)b_{j}-n(b_{i}-b_{j})}{\nu_{0}(a_{1})}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}$
-$\leq$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
-Combining type I, type II, type III cases, we conclude that
-$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
-On the other hand, by Theorem 1.3, we know that
-$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
-$=$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
-$=$ $\dim\mathbf{G}-\frac{(n-1)\tau}{\nu_{0}(a_{1})}.$
-This completes the proof of Theorem 1.7. ∎
-Proof of Theorem 1.8.
-It is known that there is a complete classification of complex-linear finite-dimensional irreducible representations of $\mathbf{G}=\operatorname{\mathbf{SL}}_{2}$ described as follows.
-Let $V_{n}$ be the complex vector space of homogeneous polynomials of degree $n$ in two variables $x$ and $y$ ($\dim V_{n}=n+1$).
-Define $\rho_{n}:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{SL}}_{n+1}(V_{n})$ by
-$\rho(g)\cdot f\left(\begin{array}[]{c}x\\ y\end{array}\right):=f\left(g^{-1}\cdot\left(\begin{array}[]{c}x\\ y\end{array}\right)\right)$
-for any $g\in\operatorname{\mathbf{SL}}_{2}$ and $f\in V_{n}$.
-Then $\{\rho_{n}\}_{n\in\mathbb{N}}$ consists of all the complex-linear finite dimensional irreducible representations of $\operatorname{\mathbf{SL}}_{2}$ up to equivalence.
-Without loss of generality, we may assume that $\rho=\rho_{n}$ for some $n\in\mathbb{N}$ and write $\{a_{t}\}_{t\in\mathbb{R}}\subset\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ as
-$\left\{a_{t}=\left(\begin{array}[]{cc}e^{\alpha t}&0\\ 0&e^{-\alpha t}\end{array}\right):t\in\mathbb{R}\right\}$
-for some $\alpha>0$.
-Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$.
-The standard basis of $V_{n}$ is $\{x^{n},x^{n-1}y,\cdots,xy^{n-1},y^{n}\}.$ We choose $\mathbb{C}\cdot x^{n}$ to be the highest weight space $V_{\beta_{0}}$ in $V_{n}$ where $\beta_{0}(a_{t})=-n\alpha\cdot t$.
-The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ is the subgroup
-$\mathbf{P}_{\beta_{0}}=\left(\begin{array}[]{cc}*&0\\ &*\end{array}\right)$
-which coincides with the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ of $\operatorname{\mathbf{SL}}_{2}$, and
-$R_{u}(\overline{\mathbf{P}}_{\beta_{0}})=R_{u}(\overline{\mathbf{P}}_{0})=\left(\begin{array}[]{cc}1&*\\ 0&1\end{array}\right).$
-The Bruhat decomposition of $\operatorname{\mathbf{SL}}_{2}$ is the following
-$\operatorname{\mathbf{SL}}_{2}(\mathbb{R})=\mathbf{P}_{\beta_{0}}\cdot w_{0}\cdot R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cup\mathbf{P}_{\beta_{0}}\cdot w_{1}$
-where $w_{0}$ is the identity element
-$w_{1}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$
-and ${}_{\mathbb{Q}}W=\{w_{0},w_{1}\}$.
-Now we consider the upper bound of $\dim_{H}S_{\tau}^{c}$. Let $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and write
-$g=\left(\begin{array}[]{cc}1&t\\ 0&1\end{array}\right).$
-Then one can compute that
-$\rho(g)\cdot(x^{n},x^{n-1}y,\dots,y^{n})$
-$=$ $(x^{n},x^{n-1}y,\dots,y^{n})\cdot\left(\begin{array}[]{cccccc}{n\choose n}(-t)^{0}&0&\cdots&0&\cdots&0\\ {n\choose n-1}(-t)^{1}&{n-1\choose n-1}(-t)^{0}&\cdots&0&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose i}(-t)^{i}&{n-1\choose i}(-t)^{i-1}&\cdots&{i\choose i}(-t)^{0}&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose 0}(-t)^{n}&{n-1\choose 0}(-t)^{n-1}&\cdots&{i\choose 0}(-t)^{i}&\cdots&{0\choose 0}(-t)^{0}\end{array}\right).$
-By Corollary 6.8 and the formula above, we have $a_{w_{0}}=1/n$ for the morphism $\Psi_{w_{0}}$ defined by
-$\rho(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R}))\cdot\{x^{n}\}.$
-Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ with
-$g=\left(\begin{array}[]{cc}1&\frac{p}{q}\\ 0&1\end{array}\right)\quad(p\in\mathbb{Z},\;q\in\mathbb{N},\;\gcd(p,q)=1).$
-Then $\rho(g)\cdot\mathbb{Z}^{n+1}\cap V_{\beta_{0}}\neq\{0\}$ and there exists a primitive integral vector $(k_{n},k_{n-1},\dots,k_{1},l)\in\mathbb{Z}^{n+1}\setminus\{0\}$ such that
-($\left(\begin{array}[]{cccccc}{n\choose n}(-\frac{p}{q})^{0}&0&\cdots&0&\cdots&0\\ {n\choose n-1}(-\frac{p}{q})^{1}&{n-1\choose n-1}(-\frac{p}{q})^{0}&\cdots&0&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose i}(-\frac{p}{q})^{n-i}&{n-1\choose i}(-\frac{p}{q})^{n-i-1}&\cdots&{i\choose i}(-\frac{p}{q})^{0}&\cdots&0\\ \vdots&\vdots&\cdots&\vdots&\cdots&\vdots\\ {n\choose 0}(-\frac{p}{q})^{n}&{n-1\choose 0}(-\frac{p}{q})^{n-1}&\cdots&{i\choose 0}(-\frac{p}{q})^{i}&\cdots&{0\choose 0}(-\frac{p}{q})^{0}\end{array}\right)\cdot\left(\begin{array}[]{c}l\\ k_{1}\\ \vdots\\ \vdots\\ \vdots\\ k_{n}\end{array}\right)=\left(\begin{array}[]{c}*\\ 0\\ \vdots\\ 0\\ \vdots\\ 0\end{array}\right).$
-One can deduce from the equation above that
-$l\cdot{n\choose i}\cdot\left(\frac{p}{q}\right)^{n-i}=k_{n-i}\;(\forall 1\leq i\leq n),\quad|l|=q^{n}$
-and $d(g)=q^{n}$. According to equation ($\ast$6) in §6, one can compute that $A_{1}=2/n$. Note that for the Weyl element $w_{1}$, $\beta_{0}(w_{1}a_{-1}w_{1}^{-1})<0$. So by Theorem 1.4, we obtain
-$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-1+\frac{n\alpha-\tau}{2\alpha}\cdot\frac{2}{n}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
-On the other hand, by Theorem 1.3, we have
-$\dim_{H}S_{\tau}^{c}\geq\dim\mathbf{G}-\frac{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
-Therefore, we can conclude that $\dim_{H}S_{\tau}^{c}=3-\tau/\beta_{0}(a_{-1})$. ∎
-Let us explain how to deduce [17, Corollary 1.3] (or equivalently [17, Theorem 1.2]) from Theorem 1.7.
+One may verify that $\beta_{0}=(-\frac{1}{n-1})\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha$. Theorem 1.8 then follows from Theorem 1.6. ∎
+Let us explain how to deduce [25, Corollary 1.3] (or equivalently [25, Theorem 1.2]) from Theorem 1.8.
 In [25], we consider a regular one-parameter diagonal subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ acting on the homogeneous space $X_{3}=\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$.
-According to [17, Definition 1.1], a point $p=g\cdot\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\in X_{3}$ is Diophantine of type $\tau$ if and only if there exists a constant $C>0$ such that for any $t>0$
+According to [25, Definition 1.1], a point $p=g\cdot\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\in X_{3}$ is Diophantine of type $\tau$ if and only if there exists a constant $C>0$ such that for any $t>0$
 $\eta(a_{t}\cdot p)\geq Ce^{-\tau t}$
@@
 $\eta(a_{t_{\epsilon}}\cdot p)<\epsilon e^{-\tau t_{\epsilon}}.$
-By [37, Corollary 11.18], for sufficiently small $\epsilon>0$, there exists a unipotent element $u_{\epsilon}\in\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\setminus\{e\}$ such that $g\cdot u_{\epsilon}\cdot g^{-1}\in\operatorname{Stab}(p)=g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})g^{-1}$ (the stabilizer of $p$) and
+By [69, Corollary 11.18], for sufficiently small $\epsilon>0$, there exists a unipotent element $u_{\epsilon}\in\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})\setminus\{e\}$ such that $g\cdot u_{\epsilon}\cdot g^{-1}\in\operatorname{Stab}(p)=g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})g^{-1}$ (the stabilizer of $p$) and
 $d_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g\cdot u_{\epsilon}\cdot g^{-1}a_{-t_{\epsilon}},e)<\epsilon e^{-\tau t_{\epsilon}}$
@@
 $\|\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g)(2\log u_{\epsilon})\|_{\mathfrak{sl}_{3}}<\tilde{C}\epsilon e^{-\tau t_{\epsilon}}$
-where $\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}$ is the adjoint representation of $\operatorname{\mathbf{SL}}_{3}$.
-This implies that $g\in S_{\tau}(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}},\{a_{t}\}_{t\in\mathbb{R}})^{c}$ by Definition 1.1.
-Conversely, let $g\in\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$ such that $g\in S_{\tau}(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}},\{a_{t}\}_{t\in\mathbb{R}})^{c}$.
+where $\rho=\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}$ is the adjoint representation of $\operatorname{\mathbf{SL}}_{3}$.
+This implies that $g\in S_{\rho}(\psi)^{c}$ for $\psi(t)=e^{-\tau t}$ by Definition 1.1.
+Conversely, let $g\in\operatorname{\mathbf{SL}}_{3}(\mathbb{R})$ such that $g\in S_{\rho}(\psi)^{c}$ where $\rho=\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}$ and $\psi(t)=e^{-\tau t}$.
 Then by Definition 1.1, for any $\epsilon>0$, there exists $t_{\epsilon}>0$ such that
 $\delta(\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}\cdot g)\mathfrak{sl}_{3}(\mathbb{Z}))<\epsilon\cdot e^{-\tau t_{\epsilon}}.$
-By [42, Proposition 3.3], for sufficiently small $\epsilon>0$, there exists a nilpotent element $n_{\epsilon}\in\mathfrak{sl}_{3}(\mathbb{Z})\setminus\{0\}$ such that
+By [78, Proposition 3.3], for sufficiently small $\epsilon>0$, there exists a nilpotent element $n_{\epsilon}\in\mathfrak{sl}_{3}(\mathbb{Z})\setminus\{0\}$ such that
 $\|\operatorname{Ad}_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}\cdot g)n_{\epsilon}\|_{\mathfrak{sl}_{3}}<\epsilon\cdot e^{-\tau t_{\epsilon}}$
@@
 $d_{\operatorname{\mathbf{SL}}_{3}}(a_{t_{\epsilon}}g\cdot\exp(2n_{\epsilon})g^{-1}a_{-t_{\epsilon}},e)<\tilde{C}\cdot\epsilon\cdot e^{-\tau t_{\epsilon}}.$
-By [17, Definition 1.1], this implies that $g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ is not Diophantine of type $\tau$.
-Consequently, Theorem 1.7 and Remark 1.10 imply [17, Corollary 1.3] when $\mathbf{G}=\operatorname{\mathbf{SL}}_{3}$.
+By [25, Definition 1.1], this implies that $g\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ is not Diophantine of type $\tau$.
+Consequently, Theorem 1.8 and Remark 1.9 imply [25, Corollary 1.3] when $\mathbf{G}=\operatorname{\mathbf{SL}}_{3}$.
+Remark 7.3.
+The arithmetic case in [82, Theorem 1.1] concerns diagonalizable flows on $\mathbf{G}(\mathbb{R})/\mathbf{G}(\mathbb{Z})$ where $\mathbf{G}$ is a simple $\mathbb{Q}$-group with $\mathbb{Q}$-rank and $\mathbb{R}$-rank both equal to one.
+Using the same argument as above, we can also derive this arithmetic case from Theorem 1.6 with the adjoint representation of $\mathbf{G}$.
+8. Connections to Diophantine approximation
+In this section, we prove Theorems 1.12 and 1.15.
+8.1. Connection to Diophantine approximation on flag varieties
+We first prove Theorem 1.12.
+To study the Diophantine approximation on $\mathbf{X}(\mathbb{R})$, or specifically, to study the Diophantine subset $E_{\mathbf{X}}(\psi)$, we need to study the $\{a_{t}\}_{t\in\mathbb{R}}$-action on the space $\mathcal{X}$ of lattices in $V(\mathbb{R})$ (see [17, Chapter 1, Section 2]).
+For any $x\in\mathbf{X}(\mathbb{R})$, write $x=\mathbf{P}(\mathbb{R})\cdot s_{x}$ for some $s_{x}\in\mathbf{G}(\mathbb{R})$ and define $\Lambda_{x}=s_{x}\cdot\mathbb{Z}^{d}\in\mathcal{X}$.
+Note that for any $g\in\mathbf{P}(\mathbb{R})$, the set $\{a_{t}\cdot g\cdot a_{-t}\}_{t\geq 0}$ is bounded in $\mathbf{G}(\mathbb{R})$.
+So for any $s_{1}$ and $s_{2}$ with $\mathbf{P}(\mathbb{R})\cdot s_{1}=\mathbf{P}(\mathbb{R})\cdot s_{2}$, the forward orbits $\{a_{t}\cdot s_{1}\mathbb{Z}^{d}\}_{t\geq 0}$ and $\{a_{t}\cdot s_{2}\mathbb{Z}^{d}\}_{t\geq 0}$ in $\mathcal{X}$ stay within bounded distance from each other for any $t>0$, and the behaviors of $\{a_{t}\cdot s_{1}\mathbb{Z}^{d}\}_{t\geq 0}$ and $\{a_{t}\cdot s_{2}\mathbb{Z}^{d}\}_{t\geq 0}$ are almost identical.
+Let $\mathcal{C}_{\chi}:=\mathbf{G}(\mathbb{R})\cdot V_{\chi}\setminus\{0\}$.
+Let $\pi_{\chi}:V\to V_{\chi}$ be the projection from $V$ to $V_{\chi}$ along all the other weight spaces in $V$.
+For any lattice $\Lambda\in\mathcal{X}$, define
+$\delta_{\chi}(\Lambda)=\inf\left\{\|v\|:v\in\Lambda\cap\mathcal{C}_{\chi},\|\pi_{\chi}(v)\|\geq\frac{1}{2}\|v\|\right\}.$
+We will need the following important correspondence theorem.
+Proposition 8.1 ([17, Proposition 3.2.4]).
+Let $x\in\mathbf{X}(\mathbb{R})$ and $s_{x}\in\mathbf{G}(\mathbb{R})$ such that $x=\mathbf{P}(\mathbb{R})\cdot s_{x}$.
+Let $\beta_{\chi}>0$ be the Diophantine exponent defined in Theorem 1.10.
+Then there exists a constant $C>0$ such that the following holds.
+Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a decreasing function, and $\Psi$ the function defined by
+$\Psi(u)=Cu^{-\beta_{\chi}}\psi(u).$
+If the inequality $d_{\textup{CC}}(x,v)\leq H_{\chi}(v)^{-\beta_{\chi}}\psi(H_{\chi}(v))$ admits infinitely many solutions $v\in\mathbf{X}(\mathbb{Q})$, then there exists arbitrarily large $t>0$ such that
+$\delta_{\chi}(a_{t}s_{x}\mathbb{Z}^{d})\leq 2e^{-t/\beta_{\chi}}\Psi^{-1}(e^{-t}).$
+If we have that there exists $t>0$ arbitrarily large such that
+$\delta_{\chi}(a_{t}s_{x}\mathbb{Z}^{d})\leq e^{-t/\beta_{\chi}}\Psi^{-1}(e^{-t})$
+then the inequality
+$d_{\textup{CC}}(x,v)\leq C^{2}H_{\chi}(v)^{-\beta_{\chi}}\psi(H_{\chi}(v))$
+admits infinitely many solutions $v\in\mathbf{X}(\mathbb{Q})$.
+Proof of Theorem 1.12.
+Since $\psi$ is decreasing, we have $\psi(t)\leq\psi(1)$ $(t\geq 1)$ which implies that $\gamma(\psi)\geq 0$.
+We first compute an upper bound for the Hausdorff dimension of $E_{\mathbf{X}}(\psi)^{c}$.
+Let $x\in E_{\mathbf{X}}(\psi)^{c}$.
+From Definition 1.11, one may deduce that there exist infinitely many $v\in\mathbf{X}(\mathbb{Q})$ such that
+$d_{\textup{CC}}(x,v)\leq(H_{\chi}(v))^{-\beta_{\chi}}\psi(H_{\chi}(v)).$
+By Proposition 8.1, this implies that there exists $t>0$ arbitrarily large such that
+$\delta_{\chi}(a_{t}s_{x}\mathbb{Z}^{d})\leq 2e^{-t/\beta_{\chi}}\Psi^{-1}(e^{-t}).$
+Note that $\delta\leq\delta_{\chi}$. So by Definition 1.1, for any $\epsilon>0$, $s_{x}\in S_{\rho}(2e^{-t(1-\epsilon)/\beta_{\chi}}\Psi^{-1}(e^{-t}))^{c}$. This implies that
+$\dim\mathbf{P}+\dim_{H}E_{\mathbf{X}}(\psi)^{c}\leq\dim_{H}S_{\rho}(2e^{-t(1-\epsilon)/\beta_{\chi}}\Psi^{-1}(e^{-t}))^{c}.$
+By Theorem 1.6, we have
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}\leq\dim\mathbf{X}-\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}\alpha(a_{-1})$
+where $\tau=\tau(2e^{-t(1-\epsilon)/\beta_{\chi}}\Psi^{-1}(e^{-t}))$. One may compute that
+$\tau=\frac{1-\epsilon}{\beta_{\chi}}-\frac{1}{\gamma(\psi)+\beta_{\chi}}.$
+Let $\epsilon\to 0$ and we obtain
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}\leq\dim\mathbf{X}-\left(\frac{1}{\beta_{\chi}}-\frac{1}{\gamma(\psi)+\beta_{\chi}}\right)\frac{1}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}\alpha(a_{-1}).$
+Now we compute a lower bound for the Hausdorff dimension of $E_{\mathbf{X}}(\psi)^{c}$.
+Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ with the same Levi factor $Z(\mathbf{T})$, and denote by $R_{u}(\overline{\mathbf{P}}_{0})$ the unipotent radical of $\overline{\mathbf{P}}_{0}$.
+In section 4, for any function $\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $\tau(\phi)\in[0,\chi(a_{-1}))$ and any sufficiently small $\epsilon>0$, we construct a Cantor type subset $\mathbf{A}_{\infty}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ such that for any $p\in\mathbf{A}_{\infty}$ we have
+$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\lesssim\phi(t_{k})\cdot e^{-\frac{\epsilon}{2}t_{k}}$
+for some divergent sequence $\{t_{k}\}_{k\in\mathbb{N}}$. Note that in the construction
+$a_{t_{k}}p\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}q_{k}$
+and there exist $v_{k}\in\rho(a_{t_{k}}q_{k})\mathbb{Z}^{d}\cap V_{\chi}$ $(V_{\chi}=V_{\beta_{0}})$ and $x_{k}\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})$ such that
+$\rho(x_{k})v_{k}\in\rho(a_{t_{k}}p)\mathbb{Z}^{d}$
+and
+$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})\leq\|\rho(x_{k})v_{k}\|\asymp e^{\chi(a_{t_{k}})}\cdot l_{k}\leq\phi(t_{k})\cdot e^{-\frac{\epsilon}{2}t_{k}}.$
+One may verify that for small $\tilde{\delta}_{0}>0$ we have
+$\|\pi_{\chi}(\rho(x_{k})v_{k})\|\geq\frac{1}{2}\|\rho(x_{k})v_{k}\|.$
+This implies that for any $p\in\mathbf{A}_{\infty}$
+$\delta_{\chi}(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})\lesssim\phi(t_{k})\cdot e^{-\frac{\epsilon}{2}t_{k}}.$
+Now we may choose
+$\phi(t)=2e^{-t/\beta_{\chi}}\Psi_{\epsilon}^{-1}(e^{-t})$
+where $\Psi_{\epsilon}(u)=Cu^{-\beta_{\chi}}\psi(u)u^{-\epsilon}$.
+By Proposition 8.1, for $x=\mathbf{P}(\mathbb{R})g$ with $g\in\mathbf{A}_{\infty}$ there exist infinitely many $v\in\mathbf{X}(\mathbb{Q})$ such that
+$d_{\textup{CC}}(x,v)\leq C^{2}H_{\chi}(v)^{-\beta_{\chi}}\psi(H_{\chi}(v))(H_{\chi}(v))^{-\epsilon}$
+which implies that $x\in E_{\mathbf{X}}(\psi)^{c}$. So we have
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}+\dim\mathbf{P}\geq\dim_{H}\mathbf{A}_{\infty}+\dim\mathbf{P}_{0}$
+and
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}\geq\dim\mathbf{X}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})(\tau+\epsilon)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
+where $\tau=\tau(\phi)=\tau(2e^{-t/\beta_{\chi}}\Psi_{\epsilon}^{-1}(e^{-t}))$. One may compute that
+$\tau=\tau(\phi)=\frac{1}{\beta_{\chi}}-\frac{1}{\gamma(\psi)+\epsilon+\beta_{\chi}}$
+and when $\gamma(\psi)<\infty$, $\tau=\tau(\psi)<\chi(a_{-1})$. Let $\epsilon\to 0$ and we obtain that
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}\geq\dim\mathbf{X}-\left(\frac{1}{\beta_{\chi}}-\frac{1}{\gamma(\psi)+\beta_{\chi}}\right)\frac{1}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}\alpha(a_{-1}).$
+Note that if $\gamma(\psi)=\infty$, then $\tau(\phi)=\chi(a_{-1})$.
+In this case, $\phi(t)e^{\chi(a_{-1})t}=2\Psi_{\epsilon}^{-1}(e^{-t})$ is an unbounded function as $\psi$ is decreasing.
+So we can choose the Cantor-type subset $\mathbf{A}_{\infty}$ constructed in Proposition 5.6 and apply the argument above again to conclude that
+$\dim_{H}E_{\mathbf{X}}(\psi)^{c}\geq\dim\mathbf{X}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\phi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
+This completes the proof of Theorem 1.12. ∎
+8.2. Connection to rational approximation to linear subspaces
+Now we prove Theorem 1.15.
+Let $1\leq k\leq l<n$.
+We consider the representation $\rho_{k}:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ $(V=\bigwedge^{k}\mathbb{C}^{n})$ defined by the $k$-th exterior product of the standard representation $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(\mathbb{C}^{n})$ of $\operatorname{\mathbf{SL}}_{n}$.
+Note that the action of $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$ on the Grassmann variety of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is transitive.
+So the representation $\rho_{k}$ is irreducible.
+Let $\{e_{1},e_{2},\cdots,e_{n}\}$ be the standard basis in $\mathbb{C}^{n}$, where the $i$-th coordinate of $e_{i}$ equals one and the other coordinates of $e_{i}$ equals zero $(1\leq i\leq n)$.
+Then
+$\{e_{i_{1}}\wedge e_{i_{2}}\wedge\cdots\wedge e_{i_{k}}:1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n\}$
+forms an integral basis of $\bigwedge^{k}\mathbb{C}^{n}$. We choose
+$V_{\chi}=\mathbb{C}\cdot e_{\chi},\quad e_{\chi}=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{k}$
+to be the weight space of highest weight $\chi$ in $\bigwedge^{k}\mathbb{C}^{n}$.
+Denote by $\mathbf{P}_{\chi}$ the stablizer of the weight space $\mathbb{C}\cdot e_{\chi}$ and by $\pi_{\chi}:\bigwedge^{k}\mathbb{C}^{n}\to V_{\chi}$ the projection onto $V_{\chi}$ along the direct sum of all the other weight spaces.
+We also choose $\mathbf{P}_{0}$ to be the upper triangular subgroup, $\mathbf{T}$ the full diagonal subgroup and $\overline{\mathbf{P}}_{0}$ the lower triangular subgroup in $\operatorname{\mathbf{SL}}_{n}$.
+Note that $\mathbf{P}_{0}$ is a minimal parabolic $\mathbb{Q}$-subgroup contained in $\mathbf{P}_{\chi}$.
+We denote by $R_{u}(\mathbf{P}_{\chi})$ (resp.
+$R_{u}(\overline{\mathbf{P}}_{0})$) the unipotent radical of $\mathbf{P}_{\chi}$ (resp.
+$\overline{\mathbf{P}}_{0}$), and write $\Phi(R_{u}(\mathbf{P}_{\chi}))$ (resp.
+$\Phi(R_{u}(\overline{\mathbf{P}}_{0}))$) for the set of $\mathbb{Q}$-roots in $R_{u}(\mathbf{P}_{\chi})$ (resp.
+$R_{u}(\overline{\mathbf{P}}_{0})$) relative to $\mathbf{T}$.
+The symbol $\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}_{\chi}))}$ stands for the sum over all $\mathbb{Q}$-roots $\alpha\in\Phi(R_{u}(\mathbf{P}_{\chi}))$.
+One may verify that for the irreducible representation $\rho_{k}$, we have $\chi=n\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}_{\chi}))}\alpha$.
+So $\rho_{k}$ satisfies the assumption of Theorem 1.6.
+Let $x_{0}$ be the $l$-dimensional subspace $\mathbb{R}$-$\textup{span}\{e_{1},e_{2},\dots,e_{l}\}$ in $X_{l}(\mathbb{R})$, and $P$ the stabilizer of $x_{0}$ in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$.
+Then the Grassmann variety $X_{l}(\mathbb{R})$ can be identified with $P\backslash\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$ by
+$P\backslash\operatorname{\mathbf{SL}}_{n}(\mathbb{R})\mapsto X_{l}(\mathbb{R}),\quad P\cdot g\mapsto g^{-1}\cdot x_{0}.$
+Let
+$a_{t}=\operatorname{diag}(e^{-t/l},\cdots,e^{-t/l},e^{t/(n-l)},\cdots,e^{t/(n-l)})\subset\operatorname{\mathbf{SL}}_{n}(\mathbb{R}).$
+The matrix $a_{t}$ has an eigenvalue $e^{-\frac{k}{l}t}$ in $\bigwedge^{k}\mathbb{R}^{n}$.
+Denote by $\pi^{+}:\bigwedge^{k}\mathbb{R}^{n}\to\bigwedge^{k}\mathbb{R}^{n}$ the projection to the eigenspace of $a_{t}$ associated to the eigenvalue $e^{-\frac{k}{l}t}$.
+We write $\nu_{0}$ for any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $\Phi(R_{u}(\overline{\mathbf{P}}_{0}))$ satisfying
+$\nu_{0}(a_{1})=\max\{\alpha(a_{1}):\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{0}))\}.$
+To study the Diophantine subset $E_{l,k}(\psi)$ in $X_{l}(\mathbb{R})$, we need to study the $\{a_{t}\}_{t\in\mathbb{R}}$-action on the space of unimodular lattices in $\bigwedge^{k}\mathbb{R}^{n}$.
+The following is a correspondence theorem which is crucial in the analysis of $E_{l,k}(\psi)$.
+Proposition 8.2 ([16, Proposition 1]).
+Let $x\in X_{l}(\mathbb{R})$ and $s_{x}\in\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$ such that $x=P\cdot s_{x}$.
+  1. (1)
+    Let $v\in X_{k}(\mathbb{Q})$ be close to $x$, and $t>0$ such that $e^{-t(\frac{1}{l}+\frac{1}{n-l})}=d(v,x)$.
+Then the pure tensor $\mathbf{v}\in\bigwedge^{k}\mathbb{Z}^{n}$ associated to $v$ satisfies
+    $\|\pi^{+}(a_{t}s_{x}\mathbf{v})\|\gtrsim\|a_{t}s_{x}\mathbf{v}\|\textup{ and }\|a_{t}s_{x}\mathbf{v}\|\lesssim e^{-t\frac{k}{l}}H(v)$
+    where the implicit constants depend on the choice of $s_{x}$.
+  2. (2)
+    Let $t>0$ and $\mathbf{v}\in\bigwedge^{k}\mathbb{Z}^{n}$ a pure tensor such that
+    $\|\pi^{+}(a_{t}s_{x}\mathbf{v})\|\geq c\|a_{t}s_{x}\mathbf{v}\|$
+    for some fixed constant $c>0$. Then the rational subspace $v\in X_{k}(\mathbb{Q})$ assocaited to $\mathbf{v}$ satisfies
+    $H(v)\lesssim e^{\frac{k}{l}t}\|a_{t}s_{x}\mathbf{v}\|\textup{ and }d(v,x)\lesssim e^{-t(\frac{1}{l}+\frac{1}{n-l})}$
+    where the implicit constants depend on the choice of $s_{x}$ and $c$.
+Lemma 8.3.
+Let $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subset\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$ be a rational element in the sense of Definition 2.1.
+If $\mathbf{v}\in\bigwedge^{k}\mathbb{Z}^{n}\setminus\{0\}$ is a primitive integral vector in $\bigwedge^{k}\mathbb{R}^{n}$ such that $\rho_{k}(g)\mathbf{v}\in V_{\chi}$, then $\mathbf{v}$ is a pure tensor in $\bigwedge^{k}\mathbb{Z}^{n}$.
+Proof.
+Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
+Then by Corollary 2.3, we may write $g=h\cdot f$ for some $h\in\mathbf{H}_{e}(\mathbb{R})$ and $f\in\mathbf{F}_{e}(\mathbb{Q})$.
+Note that $h$ fixes every element in $V_{\chi}$.
+So we have $\rho_{k}(f)\mathbf{v}\in V_{\chi}(=\mathbb{C}\cdot e_{1}\wedge\cdots\wedge e_{k})$.
+We know that $\{\rho(f^{-1})e_{1},\rho(f^{-1})e_{2},\cdots,\rho(f^{-1})e_{k}\}$ spans a linear $\mathbb{Q}$-subspace $W_{f}$ in $\mathbb{R}^{d}$, and we may take an integral basis $\{f_{1},f_{2},\cdots,f_{k}\}\subset\mathbb{Z}^{n}$ for $W_{f}$.
+Then $f_{1}\wedge\cdots\wedge f_{k}\in\bigwedge^{k}\mathbb{Z}^{n}$ is a primitive integral vector in $\bigwedge^{k}\mathbb{R}^{n}$, and $f_{1}\wedge\cdots\wedge f_{k}\in\rho_{k}(f^{-1})V_{\chi}$.
+It implies that $\mathbf{v}=f_{1}\wedge\cdots\wedge f_{k}$.
+This completes the proof of the lemma.
+∎
+Proof of Theorem 1.15.
+Since $\psi$ is decreasing, we have $\psi(t)\leq\psi(1)$ $(t\geq 1)$ which implies that $\gamma(\psi)\geq 0$.
+We first compute an upper bound for the Hausdorff dimension of $E_{l,k}(\psi)^{c}$.
+Let $x\in E_{l,k}(\psi)^{c}$ and write $x=P\cdot s_{x}$ $(s_{x}\in\operatorname{\mathbf{SL}}_{n}(\mathbb{R}))$.
+From Definition 1.14, one may deduce that there exist infinitely many $v\in X_{k}(\mathbb{Q})$ such that
+$d(x,v)\leq H(v)^{-\frac{n}{k(n-l)}}\cdot\psi(H(v)).$
+By Proposition 8.2(1), for any such $v\in X_{k}(\mathbb{Q})$, there exists $t>0$ such that
+$e^{-t(\frac{1}{l}+\frac{1}{n-l})}=d(v,x)$
+and the pure tensor $\mathbf{v}\in\bigwedge^{k}\mathbb{Z}^{n}$ associated to $v$ satisfies
+$\|\pi^{+}(a_{t}s_{x}\mathbf{v})\|\gtrsim\|a_{t}s_{x}\mathbf{v}\|\textup{ and }\|a_{t}s_{x}\mathbf{v}\|\lesssim e^{-t\frac{k}{l}}H(v)$
+where the implicit constants depend on the choice of $s_{x}$. Let $\Psi(u)=u^{-\frac{n}{k(n-l)}}\cdot\psi(u)$. Then we have
+$e^{-t(\frac{1}{l}+\frac{1}{n-l})}=d(v,x)\leq\Psi(H(v))\implies H(v)\leq\Psi^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})})$
+and
+$\|a_{t}s_{x}\mathbf{v}\|\lesssim e^{-t\frac{k}{l}}H(v)\leq e^{-t\frac{k}{l}}\cdot\Psi^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})})$
+which implies that
+$\delta\left(a_{t}s_{x}\bigwedge^{k}\mathbb{Z}^{n}\right)\lesssim e^{-t\frac{k}{l}}\cdot\Psi^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})}).$
+So by Definition 1.1, for any $\epsilon>0$, $s_{x}\in S_{\rho_{k}}(e^{-t(1-\epsilon)\frac{k}{l}}\cdot\Psi^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})}))^{c}$. This implies that
+$\dim P+\dim_{H}E_{l,k}(\psi)^{c}\leq\dim_{H}S_{\rho_{k}}(e^{-t(1-\epsilon)\frac{k}{l}}\cdot\Psi^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})}))^{c}.$
+By Theorem 1.6, we have
+$\dim_{H}E_{l,k}(\psi)^{c}\leq\dim X_{l}-\frac{\tau}{\chi(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}_{\chi}))}\alpha(a_{-1})$
+where $\tau=\tau(e^{-t(1-\epsilon)\frac{k}{l}}\cdot\Psi^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})}))$. One may compute that
+$\tau=(1-\epsilon)\frac{k}{l}-\frac{1/l+1/(n-l)}{\gamma(\psi)+n/(k(n-l))}$
+and $\nu_{0}(a_{1})=1/l+1/(n-l)$. Let $\epsilon\to 0$ and we obtain
+$\dim_{H}E_{l,k}(\psi)^{c}\leq(l-k)(n-l)+\frac{n}{n/(k(n-l))+\gamma(\psi)}.$
+Now we compute a lower bound for the Hausdorff dimension of $E_{l,k}(\psi)^{c}$.
+Let $\Lambda_{0}=\bigwedge^{k}\mathbb{Z}^{n}$.
+In section 4, for any function $\phi:\mathbb{R}\to\mathbb{R}$ with $\tau(\phi)\in[0,\chi(a_{-1}))$ and any sufficiently small $\epsilon>0$, we construct a Cantor type subset $\mathbf{A}_{\infty}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ such that for any $p\in\mathbf{A}_{\infty}$ we have
+$\delta(\rho_{k}(a_{t_{i}}\cdot p)\Lambda_{0})$ $\lesssim\phi(t_{i})\cdot e^{-\frac{\epsilon}{2}t_{i}}$
+for some divergent sequence $\{t_{i}\}_{i\in\mathbb{N}}$. Note that in the construction
+$a_{t_{i}}p\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{i}}q_{i}$
+and there exist a primitive integral vector $v_{i}\in\Lambda_{0}\setminus\{0\}$ and $x_{i}\in B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})$ such that
+$\rho_{k}(a_{t_{i}}q_{i})v_{i}\in\rho(a_{t_{i}}q_{i})\Lambda_{0}\cap V_{\chi}\;(V_{\chi}=V_{\beta_{0}}),\quad\rho_{k}(x_{i})\rho_{k}(a_{t_{i}}q_{i})v_{i}\in\rho_{k}(a_{t_{i}}p)\Lambda_{0}$
+and
+$\delta(\rho_{k}(a_{t_{i}}\cdot p)\Lambda_{0})\leq\|\rho_{k}(x_{i})\rho_{k}(a_{t_{i}}q_{i})v_{i}\|\asymp e^{\chi(a_{t_{i}})}\cdot l_{i}\leq\phi(t_{i})\cdot e^{-\frac{\epsilon}{2}t_{i}}.$
+One may verify that for small $\tilde{\delta}_{0}>0$
+$\|\pi_{\chi}(\rho_{k}(x_{i})\rho_{k}(a_{t_{i}}q_{i})v_{i})\|\geq\frac{1}{2}\|\rho_{k}(x_{i})\rho_{k}(a_{t_{i}}q_{i})v_{i}\|$
+and hence
+$\|\pi^{+}(\rho_{k}(a_{t_{i}}p)v_{i})\|\geq\frac{1}{2}\|\rho_{k}(a_{t_{i}}p)v_{i}\|.$
+Moreover, by Lemma 8.3, $v_{i}$ is a pure tensor in $\bigwedge^{k}\mathbb{Z}^{n}$. Denote by $y_{i}\in X_{k}(\mathbb{Q})$ the $k$-dimensional $\mathbb{Q}$-subspaces associated to $v_{i}$.
+Now we may choose
+$\phi(t)=e^{-t\cdot\frac{k}{l}}\cdot\Psi_{\epsilon}^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})})$
+where $\Psi_{\epsilon}(u)=u^{-n/(k(n-l))}\psi(u)u^{-\epsilon}$.
+Then by Proposition 8.2(2), for $x=P\cdot g\in X_{l}(\mathbb{R})$ with $g\in\mathbf{A}_{\infty}$, there exist infinitely many $y_{i}\in X_{k}(\mathbb{Q})$ and sufficiently large $t_{i}\in\mathbb{R}_{+}$ such that
+$H(y_{i})\lesssim e^{\frac{k}{l}t_{i}}\|\rho_{k}(a_{t_{i}}g)v_{i}\|\lesssim e^{\frac{k}{l}t_{i}}\cdot\phi(t_{i})e^{-\frac{\epsilon}{2}t_{i}}\implies H(y_{i})\leq\Psi_{\epsilon}^{-1}(e^{-t_{i}(\frac{1}{l}+\frac{1}{n-l})})$
+and
+$d(y_{i},x)\lesssim e^{-t_{i}(\frac{1}{l}+\frac{1}{n-l})}\leq\Psi_{\epsilon}(H(y_{i}))=H(y_{i})^{-n/(k(n-l))}\psi(H(y_{i}))H(y_{i})^{-\epsilon}$
+which implies that $x\in E_{l,k}(\psi)^{c}$. So we have
+$\dim_{H}E_{l,k}(\psi)^{c}+\dim P\geq\dim_{H}\mathbf{A}_{\infty}+\dim\mathbf{P}_{0}$
+and
+$\dim_{H}E_{l,k}(\psi)^{c}\geq\dim X_{l}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})(\tau+\epsilon)}{\chi(a_{-1})\nu_{0}(a_{1})}$
+where $\tau=\tau(\phi)=\tau(e^{-t\cdot\frac{k}{l}}\cdot\Psi_{\epsilon}^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})}))$. One may compute that
+$\tau=\frac{k}{l}-\frac{1/l+1/(n-l)}{\gamma(\psi)+n/(k(n-l))+\epsilon}$
+and $\nu_{0}(a_{1})=1/l+1/(n-l)$ and when $\gamma(\psi)<\infty$, $\tau=\tau(\phi)<\chi(a_{-1})=k/l$. Let $\epsilon\to 0$ and we obtain that
+$\dim_{H}E_{l,k}(\psi)^{c}\geq(l-k)(n-l)+\frac{n}{n/(k(n-l))+\gamma(\psi)}.$
+Note that if $\gamma(\psi)=\infty$, then $\tau(\phi)=\chi(a_{-1})$. In this case, the following function
+$\phi(t)e^{\chi(a_{-1})t}=\Psi_{\epsilon}^{-1}(e^{-t(\frac{1}{l}+\frac{1}{n-l})})$
+is unbounded as $\psi$ is decreasing. So we can choose the Cantor-type subset $\mathbf{A}_{\infty}$ constructed in Proposition 5.6 and apply the argument above again to conclude that
+$\dim_{H}E_{l,k}(\psi)^{c}\geq(l-k)(n-l)+\frac{n}{n/(k(n-l))+\gamma(\psi)}=(l-k)(n-l).$
+This completes the proof of Theorem 1.15. ∎
+Acknowledgements. The author thanks Nicolas de Saxce, Reynold Fregoli and Barak Weiss for valuable discussions.